A Complete Bibliography of Publications in the International Journal of Parallel, Emergent and

Distributed Systems: IJPEDS

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

02 March 2022 Version 1.30

Title word cross-reference

(4n − 9)/3 [YYN10]. 0 [ESAA18, Hu18]. 1 [ESAA18, Hu18, SCK15]. 2 [ABMOK10, ABMM^{+}12, SNMSA10, SG17, SCK15, WW05]. 3 [ABN^{+}06, CG12, EMSY12, EPR06, VPS14, WL06]. g [YHL^{+}20, ZM21a]. K [LC19, WLX21, BBFN10, FY20, FMD^{+}22, Khe16, ZWMR20]. n [FY20, ZWMR20]. P [IJT^{+}21, SWW21, WLX21, SO18]. st [CT08]. t [ERA07].

-adic [SO18]. -ary [FY20, ZWMR20]. -based [BE21]. -cubes [FY20, ZWMR20].

-D [WW05]. -diagnosable [ERA07]. -dimensional [SG17]. -exclusion [BBFN10]. -extra [ZM21a]. -good-neighbour [YHL^{+}20]. -means [LC19, WLX21]. -nearest [Khe16]. -order [CT08].

1 [RCSQ09].

2 [BM07, WQZZ20]. 2008 [DS09]. 25th [Sto10].

6D [SB22].

802.11 [MBB10, SCCrL06, SK14]. abortion [LG20]. academic [AKT21]. Accelerating [FFHK19, FA08, MUIN13, UIN14]. acceleration [OMKN17, TNTI16]. accelerator [ESAA18, SEM14]. access [AWGW08, BB09a, CL06, EMERF15, MUIN13, YHY05]. accesses [ST11]. Accuracy [RCSQ09, Aba19, DSI11, GST07]. Accurately [HSL10]. achieving [BWHR07]. acoustic [HBB^{+}17]. across [Khe16, RGS06]. activation [DBD11, SK18a]. active [PB13]. actor [AGB10, DTY21]. actual [OCV15]. actuators [Vol17]. Ad [MB10, AAH08, BBFN10, Che07, DW05, DB08, HHA21, IS14, MBB10, NJ19, RGS06, SCCrL06, SK14, SS18, SKZ12, WCW11, YHY05, XM10]. adaptation [KN15, MRA^{+}06, Suz05]. Adaptive [dSOK17, BH11, BMPV14, CdO16, Ila12, MMK21, OGD11, WLWL12, WWLW18]. addresses [TLL21]. adic [SO18]. adjustment [TSK21]. administration [HR21]. advance [SCPB09]. Advanced [DA16, SSHM17]. advances [Cho08, FNB11, VS18b]. advertising [PRMF17]. affected [CB10]. against [YM15, ZYM17]. agent [FCL^{+}11, Jon16, JMA17, KX05, TP11]. agents [BGLB21, CL06, CCLT08, Kuz18]. Aggregation [ZSL22, OJ10, ZPW10]. aging [PGL16]. AGWMS [CCLT08]. AI [SSHM17]. Algorithm [TLL21, ADL16a, ADL16b, AAAA16, BCJ06, BBFN10, BS17, CSNB11, CT08, CG12, DA09, DB08, EMSY12, EDN05, ERA07, GACdM16, GKB11, GSSR17, HL16, HRF10, Hu18, IA11, JZZ11, KK18, KLMW10, Khe16, Kon18, KKK21, LL18, LLWZ20, LC19, LLW06, MBCB18, Mah11, Mic19, NDP12, SPBD05, SS18, SK18b, SR12, SPD21, SMOK06, VSB07, WWLW18, WJZK20, WLX21, XFCH08, XCW13, XZL^{+}20, YYN10, MJAT21]. Algorithmic [LR06, DP09, GG11, GWJ09, ZBHOHQ19]. algorithms [BGLB21, CSA14, DTY21, EFBAY18, EMERF15, Fer06, JJZ12, KG20, LS19, Li10a, Li10b, LCG18, LZ07, MS16, RCS19, SD07, YHT18, ZYLH09, Zha20]. alike [KF07]. all-optical [DBBC19]. allocation [ABMOK10, ABMM^{+}12, DA09, LG20]. along [LT15]. Altix [Fat10]. ambient [HBB^{+}17]. ambiguity [SG17]. amoeba [WWLW18]. among [RG10, YW11]. analysing [BRW^{+}15]. analysis [Aba15, ACY05, BMT19, BH17, BMPV14, CML10, CFK17, CSAdS15, EdM16, ERA07, GR07, HYJ16, LZ05, LZL20, SSHM17, SSS^{+}17, SKP13, SD07, WJZK20]. analytic [OMKN17]. Analytical [GS11, LL08, GSSR17]. analytics [ACV19]. Analyzing [GSSR17]. ANN [BE21]. Announcement [Ano15a]. anomaly [CSAdS15]. ant [AP18, LL18, UIN14]. antennas [LZ05]. anticipation [MJGA17]. anycast [JTZX05]. Apicalis [MBCB18]. Application [EdM16, Fat10, GG11, HSSS09, JWWZ16, Mon07]. Applications [Jon16, YR11, AZS17, AK18, CWT14, Kon18, LL08, LZG^{+}20, OGD11, PK07, SK15, SHSS09, SO18, Suz05, YW11, Yam22, Zha18, dMDBS09]. Applied [RH07, MB10]. applying [LL18]. approach [AIN13, AAK16, AOKK08, BGLB21, BH17, DEBF20, DWZ20, ECL^{+}14, ER14, FHC21, GR07, HAC21, JS20, JWZL13, KBAO19, KKKK18, LL18, LHIN09, MM15, NNRS15, PB13, PGL16, RSR15, SS13, SMOK06, SSS18, TAMM15, TP11, VBA12, WJH20, WS13, YS09, dAGG^{+}19]. approaches [Che07, DNPT12]. Approaching [LX08]. appropriate [LC15]. approximate [Nak14a]. Approximation [LCG18, GACdM16]. approximations [RCSQ09]. aquifer [CS16]. architect [Raw20]. architecture [ABN^{+}06, ABF15, AGK^{+}19, CTS^{+}11, FBC^{+}15, GB08, HLY11, JEH09, KK17, WBP16]. Architectures [AZS17, AR14, BMPV14, GPZ08, HYJ16, LR06, RS21, SRTE09, TJKY10, ZSL22, Zho06]. area [BS17, DMN21, SSZC18, WQZZ20]. arithmetic [BP09]. arrangement [WM20]. array [IJT^{+}21, SY07]. arrays [Gor06, PGFA17]. art [TNTI16, Gor06]. artificial [BE21, DSI11, PGFA17]. ary [FY20, ZWMR20]. ASIC [CFZ^{+}10]. ASM [BPV18]. ASM-based [BPV18]. aspects [GSSR17, Nos18, Yu15]. Assamese [SS13]. Assessing [CVFC19, DNPT12]. Assessment [Fat10]. assignment [CSNB11, Li10b, TV20]. assignments [Kim16]. assist [SSZC18]. assisted [CMM17, PGFA17]. associative [SO17, TAO17]. astroinformatic [KZS17]. asymmetric [SPBD05]. Asynchronous [ZSYG18, Sk22]. attacks [HHA21, YM15]. attenuation [HBB^{+}17]. attractor [SSZC18]. attractors [KKK^{+}18]. attributes [ZYW20]. attributive [KA12]. auctions [KA12]. augmented [ACL13, WB15]. Australian [AP12]. authentication [BMM12]. automata [AIHA05, CCO19, CS16, CVFC19, DEBF20, KBAO19, Mar18, TLO19, Wac19]. Automatic [Yam21a, AMuHK15, EdM16, KZS17, WLL20, Yam21b, Yam22]. automaton [AS16, CSA14, DH20, KF07, Kuz18, Mar17]. automaton-based [CSA14]. autonomic [Suz05]. Autonomous [AGB10, BS17]. availabilities [LT15]. availability [SS18]. available [SCCrL06]. Average [SK14, CB10]. avoid [CFZ^{+}10]. aware [HR21, LK08, RGS06, RS21, SS12, TBMB18, TV20]. awareness [AXM^{+}15].

backbone [AKT21]. bacterial [BGLB21]. balancing [LGW13, XCW13]. band [HBB^{+}17]. bandit [KT17]. bandwidth [JZZ11, LK08, SCCrL06]. bandwidth-efficient [LK08]. based [ADL16a, ADL16b, Aga11, AA21, AAAA16, AGM13, AXM^{+}15, BPS^{+}11, BE21, BBFN10, BPV18, BMM12, CL06, CWT14, CG12, CTS^{+}11, CSA14, CSAdS15, DA09, DMN21, DSP20, DD11, EMSY12, EN16, EG18, FVCD05, FK17, GACdM16, GR07, GBMB19, Guo06, HB14, HYJ16, HAC21, HLZ^{+}12, IS14, JTZX05, KMS10, KKK21, KK17, Kuz18, LG21, LS13, LLZX20, LLWZ20, LW21, LMS16, LC19, LZL20, MM15, MPSHH08, NGA13, NZCX21, PL19, RG10, SNMSA10, SHB17, SCJM14, SCCrL06, Sha06, SK18b, SA19, Sk22, SWW21, TLO19, TZG16, TSK21, TLL21, VRV12, VBA12, WJH20, WCY15, WWLW18, WLL20, WJZK20, WQZZ20, WS13, WM09, WM13, XM18, XFCH08, XYC^{+}15, XZL^{+}20, YHT18, YS09, YW11, YM15, YHY05, ZP21, ZKG17, ZSYG18, ZYW20, ZPW10, ZYM17, dSOK17]. basis [CG12]. batch [DNPT12]. Bayesian [HGBP19]. BDI [CCLT08]. BE [DHRCB10]. beacon [Miš07]. behavior [Mon07]. behaviour [HHY13, JJZ12, MJGA17, SHB17, SK18b, ZBHOHQ19, dMDBS09]. behavioural [VBZV21]. behaviours [Dis09, NZCX21]. benchmarks [Fat10]. benefits [AA22]. Benes [LZ07]. Best [MW08, Xia08, TJKY10]. better [JJZ12, LG20]. between [EFF^{+}19, EFBPMMOA12, KHUC20]. beyond [NA07]. biased [GBMB19]. Bibliometric [LZL20]. big [Aba15, BTK19, FNRV15, KZS17, LZL20, LLCZ16, TSR20, TAMM15, WWKK15, BMT19, TV19, Yu15, dAGG^{+}19]. bigraphical [SHB17]. binarily [Li05]. binary [AR14, AIHA05, EdM16, ESAA18]. binocular [EPR06]. bio [AKA^{+}20]. bio-inspired [AKA^{+}20]. Biocomm [BB09a]. Biocomputing [BH17]. Bioinspired [BS17]. biointerfaces [AZS17]. biological [CLM^{+}17]. Biologically [Suz05, Wal08]. Biologically-inspired [Suz05]. biology [BGLB21]. biomedical [BB09a]. Biomorphs [AS18]. biosensing [AZS17]. biosensor [SAN^{+}17]. Biosensors [Vol17]. bit [ADL16a, ADL16b]. bit-vector [ADL16a, ADL16b]. black [KN15]. black-box [KN15]. blockchain [JW22]. blockchain-powered [JW22]. Blocking [KHUC20, HSL10]. blocks [WW05]. Blood [JWWZ16]. Bluetooth [JJZ12]. Board [Ano13, Ano14, Ano15b, Ano16]. Book [Mot06]. Boolean [Bul09, Bul16, DEBF20, FK17, Pan18]. Bounded [EFBPMMOA12, ZBHOHQ19]. bounds [Che06]. box [KN15]. boxes [KG20]. brain [EMSY12]. Brief [MSTMCVL20, uDZD16]. broadcast [Che06, LS19, RG10, SOKM05, ZSYG18]. broadcast-based [ZSYG18]. Broadcasting [SKZ12, FVCD05]. browsing [PRMF17]. Bruijn [CSNB11]. Brusselator [Suz07]. BSP [GG11, Mar09]. BSP2OMP [Mar09]. buffer [SMOK06]. buffered [VRVG11]. buffers [AOKK08]. Building [CLM^{+}17]. buildings [NZCX21]. built [Pan18]. burnt [CHJ^{+}14, CKOX16, DSM21]. bus [SY07]. business [Iha08]. Byzantine [ABS21, KLMW10, Zha16]. Byzantine-fault-tolerant [KLMW10]. Byzantine-tolerant [ABS21].

C [BB09b]. cable [CR13]. caches [ST11]. caching [CZXC06]. Caledonia [MAG^{+}19]. call [TP11]. call-centre [TP11]. cameras [RP07]. can [Mar22, SSZC18]. cancellation [WB15]. canonical [HRR^{+}14]. capillarity [Nos18]. capture [Dis09]. cardiovascular [LZL20]. Carlo [HRR^{+}14, WWLW18]. carpool [DWZ20]. Cartesian [QTCM20]. case [SAN^{+}17, WvTB^{+}07]. cases [FBC^{+}15]. cash [SRTE09]. catalysis [Wal10]. catastrophe [DHRCB10]. Cauchy [LLWZ20]. causal [HRF10, OT17]. Cayley [XZY20]. Celebration [Sto10]. Cell [DHRCB10]. Cell/BE [DHRCB10]. cells [IP08]. Cellular [CS16, CSA14, KBAO19, AS16, AIHA05, CCO19, CVFC19, DSP20, DH20, DEBF20, KF07, Kuz18, Mar17, Mar18, MOZH20, TLO19, Wac19]. cellular-automaton-alike [KF07]. center [ZM21c]. centralised [SR12]. centre [TP11]. centric [HSSS09]. century [WA20]. CFD [WvTB^{+}07]. challenge [SBC^{+}05]. challenges [CLM^{+}17]. change [HB14, KBAO19]. changes [ABMM^{+}12]. channel [CYG22, LZ05]. channels [HRF10, HBB^{+}17, Li10a]. Chaos [MTSA10, PSVK20]. Chaos-enhanced [PSVK20]. chaotic [HHY13, MBCB18, SB22, VSPK18]. character [TNTI16]. characterisation [BPV18]. characteristics [BKWA21, Ben07]. charger [JCL21]. charging [JWZL13, MCYG13]. checkpointing [WM09, WM13, dSOK17]. checkpoints [WM09, WM13]. chemical [KHUC20]. chemistry [SO17]. chemotaxis [GBMB19]. chemotaxis-based [GBMB19]. Cheng [Ano17]. Chinese [LZL20]. chip [RNMQ12, SNMSA10, TJKY10, VPS14, XM18]. chip-multiprocessor [RNMQ12]. Choice [AKT21, PSVK20]. chromatin [CEF^{+}22]. Chua [KKK^{+}18]. CICQ [LLW06]. circuit [KKK^{+}18, SO19, TAO17]. circuits [WFK18]. cities [DWZ20]. city [HR21]. clairvoyant [SD07]. clans [ZSL22]. Class [VRV12, GS11, GSKD21]. Class-based [VRV12]. classical [SBC^{+}05, SBC^{+}06]. classification [KZS17, SSS^{+}17]. classifier [Bul09]. clients [TLL21]. climate [KBAO19]. clock [MJGA17]. Cloud [AA21, AAK16, AGK^{+}19, AP18, ECL^{+}14, LZWD15, Raw20, uDZD16, RMK^{+}19, SHB17, Sin17, WJH20, XYC^{+}15, dAGG^{+}19]. Cloud-based [AA21, SHB17]. cloud-hosted [uDZD16]. Clouds [TV19, ACV19]. Cluster [LLZX20, MPSHH08, AKE13, Fat10, Miš07, SRG^{+}11, ZMX^{+}13]. Cluster-based [LLZX20, MPSHH08]. clustered [WL06, ZW08]. clusterheads [ZW08]. clustering [BTK19, ZPW10]. clusters [LZG^{+}20]. cluttered [DMN21]. CMOS [WFK18]. CMOS/memristor [WFK18]. CNN [DD11]. CNN-based [DD11]. co [BB09a, YWGH13]. co-design [BB09a]. co-scheduling [YWGH13]. coastal [CS16]. code [AR14, Fer06]. codes [BP09, EdM16]. Coding [ZBHOHQ19, SKP13]. Coding-theorem [ZBHOHQ19]. coevolutionary [Wal09]. coexistence [KK17]. cognition [Luk18, Wal08]. cognitive [KT17, Wal10]. coherence [LYH^{+}20, ST11]. Cohort [KK18, SK18b, KKKK18]. collaboration [TAMM15]. collective [HSL10]. Collision [NGA13]. Collision-based [NGA13]. colony [AP18, LL18, UIN14]. coloring [LL18]. coloured [ZSP22]. combinatorial [Sto06]. combined [LLW06]. combining [SHSS09]. comfort [WvTB^{+}07]. Communication [Zha20, ABMM^{+}12, BKWA21, BB09b, CR13, DA09, EFBPMMOA12, Fer06, HRF10, HSL10, HBB^{+}17, KHUC20, LG21, Li10a, MRA^{+}06, Nas10, QK20, SOKM05, WCC05, dMDBS09]. Communication-efficient [Zha20]. communications [VPS14, MMT^{+}15]. communities [BF06]. compared [Cor21]. Comparison [MMK21, Aga11, EN16, PdP18, RW20, SEM14, YYD11, ZWMR20]. comparison-based [EN16]. compensators [GSSR17]. Compiler [Mar09]. complete [SZG^{+}20]. complex [Ila12]. complexity [CVFC19, IP08, LFL19, CDF^{+}07]. Component [XZY20, FCL^{+}11, VBZV21]. component-level [FCL^{+}11]. components [BB09b, CLM^{+}17, KKKK18, PB13]. composer [AIHA05]. composition [GG11]. Comprehensive [Aba15]. compression [AAAA16, Sha06]. compression-based [AAAA16]. computation [Akl14, Ben07, Bul16, Bur21, LLCZ16, MUIN13, NA06, NA07, Oga17, SK14, SBC^{+}05, SBC^{+}06]. Computational [WvTB^{+}07, CDF^{+}07, Akl20a, BMPV14, CEF^{+}22, Cor21, DCL17, DSI11, FNB11, SS13]. computations [BM07, Guo06]. compute [IA11]. Computer [Sha20, Höl20, JWWZ16, Nan20, Sie20]. computers [HL16, Yam20, ZZ22, CDF^{+}07]. Computing [DS09, IP08, LDB^{+}14, NA12, Saf18, ADOKM10, AAH08, AK18, Chi09, ECL^{+}14, FCL^{+}11, GJB15, GWJ09, Höl20, Jon16, Khe16, Kon18, Li10c, LC15, LX08, MTW12, MSTMCVL20, MMS^{+}17, MTSA10, NA07, NGA13, OA17, PL19, PK07, QK20, RH07, SJ17, Sch18, SHSS09, Sin17, SBC^{+}05, TV20, WA20, Wat20, Xu06, XYC^{+}15, YS09, Zha18, Zha21]. Concentric [WCW13]. Concurrent [BCL09, SCK15, YWGH13]. Conditional [YYD11, CHJ^{+}14, YZ13, YHL^{+}20]. conditions [GBMB19, SK18a]. Conference [Xia08]. configurable [BF06, JEH09]. configuring [FCL^{+}11]. conjugate [LR06, SEM14]. connected [ABMM^{+}12, IA11, LS13]. connectivities [DSM21]. connectivity [DYZ09, GHMSR12, HH13, LHL^{+}20, QTCM20, XZY20, ZM21a, ZM21b, ZM21c]. consensus [LAA13, NOdORA20, ZS18]. conservation [DA16]. Conserving [DA16]. Consistency [CZXC06, Guo06, HLY11, LT15, RGS06]. consolidation [AP18]. Constrained [KKKK18]. constraint [CTS^{+}11, WJZK20]. constraints [Che07, FVCD05, SMOK06]. construction [GKST10]. consumption [GHMSR12]. contact [LG21]. containers [NOdORA20]. contents [Raw20]. Context [MRA^{+}06, DA09, HR21]. context-aware [HR21]. contextual [IJT^{+}21]. contiguous [ABMOK10]. control [ACY05, BB09a, BP09, Bul20, CL06, CCO19, EFBPMMOA12, FBC^{+}15, JTZX05, SB22, SCJM14, TJKY10, YHY05]. controllability [DEBF20]. controllable [YHY05]. conundrum [Wal09]. Conventional [Sch17, NA07]. cooperation [PWL09, WCL14]. Cooperative [AAH08, CZXC06, CWT12, CWT14, WW17]. core [AIN13, BMPV14, CIM^{+}09, Mic19, SEM14]. cores [KN15]. correction [CWT14]. correlation [EFF^{+}19]. Corrigendum [Ano18a, Ano18b, Ano18c]. COSMOS [BRW^{+}15]. Cost [ECL^{+}14, BE21, DBBC19, RSR15, WWKK15]. cost-efficient [WWKK15]. Cost-oriented [ECL^{+}14]. count [CTS^{+}11]. Counting [RP07]. coupled [MJGA17]. coverage [CH14, DP09, LZS^{+}10, MOZH20, NDP12]. CPS [GG11]. CPU [WFK18, YW11]. CPU- [YW11]. crab [NGA13]. crash [CB10]. crash-affected [CB10]. Cray [Fat10]. creation [SRG^{+}11]. criteria [SR13]. criterion [SJ06]. Cross [GSKD21, BB09a, XM18]. cross-layer [BB09a, XM18]. Cross-project [GSKD21]. crosspoint [LLW06]. cruncher [BCL09]. cryptographic [CFZ^{+}10, KG20]. cryptography [YR11]. cryptology [DSP20]. CT [Gor06]. cube [JW06, YYN10]. cubes [ACL13, FY20, YZ13, ZWMR20]. cubic [SZG^{+}20]. cuckoo [GSSR17, LLZX20]. CUDA [AMuHK15, AAAA16]. CUDA-based [AAAA16]. cultures [CLM^{+}17]. current [RNMQ12]. customisable [WBP16]. cut [WS13]. CVE [HLY11]. cyber [Guo14, JWZL13, WCL14, XYC^{+}15]. cyber-physical [Guo14, JWZL13, WCL14]. Cyclic [QTCM20, ZM21c].

D [SNMSA10, ABMOK10, ABMM^{+}12, ABN^{+}06, CG12, EMSY12, EPR06, VPS14, WW05]. DAGs [Li05]. Data [ACV19, Che07, Dje07, Aba15, AGM13, BTK19, DTY21, EFF^{+}19, ET09, FNRV15, FHC21, GSKD21, Khe16, KZS17, LK08, LW21, LT15, LZL20, LZG^{+}20, LLCZ16, MM15, NJ19, SA19, SS19, SCPB09, TSR20, TAMM15, WWKK15, WCW13, ZM21c, ZPW10, BMT19, TV19, Yu15, dAGG^{+}19]. data-intensive [LZG^{+}20]. data-set [WCW13]. database [BF06, LH10, OMKN17, PdP18]. databases [BD14, uDZD16, VBZV21]. datapath [FFHK19]. DCCLC [YYD11]. deadlock [ADL16a, ADL16b, SR12, WW05]. deadlock-free [WW05]. Decentralised [GJB15, JWZL13, Zha20]. Decentralized [ZS18, ZSYG18]. Decision [NZCX21, KKK21, Luk18, NNRS15, NKC18, SJ06]. decision-making [KKK21, NKC18]. Decomposable [Khe16]. decomposition [BB09b, FMD^{+}22]. decrease [LLWZ20]. Deep [SOL21, TLL21, XZL^{+}20]. defect [GSKD21]. defense [CWT12]. delay [CSW11, KMS10, LG21, SK14, XC12]. delegation [CSW11]. deletion [IJT^{+}21]. delivering [FNRV15]. delivery [AA21, IBCC15, LW21]. Delta [VRV12]. demand [BRW^{+}15, LAWS14]. Dependable [LHIN09, Dis09]. deployment [JCL21, NDP12, OCV15]. derivation [EFBAY18]. describing [BB09b]. descriptors [DSP20]. Design [Li10a, Miš07, SO19, TAO17, ACV19, BB09a, BGLB21, HSSS09, KKKK18, Mah11, SR13]. designs [AAAA16]. destination [WS13]. destination-based [WS13]. Detecting [HHA21]. Detection [LYH^{+}20, MAO19, ADL16a, ADL16b, AKE13, BCJ06, BE21, CB10, CSAdS15, ERA07, HB14, IS14, NNRS15, SY07, SS17, SR12, SR13, WS13]. detector [LAA13]. Determining [BJ19, LC15]. deterministic [JS20, SCK15, Sk22]. detouring [LGW13]. Developing [SP09, Kon18, MJ21]. development [Fre16, PB13, SAN^{+}17, SPD21]. device [MMK21]. devices [BKWA21, Cor21].

Diagnosability [CQS20a, WM20, ZWMR20, FY20, RW20, YYD11, YHL^{+}20, ZM21a]. diagnosable [EDN05, ERA07]. diagnosis [ER14, EN16, EG18, HL16, YYN10]. DIBS [JZZ11]. different [KZWS19, LT15, RGS06, Vre17]. differentiation [JZZ11]. diffusion [AIHA05, SO17, TLO19]. digital [BDT07, Höl20]. dilated [LZ07]. dilemma [AS16]. dimensional [BM07, KN15, MJAT21, SG17, SA19]. dimensions [LT15]. Directing [GBMB19]. direction [LS13]. direction-based [LS13]. discharge [LVZM18]. disconnected [FHC21]. discovery [Hu18, MPSHH08]. discrete [KK18, Nak14a, TLO19, Vre17]. disease [LZL20]. diseases [LYH^{+}20]. disjoint [US16]. disorderly [KHUC20]. dispersion [CSA14]. disruption [DKN10]. dissected [OJ10]. dissemination [LK08]. distance [MUIN13]. distinguishing [EFBAY18]. Distributed [LGW13, QK20, SRTE09, WCW11, XCW13, AKE13, BD14, BTK19, Bul20, COZ08, DB08, DP09, DTY21, FNB11, GACdM16, GPZ08, Guo06, HAC21, HLZ^{+}12, IS14, JJZ12, Kim16, KX05, LLZX20, LMS16, LC15, MDJ10, PB13, SN14, SCK15, SR12, SWW21, SMOK06, WBP16, WJZK20, WvTB^{+}07, Xu06, YS09, ZP21, ZS18]. distribution [GSSR17, HWS08, RCSQ09, SWW21, ZBHOHQ19]. distributions [CdO16]. distributive [Bur21]. disturbance [LLWZ20]. dividing [CML10]. Divisible [SMOK06]. domain [BB09b]. dominating [IA11]. DOT [LLCZ16]. double [KZS21]. DQcube [ZM21a]. Drawing [ZKG17]. driven [AKT21, GKB11, IBCC15]. dual [JZZ11, JW06, YZ13]. dual-cube [JW06]. dual-cubes [YZ13]. during [MSTMCVL20]. dynamic [AP18, CFQS12, Dis09, KKKK18, KKK21, LLZX20, SPBD05, TSK21, TLL21]. Dynamical [LVZM18, ACY05, Bul09]. dynamically [BF06]. dynamics [BMB^{+}19, Car18, CCO19, KPK13, LFL19, MJGA17, SN14].

eager [RBU13]. eavesdroppers [ZYM17]. echo [Kon18]. Eddie [Ano17]. Eden [Wac19]. edge [BE21, HH13, TV20, ZM21b]. edge-connectivity [ZM21b]. edges [FY20, WM20, ZWMR20]. editor [Ano18e, Xia08]. Editorial [ATA07, ADOKM10, Ano05, Ano09, Ano13, Ano18d, DB21, Li10c, Sto10, Sto15, Ano14, Ano15b, Ano16]. editors [VS18a]. effect [ABMM^{+}12, ACY05, SG17, Vre17, WB15]. effective [Car18]. Effects [FMD^{+}22]. Efficacy [OT17]. efficiency [FMD^{+}22, LZG^{+}20, NJ19, RMK^{+}19, RCSQ09]. Efficient [CEF^{+}22, DBD11, DA16, RNMQ12, SA19, AKE13, ACV19, BCJ06, CR13, HRF10, HDK05, IS14, LK08, MUIN13, MBB10, Mic19, MJ21, SK15, SKP13, SCPB09, TBMB18, VSB07, VBA12, WWKK15, XC12, YHWZ12, Zha20]. effort [TJKY10]. eigensolver [GPZ08]. elastic [RMK^{+}19, SHB17]. election [SS18, Sk22]. Electric [MCYG13, JWZL13, MMS^{+}17, Oga17]. electrical [MBCB18, SWW21, Vol17]. electron [SO19, TAO17]. electronic [SRTE09]. element [BWHR07, BM07, KTD^{+}19]. elements [BJ19]. elimination [RCS19]. ELM [SSS18]. embedded [GHSR07]. emergence [MJGA17, ZBHOHQ19]. Emergent [QK20, Zha18, KF07, VBZV21, ZP21]. emerging [OA17, Wal09]. emotion [ZYW20]. empirical [GSKD21, SD07]. employing [BKWA21]. emulated [GST07]. emulated- [GST07]. emulation [DV08]. enabled [BMM12, Miš07]. enabling [AAH08]. encoding [AAAA16]. encounter [CWT14]. encounter-based [CWT14]. encrypt [Akl20b]. encryption [AA20, CFZ^{+}10, Raw20]. end [SK14]. end-to-end [SK14]. Energy [AKE13, CR13, HDK05, ZD19, ZPW10, ACV19, AZS17, Che06, DA16, GHMSR12, LS19, LZG^{+}20, MT08, MBB10, RSR15, RS21, SKP13, TBMB18, XC12, DA16]. energy-aware [RS21]. Energy-efficient [AKE13, CR13, ACV19]. Energy-optimal [ZPW10]. energy-saving [MT08]. engineering [KK18, WLL20]. Enhanced [CSW11, LL08, LHL^{+}20, PSVK20, WM09, YHL^{+}20]. enhancement [HGBP19, WLWL12]. Enhancing [DTY21, LMS16]. enriched [BDT07]. ensemble [HRR^{+}14]. ensembles [CMM17]. ensure [SS18]. ensuring [TSK21]. Entropy [SN14]. environment [AZS17, AK18, HSSS09, Kon18, SRG^{+}11, SMS15, WWLW18]. environment-sensitive [AK18]. environmental [JMA17]. environments [AAH08, AAK16, ATARY09, Chi09, LC15, Nas10, SCJM14, SPBD05]. enzymatic [Pri17b]. enzyme [FK17]. enzyme-based [FK17]. EOMP [MDJ10]. epigenetic [Wal10]. equipped [Ste21]. Erlang [Kar12]. erosion [Saf18]. Error [BP09, CWT14]. estimating [MAG^{+}19]. Estimation [AKA^{+}20, BPS^{+}11]. Euclidean [MUIN13]. Evaluation [MWHK20, AP12, ADOKM10, ACV19, GS11, KG20, Li10a, Li10b, Mah11, MJ21, NOdORA20, SCCrL06, SEM14, SZG^{+}20, TAO17, VRVG11, WJZK20, Yam21b, Yam22]. evaluations [MS16]. events [Iha08]. Eventually [CB10]. evolution [Bul16, OCV15]. Evolutionary [KZS17, CdO16, ERA07]. exact [ZZ22]. exactly [MDJ10]. exchange [BJ19]. exchanged [RW20]. excited [KKK^{+}18]. exclusion [BBFN10, DB08]. execution [EdM16, HLZ^{+}12, PL19]. exhaustive [SMS15, TNTI16]. expanded [ZWMR20]. experiences [PK07]. Experimental [Aga11]. experiments [CEW^{+}10, EFBAY18, FBC^{+}15]. expert [WLL20]. explained [Akl16]. explicit [MAG^{+}19]. Exploiting [MT08]. Exploration [AMuHK15]. explore [CEF^{+}22]. Exploring [Bul20, AK18]. exponential [RCSQ09]. expression [MM15]. expressions [Pri17b]. Extending [CdO16]. extensible [GB08]. extension [BSG^{+}16, CTS^{+}11, GHSR07]. Extra [HH13, LHL^{+}20, ZM21a, ZM21b]. Extracting [dMDBS09]. Extraction [Iha08, WQZZ20]. extreme [LLWZ20]. eye [EPR06].

factorization [WJZK20]. failure [LAA13]. Fair [EMERF15, TJKY10, VRV12]. fake [ZYM17]. fake-message [ZYM17]. false [LG20]. families [JWWZ16, Mar18]. Fast [BPS^{+}11, LZ07, DBBC19, HL16, LC15, SA19, XCW13, dAGG^{+}19]. Fast-Sec [dAGG^{+}19]. Fault [DYZ09, ER14, SB22, SWW21, US16, WW05, ZW08, ECL^{+}14, ERA07, EN16, EG18, JW06, KLMW10, LL08, NNRS15, SS17, YZ13, Zha16]. Fault-tolerance [ZW08]. Fault-tolerant [US16, WW05, JW06]. faults [ACL13, EDN05, ZM21b]. Faulty [ZM21a]. FDFM [AIN13]. FDR [SS17]. feasibility [ZD19]. feature [MJAT21, ZYW20]. federated [FBC^{+}15, JW22]. feed [Sha06]. feed-forward [Sha06]. FELIX [FBC^{+}15]. FEM [GST07]. Ferlo [EFF^{+}19]. ferrying [FHC21]. FET [KK17]. Fifty [Höl20]. file [AGM13, AGK^{+}19, HLZ^{+}12, WCW13, ZMX^{+}13]. filter [LLZX20, SA19]. filtering [LK08, Pri17b]. filters [DH20]. Finding [QA07, SSZC18]. fingerprint [TZG16]. fingerprint-based [TZG16]. finite [AOKK08, BWHR07, BM07, CB10, EFBAY18, KTD^{+}19, VRVG11]. finite-buffered [VRVG11]. finite-element [KTD^{+}19]. firefly [LC19, XZL^{+}20]. fitness [Ric18, ZD19]. Flexible [CTS^{+}11, YHT18, XC12]. Flexilink [MWHK20]. flooding [MD13]. flow [BM07, TJKY10]. fluidic [FK17]. FMSLPP [ZYM17]. fog [AKA^{+}20]. folksonomy [AXM^{+}15]. forecast [TB16]. forecasting [XFCH08]. forensic [BH17]. formal [Wal08]. formalise [SO18]. formation [TSK21]. forward [Sha06]. forwarding [CSW11, KMT18]. Foundations [SK18a]. FPGA [AIN13, CFZ^{+}10, Yam21a]. FPGA/ASIC [CFZ^{+}10]. FPGAs [ET09]. fractal [Ric18]. fractional [SB22]. framework [AA20, BSG^{+}16, BTK19, DP09, FBC^{+}15, HMW^{+}20, KTD^{+}19, KX05, LS19, MSM^{+}12, TZG16]. frameworks [AM11]. free [BPV18, ET09, WW05]. frequencies [MJGA17]. frequent [LC15]. friction [Nos18]. Fully [AA20, WM13]. function [DBD11, KKKK18, OT17, TAMM15]. functions [CG12, Pan18]. future [AZS17, CQS20a].

G [WL06]. GaAs [KK17]. GaAs-based [KK17]. gadgets [ZD19]. Gait [SSS^{+}17]. game [AAK16, EG18, JW22, OT17]. Garden [Wac19]. GarQ [SCPB09]. gates [APA20, Pri17b, Sch17]. gateway [KMS10]. gateway-based [KMS10]. gaze [EPR06]. GCA [JEH09]. GCC [Xia08]. gene [MM15]. general [AA20, BCL09]. generalised [DYZ09, SR12, YYN10]. Generalized [YMLC07]. generated [XZY20]. generation [AR14, BKWA21, BH11, Fer06]. generations [Chi09, SWW21]. generator [TNTI16]. Generic [GWJ09, HSSS09]. Genetic [Sin17, BGLB21, Bul09, EDN05, GKB11, SMOK06]. Genetic-variable [Sin17]. genomic [FNRV15]. geo [TV20]. geo-aware [TV20]. geographic [LS13, LGW13]. geometric [DKN10]. geothermal [SSZC18]. German [Sie20]. Getting [DV08]. Global [KMS10, CEW^{+}10, JW06, MJGA17, SG17, WLL20, ZYM17]. globally [TSK21]. Gnutella [HWS08]. good [YHL^{+}20]. Google [GB08]. governing [SN14].

GPGPUs [CS19]. GPU [ADL16a, ADL16b, AKT21, BCL09, EILB19, MUIN13, TNTI16, UIN14, YW11, Yam21b, Yam22]. GPU-based [ADL16a, ADL16b, YW11]. GPU-LMDDA [ADL16a]. GPU-OSDDA [ADL16b]. GPUs [DD11, Nak14b]. gradient [LR06, SEM14]. grammars [IJT^{+}21]. grand [SBC^{+}05]. graph [Akl20b, AGM13, DEBF20, LG21, MM15, VBZV21]. graph-based [LG21]. graph-theoretic [MM15]. graphics [ABN^{+}06, CIM^{+}09, HRR^{+}14]. graphs [CFQS12, CHJ^{+}14, CKOX16, DSM21, QTCM20, WM20, XZY20, YYD11]. Gray [AR14]. greedy [DWZ20]. grey [DBD11]. grey-level [DBD11]. grid [BH11, DA09, HB14, HLZ^{+}12, MCYG13, OGD11, SCJM14, SS17, SRG^{+}11, SCPB09, VBA12, ZPW10, COZ08, GB08, PK07, Tra11]. grid-based [DA09, ZPW10]. grids [BWHR07, DSI11, GKB11, KR17, LAWS14, LSB^{+}14]. Gridscape [GB08]. Group [LK08, LG21, SK18b, TSK21, Wac19, WCC05]. Group-aware [LK08]. groups [Kim16, Sk22]. guard [LZ05]. Guest [Li10c, Xia08]. guide [Bul16, HSSS09].

Hadoop [WCW13]. half [WA20].

Hamiltonian [DYZ09].

Hamiltonian-connectivity [DYZ09]. Hamiltonicity [DYZ09]. handoff [LZ05]. Hardware [ESAA18, KF07, LG20, OA17, GST07]. hardware-oriented [GST07]. harmony [ESAA18]. harvesting [ZD19]. hats [ZKG17]. healthy [SSS^{+}17]. heap [GACdM16]. heterogeneous [Chi09, DNPT12, GACdM16, KKK21, MS16, MT08, MPSHH08, Nas10, SPBD05, SS12]. heuristic [Li10b, Mah11]. Hidden [KKK^{+}18]. Hierarchical [BWHR07, EMSY12, MS16, HLY11, Nas10, ZYW20]. Hierarchical-based [EMSY12]. hierarchy [WCC05]. High [KR17, ECL^{+}14, HL16, LZS^{+}10, LH10, MJAT21, NZCX21, SNMSA10, SP09, SHSS09, SS18, DS09]. high-dimensional [MJAT21]. High-performance [KR17, SNMSA10, SHSS09]. high-rise [NZCX21]. Highly [ZMX^{+}13, DW05, SKZ12, Wal08]. Hilbert [HLP07]. Hilbert-order [HLP07]. History [Yam20, CL06, MSTMCVL20, Sie20]. HMM [EMSY12]. hoc [AAH08, BBFN10, Che07, DW05, DB08, HHA21, IS14, MBB10, RGS06, SCCrL06, SK14, SS18, SKZ12, YHY05, NJ19, WCW11, XM10, MB10]. home [Höl20]. homogeneous [GACdM16]. homomorphic [AA20]. honeycomb [ZYLH09]. hop [CTS^{+}11]. hop-count [CTS^{+}11]. Hopfield [ER14]. host [DSI11]. hosted [uDZD16]. HPC [ECL^{+}14, MSM^{+}12]. HPRD [LH10]. hub [RSR15]. huge [AKT21]. Human [LFL19, JWWZ16]. human-computer [JWWZ16]. Hybrid [MOZH20, BWHR07, JWWZ16, LW07, LZQL21, SS13, SPBD05, WCL14, WFK18, XFCH08, YWGH13]. hydro [HBB^{+}17]. hydrological [XZL^{+}20]. hyper [RCSQ09, SB22]. hyper-chaotic [SB22]. hyper-exponential [RCSQ09]. hyperbolic [Mar18]. hypercube [HH13, YHL^{+}20]. hypercube-like [HH13]. hypercubes [LHL^{+}20, RW20, US16]. hypergraph [MTW12]. hyperspectral [WJZK20]. hypertorus [ZSP22]. Hypothesis [NNRS15, HB14]. HyTeG} [KTD^{+}19].

I/O [RBU13]. IDA* [Mah11]. identical [DNPT12]. identification [MBCB18]. identifier [HYJ16]. identifier/locator [HYJ16]. identifying [EDN05, MM15, TLL21]. identity [BMM12]. identity-based [BMM12]. IEEE [AWGW08, MBB10, SCCrL06, SK14]. IFR [LL08]. II [GB08, SBC^{+}06]. image [CMM17, DBD11, DSP20, DD11, Gor06, KF07, LYH^{+}20, LZQL21, WJZK20]. image-based [DSP20]. images [AIHA05, WQZZ20]. IMAKA [SK15]. IMAKA-Tate [SK15]. imbalance [GSKD21]. immune [BMB^{+}19]. impact [HBS08, KBAO19, YW11]. impacts [MAG^{+}19]. implement [CCLT08]. Implementation [KK17, WJZK20, AIN13, AAAA16, ACV19, BCL09, CEF^{+}22, GG11, Mic19, SSHM17, SO17]. implementations [AA20, Nak14a]. implemented [DH20, NGA13]. Implementing [Kar12, LAA13, YHT18]. implications [NA12]. importance [BJ19]. Improve [PWL09, RBU13, SS12, YWGH13]. Improved [KG20, LZQL21, KMS10, SR12, TLO19, WLX21, Yam21b]. Improvement [ATARY09, Aba19, IBCC15]. Improving [CCO19, CWT14, DSI11, LZG^{+}20, HLY11]. in-memory [OMKN17]. incentive [WCL14]. incentives [KA12]. incident [MJGA17]. Incorporating [NOdORA20]. increasing [LT15]. indexing [Urr08]. Indoor [WvTB^{+}07, LS14, SK15]. indulgent [Sk22]. industrial [WLL20]. infection [BMB^{+}19]. infinite [ZSP22]. Influence [TTS^{+}21]. information [CVFC19, CSAdS15, JW06, NA06, Nos18]. informed [WM13]. infrastructure [FK17]. inhibition [SK18a]. initial [GBMB19, SBC^{+}06]. initialisation [WLX21]. input [APA20, LLW06]. input-crosspoint-queued [LLW06]. insertion [IJT^{+}21]. insolation [SSS18]. inspired [AKA^{+}20, SO19, SK18b, Suz05, Wal08]. instance [MJAT21]. Insulin [SAN^{+}17]. insurgent [Ila12]. integer [SB22]. integral [RNMQ12]. integrated [BRW^{+}15, KX05, LHIN09]. Integrating [CFZ^{+}10]. integration [CLM^{+}17, GB08, KMS10, LDB^{+}14]. integrative [MAG^{+}19]. Intel [EdM16]. Intelligence [SK18b, KK18, KKKK18, SMS15, NZCX21]. Intelligent [MJ21, FHC21, LWZ08]. intensive [LZG^{+}20]. inter [SK18b]. inter-group [SK18b]. interacting [Ben07]. Interactive [KPK13, WvTB^{+}07]. interconnect [ZSP22]. interconnection [CQS20a, GS11, HL16, Miš07, QA07, Ste21, VRVG11]. intermittently [LS13]. internal [VRVG11]. internal-priority [VRVG11]. International [QK20]. Internet [HYJ16, OCV15, WCC05, Zho06]. interpolations [FMD^{+}22]. interpreter [DD11]. interval [JZZ11]. intra [SK18b]. intra- [SK18b]. intree [DNPT12]. intree-shaped [DNPT12]. Introduction [Xia08]. intrusion [CS16, IS14]. invariance [Ric18]. investigation [Suz07]. ionising [SR13]. IoT [BKWA21]. IP [LW21]. IP-based [LW21]. IPART [WLL20]. IPv6 [WCY15]. IPv6-based [WCY15]. ISRL [LWZ08]. Issue [ADOKM10, Yu15, Ano06, Cho08, FNB11, Guo14, OA17, QK20, RH07, Tra11, VS18b, Xu06, ZP21]. issues [FNRV15, IP08, Miš07].

Japan [Yam20]. Jaya [Mic19]. job [HLZ^{+}12, PL19]. Journal [QK20]. journey [Akl20a]. Journeys [SBC^{+}05, SBC^{+}06].

kernel [BMPV14]. Kinect [KPK13]. knapsack [ESAA18, Hu18, RNMQ12]. knowledge [ABF15]. KpyrRec [Urr08]. kriging [MOZH20]. Kubernetes [NOdORA20].

landscape [Aba15]. landscapes [Ric18]. lane [Ste21]. language [AGM13, DD11, Yam22]. languages [SHSS09]. large [ABS21, BWHR07, CEW^{+}10, EDN05, FBC^{+}15, JWZL13, KG20, SS18, Sk22]. large-scale [ABS21, CEW^{+}10, FBC^{+}15]. last [Sha20]. latency [MWHK20]. lateral [SK18a]. Lattice [PGL16]. law [GKST10]. layer [BB09a, LMS16, YHT18, XM18]. layer-based [LMS16]. layered [SO19]. layers [DH20]. layout [WCW13]. leader [SS18, Sk22]. leads [KHUC20]. learn [Mar22]. Learning [ZCT07, Bul09, Car18, GSKD21, JW22, LWZ08, RSR15, SK18b, SSS^{+}17, SOL21, TLL21, VSB07, XZL^{+}20, ZSYG18]. less [LG20]. level [DBD11, EN16, EG18, FCL^{+}11, KZWS19]. libraries [BDT07]. library [HSSS09]. library-centric [HSSS09]. lifetime [CTS^{+}11, HWS08]. light [DA09]. like [HH13, WLX21, ZBHOHQ19]. limitations [OCV15]. limited [JW06, LFL19]. limited-global [JW06]. limits [NA07]. linear [BCL09, CIM^{+}09, GKB11, SY07, SMOK06, ZSL22, ZCT07]. Linearly [ACL13]. Link [DA16]. list [SCK15]. Listing [Sto06]. LMDDA [ADL16a]. load [DSI11, LGW13, SMOK06, XCW13]. load-balancing [XCW13]. local [DSP20, IA11, RW20, TNTI16, TSK21, YHY05]. localization [Aba19, LC19]. Localized [RCS19]. location [LS14, SWW21, YM15, ZYM17]. location-based [YM15]. location-tracking [LS14]. locator [HYJ16]. logic [APA20, FK17, KK17, Pri17b, Sch17, ZKG17]. Logical [Nos18]. Logics [Sch18, SO18]. logs [Iha08]. long [LL08]. long-running [LL08]. loop [Yam21a]. loss [IBCC15]. low [LS19, MWHK20, SNMSA10]. low-energy [LS19]. low-power [SNMSA10]. Lower [Che06].

M [RCSQ09]. MAC [BB09a, CYG22, TLL21]. machine [AP18, LZWD15, Nak14a, Nak14b, SSS^{+}17, Wal08, dSOK17]. machines [Bur21, EFBAY18, EN16, Fat10, FA08, Mar22, Pan18, Wal09, Wal10, YWGH13]. macromolecular [MMS^{+}17]. macular [LYH^{+}20]. Maeda [PGFA17].

Maeda-Makino [PGFA17]. mahjong [SSHM17]. making [KKK21, NKC18]. Makino [PGFA17]. Managed [OMKN17, NOdORA20]. management [AGM13, COZ08, GR07, Kar12, KX05, RS21, TP11, XYC^{+}15]. managing [ABF15]. MANET [IS14, CZXC06, DA16, HBS08]. MANETs [TSK21, YHWZ12]. manifolds [ZCT07]. many [ACL13, SYLS07]. many-to-one [SYLS07]. map [KKK21, MUIN13]. map-based [KKK21]. mapping [CCLT08, CWT12, CWT14, HYJ16, XM18]. MapReduce [HAC21]. Maps [GB08]. markers [MM15]. massive [HSL10]. massively [BSG^{+}16, JEH09, KF07, SEM14]. Matching [CHJ^{+}14, AGB10, CKOX16, DYZ09, Nak14a, SA19, SMS15, YMLC07]. mathematical [NKC18]. mathematics [SK18a]. matrices [HLP07]. matrix [IJT^{+}21, WCW13, WJZK20]. Maximal [ET09]. maximisation [AAK16, NDP12]. maximizing [WW17]. maximum [QA07]. MaxSAT [WJH20]. means [LC19, WLX21]. measures [Ric18]. measuring [HSL10, NJ19]. mechanical [KKKK18]. mechanism [Bul16, HHY13, LGW13, TJKY10, WCL14, ZMX^{+}13]. media [BRW^{+}15]. mediated [JMA17]. medicines [LZL20]. medium [BB09a, YHY05]. Meet [Ano18e, VS18a]. Membrane [Zha21, IP08]. memory [AS18, Car18, EMERF15, GPZ08, Kuz18, LG20, LR06, MUIN13, Nak14a, Nak14b, OMKN17, SO17, TAO17]. Memristive [VS18b, Car18]. memristor [AK18, WFK18, YHT18]. memristors [AA22, Vol17]. merging [FFHK19]. mesh [ABMOK10, ABMM^{+}12, Aga11, CG12, DBBC19, DV08, LHIN09, SNMSA10, ZW08]. mesh-based [Aga11, SNMSA10]. mesh-connected [ABMM^{+}12]. meshes [RGS06, WW05, ZYLH09]. mesoscopic [Suz07]. message [Fer06, SCK15, ZMX^{+}13, ZYM17]. message-passing [Fer06, ZMX^{+}13]. messages [BMM12]. meta [GKB11, YS09]. meta-pipeline [YS09]. meta-scheduling [GKB11]. metadata [AGM13]. metaheuristic [MBCB18]. metaheuristics [EILB19]. metatypes [GHSR07]. Method [CCLT08, CML10, KZS21, WWLW18, WLX21, Yam21a, Yam21b, Yam22]. methodology [SK18b]. metric [NJ19, SJ06]. metrics [MB10]. Mexico [MSTMCVL20]. microblogging [Fre16]. microsensor [ZPW10]. middleware [ABN^{+}06, COZ08, LDB^{+}14]. migration [LL08, LZWD15]. migration-enhanced [LL08]. Min [GACdM16]. Min-heap-based [GACdM16]. miner [Mon07]. Minimising [GHMSR12]. minimum [Che06, XC12]. mining [AMV07, BTK19, Dje07, LC15]. mise [Car18]. missing [FY20, WM20, ZWMR20]. mixed [KK18, WJZK20, WQZZ20, GST07]. mixed-precision [GST07]. MM* [WM20]. MMPP [RCSQ09]. MMPP/M/1 [RCSQ09]. mobile [ABN^{+}06, AM11, BF06, CZXC06, CL06, CSA14, CFK17, DB08, HR21, HDK05, IS14, KX05, LS13, LZ05, MDJ10, NJ19, SS18, Sin17, SKZ12, TZG16, TV20, WCC05, WL06, WCL14, Xu06]. Mobility [WCY15, Aba19, GHMSR12, RGS06]. mobility-aware [RGS06]. Model [Kuz18, AGM13, Ben07, Bul20, COZ08, CH14, DCL17, DA16, DTY21, EFF^{+}19, GS11, JEH09, JW06, JMA17, LLCZ16, MTW12, MAG^{+}19, RW20, RMK^{+}19, SHB17, SR12, TLO19, WM20, WM09, WM13, XFCH08, YYD11, YHL^{+}20, ZWMR20]. model-based [AGM13, WM09, WM13]. Modeling [Sha06, HBB^{+}17, Pri17b]. Modelling [BKWA21, BMB^{+}19, Bur21, HHY13, Ila12, Nas10, ST11, VRVG11, VBZV21, YS09, AAK16, AOKK08, DHRCB10, DH20, KBAO19, PGL16, SOKM05, ZD19]. models [BMT19, CEF^{+}22, FNB11, Iha08, KN15, Luk18, Nak14a, Nak14b, NKC18, Pan18, RGS06, Wal08]. modern [SEM14]. modes [BKWA21]. Modified [VSPK18, BPS^{+}11, ER14, GSSR17]. Modulation [SG17, SKP13]. mold [SO19]. molecular [Oga17]. molecule [Oga17]. Monetary [KA12]. monitoring [GB08, HB14, XFCH08]. monotone [WW05]. Monte [HRR^{+}14, WWLW18]. most [DV08, MM15]. motion [BPS^{+}11, Kuz18, Vre17]. motorways [AP12, AAAS^{+}13]. mould [AP12, AAAS^{+}13, Jon16, MJGA17]. movement [DTY21, KZWS19, KHUC20]. movie [Mon07]. MPSoC [RS21]. MRI [EMSY12]. Multi [APA20, AP18, DH20, JTZX05, SK18b, ADL16a, AWGW08, AKT21, BMPV14, CG12, CYG22, DW05, DB08, ERA07, GS11, GSSR17, Jon16, JMA17, KLMW10, KN15, KA12, LZS^{+}10, MTW12, Mic19, ST11, SO19, SHSS09, SEM14, Ste21, TTS^{+}21, TP11, WBP16, YHT18, YWC09]. multi-agent [Jon16, JMA17, TP11]. multi-attributive [KA12]. multi-channel [CYG22]. multi-class [GS11]. Multi-Cohort [SK18b]. multi-core [BMPV14, Mic19, SEM14]. multi-cores [KN15]. multi-dimensional [KN15]. multi-GPU [AKT21]. Multi-input [APA20]. multi-lane [Ste21]. multi-layer [YHT18]. multi-layered [SO19]. Multi-objective [AP18, MTW12]. multi-paradigmatic [SHSS09]. multi-path [DW05, YWC09]. multi-processor [ERA07]. multi-rate [AWGW08]. multi-resolution [TTS^{+}21]. multi-routing [DB08]. Multi-scale [DH20, CG12]. Multi-shared-trees [JTZX05]. multi-tenant [WBP16]. multi-tree [LZS^{+}10]. multi-type [GSSR17]. multi-unit [ADL16a]. multi-versioned [ST11]. multi-writer [KLMW10].

Multicast [RGS06, HRF10, JTZX05,

MDJ10, TSK21, YHWZ12]. multicomputer [SOKM05]. multicomputers [ABMOK10, ABMM^{+}12]. multicore [AR14, EMERF15, FFHK19]. multidimensional [Urr08]. multigrid [KTD^{+}19]. multihop [Li10a]. multilayer [EFF^{+}19, MOZH20]. Multimedia [BDT07, Dje07, GR07, HMW^{+}20]. multiphysics [BSG^{+}16]. multiple [BBFN10, GSSR17, IA11, Li10a, PSVK20, YW11, ZCT07]. multiple-choice [PSVK20]. multiplication [HLP07]. multiprocessor [RNMQ12]. multistage [GS11, Ste21, VRVG11]. multithreaded [Mah11]. MUSIG [DBD11]. mutation [LLWZ20]. mutual [DB08].

name [SA19]. named [SA19]. nano [AZS17]. nano-biointerfaces [AZS17]. nanowire [KK17]. narrowly [TLL21]. nation [KR17]. nation-sized [KR17]. native [GST07]. native- [GST07]. natural [Ben07, Saf18]. Navigation [Mar18]. nearest [Khe16]. negative [WJZK20]. neighborhood [Sin17]. neighbour [YHL^{+}20]. neighbours [Khe16]. net [GR07, KZS21, Pan18, VBZV21, ZP21, ZZ22]. net-based [GR07]. net-tools [VBZV21]. nets [SOL21, ZSP22, ZYW20]. network [AKE13, AKT21, BE21, CTS^{+}11, DSP20, DKN10, ER14, EPR06, HL16, KX05, KKK21, KK17, LZS^{+}10, MOZH20, PL19, Sha06, SR13, SO17, SS12, SWW21, Suz05, TJKY10, TAMM15, WL06, WCW11, XM18, YHT18, ZM21c, Li10c, MMT^{+}15]. network-on-chip [TJKY10]. networked [ADOKM10, EFBPMMOA12, Pri17b]. Networking [FNRV15, SA19, Yu15]. Networks [Ano06, AIN13, AGB10, AOKK08, ABS21, AK18, BB09a, BBFN10, BJ19, Bul09, Bul16, Car18, CFQS12, CSNB11, Che06, Che07, CSW11, CR13, CH14, CYG22, CQS20a, Cho08, CSA14, CFK17, CLM^{+}17, DW05, DB08, DBBC19, DP09, DV08, DYZ09, DSI11, FHC21, GS11, GKST10, GHMSR12, HHA21, HGBP19, HH13, HDK05, Ila12, IS14, IA11, JJZ12, JCL21, KMT18, LG21, LS19, LW07, LWZ08, Li10a, Li10b, LS13, LGW13, LW21, LZ05, LC19, LZ07, LHIN09, MWHK20, MMK21, MT08, MD13, MPSHH08, MOZH20, MBB10, Miš07, NNRS15, NDP12, NJ19, PL19, PWL09, QA07, RG10, RCS19, SNMSA10, SCCrL06, SK14, SOKM05, SS18, SKP13, SS17, SYLS07, Ste21, SKZ12, SZG^{+}20, TBMB18, TLL21, VRVG11, Vol17, WWKK15, WLWL12, WW17, WS13, XC12, YMLC07, YYN10, YM15, YHY05, ZSYG18, ZS18, Zha20]. networks [ZM21b, ZCT07, ZPW10, ZW08, VS18b, VRV12, XM10]. networks-on-chip [SNMSA10]. neural [AIN13, BE21, DSI11, ER14, EPR06, MOZH20, PL19, SO17, SWW21, TLL21, XM18, YHT18]. neuromorphic [YHT18]. neuron [PGFA17]. neuroscience [SK18a]. next [BH11]. next-generation [BH11]. Neyman [SR13]. NFS [RBU13]. NFV [TAMM15]. NK [Bul20]. NN [FMD^{+}22]. No [LAWS14]. node [ABS21, SCCrL06, US16, WCL14, XC12]. node-based [SCCrL06]. node-disjoint [US16]. nodes [EFBPMMOA12, LC15]. Noise [CMM17, PGFA17, HBB^{+}17, KK17, ZKG17]. Noise-assisted [PGFA17]. noise-based [ZKG17]. noise-coexistence [KK17]. noisy [ATARY09]. Non [Oga17, LLW06, MM15, SBC^{+}05, SBC^{+}06, WJZK20]. non-classical [SBC^{+}05, SBC^{+}06]. non-negative [WJZK20]. non-redundant [MM15]. Non-symmetric [Oga17]. non-uniform [LLW06]. nonblocking [HSL10]. nondeterministic [EFBAY18]. Nonlinear [MMS^{+}17, DSP20, LLWZ20, MAO19]. nonparametric [ATARY09]. nonsymmetric [GPZ08]. Nonuniversality [Akl16]. normal [SSS^{+}17]. North [Akl20a].

NoSQL [PdP18]. note [RW20]. notes [Ano18d, MSTMCVL20]. Novel [TAMM15, JWWZ16, LL18, MM15, SNMSA10, VRVG11, WCC05, WCL14]. NSF [Fat10]. NUMA [KN15]. number [BCL09, Chi09, LC15, WW17]. numerical [LSB^{+}14, PGL16].

O [RBU13]. Object

[DS09, AIHA05, SPD21]. Object-Oriented [DS09]. objective [AP18, GSSR17, MTW12]. objects [Sto06]. obstacles [DMN21, OCV15]. offline [DBBC19]. offloading [Yam21a, Yam21b, Yam22]. on-chip [VPS14]. once [MDJ10]. one [SYLS07]. online [JWZL13]. onto [XM18]. ontologies [BDT07]. OpenCL [KPK13]. OpenCOPI [LDB^{+}14]. OpenMP [Mar09]. operating [GHSR07, Ste21]. operation [BKWA21, CYG22, LT15, MCYG13]. operational [RSR15]. operations [FK17, HSL10]. operators [TAMM15]. Opportunistic [KMT18, TBMB18, YHWZ12]. Opteron [Fat10]. Optical [WWKK15, DBBC19, Li10b, LYH^{+}20, LZ07]. Optimal [Nak14a, SR13, WWLW18, ZPW10]. Optimisation [HLZ^{+}12, MJAT21, YWGH13, AMuHK15, AKA^{+}20, BPS^{+}11, CEW^{+}10, DWZ20, KZWS19, LZQL21, MMK21, Mic19, PSVK20, SJ17, SK18b, UIN14]. optimisations [LSB^{+}14]. Optimised [DA16, LG20, SMS15, DBD11, VPS14, WLX21]. Optimising [RSR15]. Optimistic [Zha16, Guo06]. optimization [JCL21, MS16]. optimizations [LR06]. optimized [Pri17b]. optimum [SYLS07]. OptiMUSIG [DBD11]. orchestration [FBC^{+}15]. order [CT08, HLP07, MAO19, SB22]. ordering [HRF10]. organisation [AGK^{+}19, GBMB19]. organised [Ila12]. organising [KKK21]. organisms [KHUC20]. organization [HDK05]. organizing [ZCT07]. Oriented [DS09, CCLT08, ECL^{+}14, LLW06, XFCH08, GST07]. oscillator [MJGA17]. OSDDA [ADL16b]. Out-of-core [CIM^{+}09]. overhead [HSL10, ST11]. overlay [LZS^{+}10, MMK21]. overlays [CFK17].

P [NZCX21, LLCZ16, SSS18]. P-DOT [LLCZ16]. P-ELM [SSS18]. p2p [MD13, LW07, LZS^{+}10, Tra11]. Pachycondyla [MBCB18]. packet [IBCC15, SYLS07]. packets [MJGA17]. padfem [BM07]. page [CML10]. painting [DMN21]. pancake [CHJ^{+}14, CKOX16, DSM21]. paper [AXM^{+}15]. Papers [Xia08, MW08]. paradigm [YHY05]. paradigmatic [SHSS09]. Parallel [AR14, BTK19, BM07, Chi09, DHRCB10, EFBAY18, Gor06, HRR^{+}14, IJT^{+}21, Mah11, PK07, QK20, SPD21, AAAA16, BCJ06, BPS^{+}11, BSG^{+}16, BB09b, EMSY12, EDN05, FNB11, JEH09, KG20, KF07, Li05, Li10c, MTW12, RNMQ12, RH07, Sha06, SEM14, SMS15, Sto06, VBZV21, Wal08, ZP21, ZSL22, dMDBS09, CDF^{+}07, DS09, HSSS09]. Parallel/High [DS09]. parallelisation [BMPV14, EdM16, FFHK19]. parallelism [ACY05, AKT21, DTY21]. parallelization [EILB19]. parameter [SN14]. parameters [MBCB18]. parking [LS14]. Partial [EG18, ER14, LVZM18]. partial-discharge [LVZM18]. partially [ET09, SD07]. participatory [KA12]. Particle [JCL21, BPS^{+}11, KPK13, LLWZ20, LZQL21, MS16, PSVK20, Vre17]. partitionable [Li05]. partitioned [Khe16]. Partitioning [AM11, LCG18, MTW12]. passing [Fer06, SCK15, ZMX^{+}13]. past [CQS20a]. Path [CL06, DW05, EdM16, YWC09]. Path-history-based [CL06]. paths [US16]. patterns [KZWS19, LVZM18, LC15, MJGA17, PRMF17]. payload [MWHK20]. Pearson [SR13]. pedestrian [NZCX21]. peeking [LAWS14]. Peer [OJ10, Aga11, CZXC06, CFK17, Fre16, LWZ08, MMK21, PWL09]. Peer-to-peer [OJ10, Aga11, CZXC06, CFK17, Fre16, LWZ08, MMK21]. penalty [KKKK18]. pentagrid [Mar17]. people [RP07]. perceptron [EFF^{+}19, MOZH20]. percolation [PGL16, RG10]. percolation-based [RG10]. PercolationNET [LZS^{+}10]. perfect [CB10]. Performance [ADOKM10, BMPV14, CFK17, DS09, ERA07, Li10b, LZ05, RG10, SOKM05, SKP13, SEM14, YW11, GST07, ABMM^{+}12, BWHR07, ECL^{+}14, GS11, HL16, HLZ^{+}12, HBS08, KMS10, KR17, Li10a, LG20, LMS16, LH10, Mah11, Miš07, PdP18, RBU13, SNMSA10, SHSS09, VRVG11, Zho06]. performance-evaluation [GS11]. performances [Nas10]. periodic [MJGA17]. Personal [Nan20, Wat20, FCL^{+}11, Yam20]. personalised [PRMF17]. perspective [AA21, AKA^{+}20, Zho06]. pervasive [FCL^{+}11]. Petri [GR07, Pan18, SPD21, VBZV21, ZP21, ZZ22, ZSP22]. Petri-object [SPD21]. Petri/Sleptsov [ZP21]. phase [WB15]. pheromone [TLO19]. phonemes [SS13]. Physarum [JMA17, MJGA17, Pan18, SJ17, Sch17]. physical [Guo14, JWZL13, WCL14]. physics [CEF^{+}22]. pipeline [AKT21, YS09]. pipelined [SY07]. pixels [WJZK20]. placement [LZWD15]. planarian [DCL17]. plane [Mar18]. planting [WQZZ20]. plants [Vol17]. platform [DNPT12, EdM16, FCL^{+}11, Fre16]. play [SSHM17]. player [SSHM17]. players [SSHM17]. PMC [YHL^{+}20]. point [AWGW08, AAAS^{+}13]. policy [AWGW08, WBP16]. policy-customisable [WBP16]. pollution [MAO19]. polycephalum [JMA17, MJGA17, Sch17]. polygonal [WW05]. polymer [CEF^{+}22]. polynomial [JS20]. polynomial-time [JS20]. population [BPS^{+}11]. population-based [BPS^{+}11]. populations [JWZL13]. portal [GB08]. positioning [AGB10, SK15]. potential [MT08]. Power [CEW^{+}10, GKST10, HB14, NJ19, SNMSA10, SCJM14]. power-law [GKST10]. powered [JW22, ZD19]. powerful [LX08]. precision [GST07]. preclusion [CHJ^{+}14, CKOX16]. predicate [CB10]. predicting [dMDBS09]. prediction [GSKD21, IBCC15, KN15, MOZH20, PL19, SSS18, XZL^{+}20]. predictions [DSI11]. Preface [CQS20b, El 19, Pri17a, ZP21, Zel18]. preprocessing [KF07]. present [CQS20a]. preservation [GHMSR12, ZYM17]. preserving [LAWS14, SK15]. preventing [WM09, WM13]. pricing [AKA^{+}20]. primitive [GG11]. prioritised [SS12]. priority [GS11, VRVG11, VRV12]. prisoner [AS16]. privacy [LAWS14, SK15, YM15, ZYM17]. privacy-preserving [LAWS14]. private [Höl20]. Privman [Pri17c]. PRNG [VSPK18]. proactive [ECL^{+}14, HBS08]. probabilistic [AS16, CH14, Cor21, Hu18]. probability [SO18, ZBHOHQ19]. problem [CT08, EN16, Hu18, JS20, LAA13, RNMQ12, TV20, UIN14]. problems [BM07, Che06, DP09, ESAA18, KK18, KT17]. process [Iha08, LZWD15]. processes [VBZV21, WBP16, XZL^{+}20]. processes-as-a-service [WBP16]. processing [AAH08, BF06, BMPV14, CMM17, DD11, FNRV15, HRR^{+}14, Pri17b, TSR20, TV19, dAGG^{+}19]. processor [AIN13, ERA07, Gor06]. processors [BE21, CIM^{+}09, EMERF15]. product [BP09, QTCM20]. programmable [WFK18]. Programming [BMT19, Wal09, AMuHK15, Bul09, CdO16, GKB11, SMOK06, Wal10, YS09, ZP21]. programming-driven [GKB11]. Programs [Mar09]. progression [HSL10]. progressive [VSPK18]. project [GSKD21]. properties [LVZM18, Oga17, YMLC07]. property [RW20]. Proportionally [TJKY10]. Proposal [CSAdS15]. Protecting [YM15]. protection [DBBC19]. protocol [BB09a, CYG22, DB08, DKN10, HBS08, JTZX05, MBB10, MDJ10, NOdORA20, Sk22, SS12, TBMB18, WCC05, WLL20, WM09, WM13]. protocols [OCV15, RG10, WLL20]. provided [KMT18]. provisioning [AM11, BH11, TSK21, WWKK15]. Proxy [Raw20]. PSO [MM15]. PSO-based [MM15]. publish [MBCK12]. publish/subscribe [MBCK12]. PUFs [AA22]. Pull [FVCD05]. Pull-based [FVCD05]. Python [BSG^{+}16].

QMDS [AGM13]. QoE [IBCC15]. QoE-driven [IBCC15]. QoS [MMK21, MB10, SS12, TBMB18, WJH20, YHY05]. QoS-based [YHY05]. quad [VRV12]. quad-priority [VRV12]. quadrilateral [AIHA05]. quadrilateral-object [AIHA05]. Qualitative [HYJ16]. quality [SP09, TSK21]. quantified [BD14]. quantitative [Vre17]. Quantum [Luk18, NA06, NA07, AS16, Cor21]. quasirandom [Chi09]. queries [BD14]. Query [uDZD16, AGM13]. queued [LLW06]. queues [RCSQ09]. queuing [VRV12].

R [CMM17]. R-SETs [CMM17]. radial [CG12]. radiation [SR13]. Raft [NOdORA20]. RAM [NM18]. random [BBFN10, GBMB19, Li05, MBB10, SG17, ST11, VSPK18]. rapid [HB14]. rate [AWGW08, WLWL12]. rational [AAAS^{+}13]. RDF [LH10]. re [CFZ^{+}10, Raw20]. re-encryption [CFZ^{+}10, Raw20]. Reaction [SO17, AIHA05]. reaction-diffusion [AIHA05]. readers [RCS19]. reading [XCW13]. real [HMW^{+}20, LCG18, PdP18, SN14, TSR20, Wal10]. real-time [HMW^{+}20, LCG18, PdP18, TSR20, Wal10]. Realisation [KZS21]. realise [Pan18]. Realising [KF07]. realization [FK17]. rechargeable [JCL21]. recognition [ATARY09, ZYW20]. recommendation [AXM^{+}15]. recommender [HR21]. reconfigurable [ET09, SY07]. reconfiguration [LZ07]. reconstruction [Gor06, TTS^{+}21, TZG16]. recording [EPR06]. recovery [DBBC19, SS17]. rectilinear [WW05]. rectilinear-monotone [WW05]. recursion [GWJ09]. recursive [FBC^{+}15, LS19, Urr08]. Redis [LLZX20]. Reducing [WB15, DBBC19, EFBAY18]. redundancy [MD13]. redundant [MM15, RCS19]. reenterable [ZSP22]. reflected [AR14]. Reflections [Cor21, Nan20, Wat20]. Reflexion [NKC18]. regeneration [DCL17]. regional [DEBF20]. register [KLMW10]. regression [ATARY09, XZL^{+}20]. regular [KLMW10, ZM21b]. regularization [TZG16]. reinforcement [LWZ08, RSR15]. related [FNRV15]. Relation [AMV07]. Relaxing [LT15]. Relay [WW17, CTS^{+}11]. relay-based [CTS^{+}11]. relaying [WL06, WLWL12]. relevance [AZS17]. relevant [MM15]. Reliability [SZG^{+}20, TSK21, BJ19, MBCK12, SS12]. Reliability-based [TSK21]. Reliable [YHWZ12, DW05, MJ21, ZMX^{+}13]. remote [WQZZ20]. Remus} [dSOK17]. Repetitions [SY07]. replication [Che07, Sin17, SS19, dSOK17]. replications [LCG18]. repositories [ABF15, AGK^{+}19]. represent [ET09]. Representation [JMA17]. Representations [Kim16]. representative [BKWA21]. repulsive [TLO19]. reputation [VBA12]. request [ABMM^{+}12]. Research [CYG22, SBC^{+}05]. reservation [LZ05]. reservations [SCPB09]. reservoir [AK18, Kon18]. resolution [SR12, TTS^{+}21]. Resource [AAK16, DA09, ADL16a, ADL16b, YW11, ZBHOHQ19]. resource-bounded [ZBHOHQ19]. resources [KBAO19, SCPB09, VBA12]. response [BMB^{+}19, CB10, JW22, LAWS14, RCSQ09]. retinal [DH20]. retransmission [HHY13]. reveal [BKWA21]. reverse [WLL20]. reversible [Sch17]. Review [Mot06, FA08, TSR20]. Revisiting [Fer06]. reward [JW22]. RFID [RCS19, WB15]. RingNet [WCC05]. rise [NZCX21]. risk [MAG^{+}19]. RNA [Bul16]. Road [LW21]. Road-safety [LW21]. robotics [TLO19]. robots [BS17]. robust [DSP20, Kar12, MBB10]. role [KZWS19]. route [WWLW18]. Routing [ZYLH09, BB09a, DW05, DB08, DKN10, HBS08, JTZX05, JW06, Li10b, LS13, LGW13, MB10, RGS06, SYLS07, Ste21, SS12, TBMB18, US16, VRVG11, WW05, YHWZ12, YWC09, DA16]. ruggedness [Ric18]. rule [AMV07]. RUN [HGBP19]. running [LL08, LZG^{+}20].

s [Höl20, KG20]. S-boxes [KG20]. safe [GHSR07]. safety [BMM12, LW21]. salesman [UIN14]. salesperson [JS20]. saltwater [CS16]. sampling [ABS21, GSKD21]. sandstone [Saf18]. satellite [EFF^{+}19, WQZZ20]. satisficing [KT17]. satisfying [Che07]. saving [CEW^{+}10, MT08]. scalability [AA21, HLY11]. Scalable [KN15, BRW^{+}15, HR21, JEH09, KTD^{+}19, TV19, WWKK15, YWC09]. Scale [Ric18, ABS21, BWHR07, CEW^{+}10, CG12, DH20, FBC^{+}15, HSL10, Suz07]. Scale-invariance [Ric18]. scan [AAAA16]. scenarios [GR07]. scéne} [Car18]. schedule [DNPT12]. Scheduling [Li05, PL19, CCO19, CH14, DWZ20, EMERF15, EFBPMMOA12, GACdM16, GKB11, GJB15, HMW^{+}20, JZZ11, JWZL13, LLW06, SPBD05, SYLS07, SD07, SCPB09, SMOK06, TSR20, VRV12, WL06, XC12, YWGH13]. scheme [AKE13, BMM12, DW05, DSP20, DBBC19, HLP07, LW07, LS13, LZ05, SS17, VRVG11, XC12, YWC09]. schemes [LG21, LGW13, uDZD16, SKP13, SRTE09, YWGH13, ZW08, CDF^{+}07]. Scholarly [AXM^{+}15]. science [Nos18]. Scientific [DS09, GWJ09, HSSS09, MSM^{+}12, SHSS09, WCW13, Wat20]. SDN [FBC^{+}15]. Sea [TB16]. seamless [SKZ12]. search [ESAA18, FMD^{+}22, GSSR17, LWZ08, LZS^{+}10, Mah11, OT17, Sin17, SMS15, TNTI16]. searching [LW07]. Sec [dAGG^{+}19]. second [MAO19]. Secure [SCJM14, LMS16, SK15, TLL21, dAGG^{+}19]. secured [VBA12]. Security [LMS16, XYC^{+}15]. Segmentation [SS13, CML10, DBD11, EMSY12, LZQL21]. Seismic [SSZC18, Zha20]. selection [AWGW08, JTZX05, MJAT21, VBA12, WW17, WWLW18]. selective [Bur21]. Self [KKK21, BBFN10, CdO16, CT08, FCL^{+}11, GBMB19, Ila12, KKK^{+}18, ZD19, ZCT07]. self-adaptive [CdO16]. self-configuring [FCL^{+}11]. self-excited [KKK^{+}18]. self-organisation [GBMB19]. self-organised [Ila12]. Self-organising [KKK21]. self-organizing [ZCT07]. self-powered [ZD19]. self-stabilizing [BBFN10, CT08]. SEM [SNMSA10]. semantic [ABF15]. semi [VSB07]. semi-supervised [VSB07]. Senegal [EFF^{+}19]. sensing [AK18, Che06, Kon18, KA12, WQZZ20]. sensitive [AK18]. sensor [Aba19, AGB10, AKE13, BB09a, Che06, CR13, CH14, CYG22, CSA14, CLM^{+}17, DP09, GHMSR12, HDK05, IA11, JCL21, KKK21, LGW13, LW21, LC19, MPSHH08, NNRS15, NDP12, RG10, SKP13, SS17, SYLS07, SR13, SKZ12, SS12, TBMB18, WS13, XC12, ZSYG18, Zha20, XM10]. Sentinel [WQZZ20]. Sentinel-2 [WQZZ20]. separation [HYJ16]. sequence [AR14, XFCH08]. sequence-based [XFCH08]. server [TV20]. Service [BH11, AGM13, ABS21, BRW^{+}15, CCLT08, GS11, LX08, MPSHH08, OGD11, TSK21, WBP16, XFCH08]. service-computing [LX08]. service-oriented [CCLT08, XFCH08]. Services [LSB^{+}14, AA21, AM11, RMK^{+}19, WJH20, Zho06]. session [HWS08]. set [WCW13]. sets [IA11, Wac19, CMM17]. SGI [Fat10]. shape [ABMM^{+}12, JMA17]. shaped [DNPT12]. shapes [ABMOK10]. shared [JTZX05, LR06]. sharing [Raw20, YW11]. Short [SSS18, JZZ11, RW20]. short-term [JZZ11]. shortest [US16]. shrinking [CML10]. shuffle [BJ19]. shuffle-exchange [BJ19]. sieve [LMS16]. signal [Nak14a, Pri17b]. signals [SB22]. Simple [Nak14b, Bul09, KLMW10, LX08, Vre17]. Simplified [LLWZ20]. Simulation [HBB^{+}17, NZCX21, XZL^{+}20, BSG^{+}16, CS16, HRR^{+}14, HSSS09, KKK^{+}18, KR17, Nan20, SN14, SPD21, TB16]. simulations [BWHR07, LSB^{+}14, WvTB^{+}07, GST07]. single [ADL16b, Oga17, SO19, TAO17]. single-electron [SO19, TAO17]. single-molecule [Oga17]. single-unit [ADL16b]. sink [ZYM17]. SIP [HHY13]. Siromoney [IJT^{+}21]. sized [KR17]. skeletons [GG11]. skip [SCK15]. SLA [DA09]. slave [Miš07]. slave-slave [Miš07]. Sleep [CH14]. Sleep-wake [CH14]. Sleptsov [KZS21, SOL21, ZP21]. Slime [AP12, AAAS^{+}13, Jon16, MJGA17, SO19]. slime-mold-inspired [SO19]. small [BPS^{+}11, MJAT21]. small-instance [MJAT21]. Smart [MMT^{+}15, DTY21, DWZ20, HR21, KR17, LS14, LAWS14, LZ05, MMK21, MCYG13, QK20, RSR15, SCJM14]. SMRS [YWC09]. snapshots [WCW11]. SOAR [MSM^{+}12]. Social [LG21, AXM^{+}15, BRW^{+}15, PWL09]. socio [SK18b]. socio-inspired [SK18b]. sockets [LMS16]. soft [SS13]. soft-computational [SS13]. software [AA21, KTD^{+}19, MJ21, SP09, Yam21a]. soil [EFF^{+}19]. solar [SSS18]. soldier [NGA13]. solution [CIM^{+}09, Hu18]. solutions [OCV15, RNMQ12]. solve [EN16]. solver [BCL09, LR06]. solvers [KTD^{+}19, GST07]. solving [ESAA18, Hu18, LAA13, ZSL22]. some [IP08]. space [BMB^{+}19, BMPV14, ET09, LZ07]. space-time [BMPV14]. sparse [BCL09, FHC21, GPZ08, HLP07]. Spatial [ZSP22, BTK19, FMD^{+}22, GBMB19, VSB07]. spatially [MAG^{+}19]. Special [ADOKM10, Ano06, Cho08, FNB11, Guo14, OA17, QK20, RH07, Tra11, VS18b, Xu06, Yu15, ZP21]. specification [ZSP22]. specifying [SHB17]. Speech [ZYW20, ATARY09]. speed [LL18, ZSL22]. speed-up [ZSL22]. SPICE [KKK^{+}18]. spikes [PGFA17]. spiking [SWW21, XM18]. splitting [SJ06, CDF^{+}07]. stabilizing [BBFN10, CT08]. stable [AGB10, CdO16]. stable-matching [AGB10]. stack [MRA^{+}06, NM18]. standard [CdO16]. Star [DSM21]. starvation [BPV18]. starvation-free [BPV18]. State [DA16, EFBAY18, Kon18]. statements [Yam21a]. states [Mar17]. static [KKKK18, SKZ12]. statistical [GKST10]. steering [Guo06, WvTB^{+}07]. Steiner [RGS06]. steps [Pri17b]. stigmergy [Vre17]. stimuli [JMA17, MJGA17]. stochastic [MMS^{+}17, NDP12, Oga17, OA17]. storage [AGK^{+}19, CFZ^{+}10, MSM^{+}12, Ste21]. storing [Raw20]. strategies [EILB19, HBS08, RBU13, SS19]. strategy [EFBPMMOA12, KT17, PSVK20]. stream [LK08, TSR20, YS09]. stream-based [YS09]. streaming [Aga11, MJGA17]. string [Nak14a]. strips [ET09]. Strong [CKOX16, RW20, SSHM17, ZZ22]. structural [GWJ09]. structure [CEF^{+}22, DSM21, ET09, SCPB09, Urr08]. structured [BMPV14, CFK17]. Study [Yam21b, Yam22, GSKD21, LL08, MD13, Mon07, SAN^{+}17, SN14, Ste21, WvTB^{+}07, ZD19]. style [SSHM17]. subjects [SSS^{+}17]. submesh [ABMOK10]. subscribe [MBCK12]. subsequence [QA07]. successive [WW17]. suites [MJ21]. sum [QA07]. Super [ZM21b]. supercomputers [WvTB^{+}07]. superhydrophobicity [Nos18]. Supernode [CS19]. supervised [VSB07]. support [EN16, GWJ09, MB10, WCY15, XZL^{+}20]. Supporting [BD14, AGM13, GS11, LZS^{+}10]. surface [Nos18, TB16]. surveillance [BS17, TLO19]. Survey [AA22, EILB19, MBCK12, uDZD16, SRTE09, SS19]. Swarm [DMN21, BPS^{+}11, FHC21, IS14, JCL21, KZWS19, LLWZ20, LZQL21, MS16, PSVK20, SMS15, TLO19]. Swarm-based [DMN21, IS14]. swarms [NGA13]. sweep [KZS21]. SWEET [Kon18]. switched [AOKK08]. switches [LLW06]. switching [BJ19]. Sybil [HHA21]. symmetric [Kim16, Oga17]. synaptic [SWW21]. synchronisation [SB22]. synchronising [AGK^{+}19]. synchronization [GR07, KKK^{+}18]. syndrome [EG18]. syndrome-based [EG18]. syndromes [ER14]. Synergies [SHSS09]. synod [Sk22]. synthesis [KZS17]. synthetic [BGLB21]. system [Aga11, AGM13, AP18, EN16, EG18, EFBPMMOA12, GHSR07, HB14, HR21, HLZ^{+}12, IS14, JW22, JWZL13, LS14, LAWS14, LLZX20, MBCB18, MMS^{+}17, SY07, SN14, SMOK06, WCW13, WB15, WLX21, XFCH08, XYC^{+}15, YHT18, ZMX^{+}13, NZCX21]. system-level [EN16, EG18]. Systemic [Ben07]. systems [ADL16a, ADL16b, ACY05, ADOKM10, BMT19, Ben07, BPV18, Bul09, CZXC06, CIM^{+}09, CFZ^{+}10, CB10, Dis09, EDN05, ERA07, FFHK19, Fer06, GACdM16, GJB15, GSSR17, Guo14, HYJ16, HLY11, IP08, Ila12, IJT^{+}21, KZWS19, Kim16, Li05, LL08, LZWD15, LMS16, MBCK12, MAO19, MDJ10, PB13, RNMQ12, SB22, SHB17, SEM14, SCK15, SWW21, TP11, Vre17, WCL14, WvTB^{+}07, WCW11, ZP21, ZSL22, Tra11, QK20].

tableaux [Kim16]. tag [XCW13]. Target [Aba19]. targeted [TLL21]. task [FFHK19, Kim16]. tasks [KF07, Li05, LCG18]. Tate [SK15]. TDCS [HMW^{+}20]. technique [AKA^{+}20, CSAdS15, DTY21, EMSY12, LMS16, MOZH20, TTS^{+}21]. techniques [MMK21, OJ10, SMS15, TSR20]. Technologies [MMT^{+}15]. technology [NM18, ZP21]. temperature [TB16]. template [SMS15]. temporal [HAC21]. tenant [WBP16]. Teraflop [BWHR07]. term [JZZ11, SSS18]. termination [BCJ06]. terms [MAO19]. ternary [WFK18]. terrorist [Ila12]. test [AAAA16, MJ21]. testing [HB14, HAC21, NNRS15, SPD21]. testing-based [HB14]. their [KHUC20, SHB17, SO18, YMLC07]. theorem [Wac19, ZBHOHQ19]. theoretic [AAK16, MM15, NNRS15]. Theoretical [Pri17b]. theory [CVFC19, CSAdS15, DEBF20, EG18, Kon18]. theory-based [CSAdS15]. Thermal [RS21]. Thermodynamic [BE21, NM18]. Thermodynamic-RAM [NM18]. thread [Sin17]. three [BM07]. three-dimensional [BM07]. threshold [KK17, SCJM14]. threshold-based [SCJM14]. throughput [NJ19]. tilings [Mar18]. Time [CFQS12, SYLS07, BMB^{+}19, BMPV14, Che07, CB10, EFBAY18, GR07, HSL10, HMW^{+}20, JS20, LCG18, LZ07, PL19, PdP18, RCSQ09, SN14, TSR20, Wal10, XFCH08]. Time-optimum [SYLS07]. Time-varying [CFQS12]. timing [FVCD05]. tissue [PGL16, WLX21]. tissue-like [WLX21]. tolerance [ECL^{+}14, LL08, Zha16, ZW08]. tolerant [ABS21, CSW11, DKN10, JW06, KLMW10, LG21, SB22, US16, WW05, YZ13]. tomographic [TTS^{+}21]. tomography [LYH^{+}20, TTS^{+}21, Zha20]. tool [WLL20]. toolkit [SRG^{+}11]. tools [Mar18, VBZV21]. topology [HBS08, SNMSA10, SJ17, VPS14]. tracker [Kon18]. trackers [EPR06]. tracking [LS14, PRMF17]. tradeoffs [RG10]. trading [AAK16]. traditional [LZL20]. traffic [BKWA21, CSAdS15, LLW06, SS12, TJKY10, WWKK15]. trailer [Mon07]. trajectories [TZG16]. trajectory [CWT12, CWT14]. transaction [BF06, Guo06]. transaction-based [Guo06]. transactional [LG20]. transformation [CS19]. transition [SKZ12]. Translating [Mar09]. Transmission [WLWL12, PGFA17, SWW21, WW17]. transmission-based [SWW21]. Transport [OCV15, MMS^{+}17]. transposition [XZY20]. travelling [JS20, UIN14]. treated [LZL20]. tree [Aga11, LZS^{+}10, OT17, RGS06]. tree-based [Aga11]. trees [JTZX05, SJ06, XZY20]. Trends [SS19]. Trichoplax [KHUC20]. True [Akl20a]. trust [VBA12]. Tunable [Wal10]. Turing [Mar22]. Twister [Fre16]. two [Mar17, Mar18, SB22, SA19]. two-dimensional [SA19]. type [GSSR17]. types [Vre17]. typing [JWWZ16].

Ubiquitous [LDB^{+}14, ADOKM10].

ultrasound [TTS^{+}21]. UMTS [KMS10]. UMTS/WLAN [KMS10]. uncertain [WWLW18]. uncertainty [NA12]. unconventional [Sch17, Sch18]. understanding [JJZ12]. unified [MSM^{+}12, MTSA10]. uniform [ABS21, LLW06]. unit [ADL16a, ADL16b, HRR^{+}14]. units [BP09]. universal [Mar17, Mar22, ZBHOHQ19]. universality [NA12]. unstructured [HLP07, LW07, LWZ08, MD13]. update [HBS08]. updating [AS16]. Uplink [WL06]. upper [Che06]. UREA [NDP12]. usage [SCJM14]. use [FBC^{+}15, GHSR07, Höl20]. useless [WM09, WM13]. user [KMT18, PRMF17]. user-provided [KMT18]. USFD [MSM^{+}12]. Using [EN16, RBU13, AGB10, AA22, AAK16, ATARY09, AP18, AK18, BPS^{+}11, BM07, BS17, CCLT08, CML10, CVFC19, DBD11, DA16, ESAA18, ER14, EG18, GSSR17, GSKD21, Gor06, IBCC15, JTZX05, Kim16, KZS21, KA12, KPK13, KKKK18, LZ05, MUIN13, PWL09, RP07, SS13, SS17, SR13, SPD21, SWW21, TNTI16, TLL21, VBZV21, WCC05, WW05, WFK18, ZYW20]. utilisation [RS21]. utility [AAK16, AKA^{+}20, GKB11]. Utilization [FK17].

value [OT17]. valued [SO18]. VANET [WCY15]. VANETs [BMM12]. variable [KK18, Sin17]. variety [Aba15]. various [Yam22]. varying [CFQS12]. VCDE [SRG^{+}11]. vector [ADL16a, ADL16b, EN16, XZL^{+}20]. vectors [AAAA16]. vehicle [MCYG13]. vehicle-to-grid [MCYG13]. vehicles [JWZL13]. vehicular [HHA21, LW21, SKZ12, YM15]. verification [Guo06, uDZD16]. versioned [ST11]. versus [RCSQ09]. vertex [LL18, QTCM20]. vertex-connectivity [QTCM20]. via [Aba19, GBMB19, LL18, SB22, Saf18, WLWL12]. video [Aga11, BDT07, IBCC15, MWHK20, Mon07, RP07]. view [AAAS^{+}13, MTSA10]. viewing [Mon07]. Virtual [LZWD15, AP18, SS17, SRG^{+}11, YWGH13, dSOK17]. virtualisation [TAMM15]. virus [BMB^{+}19]. vision [Sha20]. visual [YR11]. Vladimir [Pri17c]. volistor [APA20].

Volume [Sto10]. volunteer [GJB15, PL19]. voting [WLL20]. vulnerability [EFF^{+}19].

wake [CH14]. waLBerla [BSG^{+}16]. walk [MBB10, SG17, VSPK18]. walking [SSS^{+}17]. walks [BBFN10]. water [KBAO19]. Watermarking [CG12]. wave [MJGA17, MAO19, BMM12]. WAVE-enabled [BMM12]. wavelength [CSNB11, Li10b, LZ07]. waypoints [SBC^{+}06]. WDM [CSNB11, DBBC19, Li10b]. weakest [LAA13]. weakly [Mar17]. weaving [Kon18]. web [AM11, PRMF17, CML10]. weighted [VRV12, CDF^{+}07]. weighting [MB10]. Whale [MJAT21]. wheat [WQZZ20]. wide [HBB^{+}17]. WiFi [DV08]. wildfire [MAG^{+}19]. winter [WQZZ20]. Wireless [Ano06, XM10, AGB10, AKE13, AAH08, Che07, CTS^{+}11, CR13, CYG22, Cho08, CSA14, CLM^{+}17, DW05, DP09, IA11, JCL21, KKK21, Li10a, LGW13, LZ05, LHIN09, MPSHH08, NNRS15, NDP12, RG10, SKP13, SS17, SYLS07, SR13, SKZ12, SS12, TBMB18, WLWL12, WS13, XC12, YHY05, ZPW10]. within [DA09]. WLAN [KMS10]. WLANs [AWGW08]. workflow [CCLT08, COZ08, GJB15, Iha08, Kar12]. workflows [DA09, DNPT12, MSM^{+}12]. working [VBZV21]. workloads [OMKN17]. workshop [MW08]. Wormhole [CWT12, AOKK08, Ste21]. wormhole-switched [AOKK08]. writer [KLMW10]. WSNs [ZYM17]. WWASN2007 [MW08].

x86 [EdM16]. Xen [dSOK17]. Xen-based [dSOK17]. XNOR [APA20].

years [Höl20, MSTMCVL20, Nan20, Sha20, Sie20, Wat20]. Young [Kim16].

References

Akl:2020:FHE

[AA20] Selim G. Akl and Ibrahim Assem. Fully homomorphic encryption: a general framework and implementations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5): 493–498, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ahmad:2021:CBS

[AA21] Amro Al-Said Ahmad and Peter Andras. Cloudbased software services delivery from the perspective of scalability. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):53–68, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Aljafar:2022:SBU

[AA22] Muayad J. Aljafar and John M. Acken. Survey on the benefits of using memristors for PUFs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):40–67, 2022. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

AlQuraishi:2016:CBP

[AAAA16] Eman AlQuraishi, Eman AlDwaisan, Alaa AlSaqaa, and Imtiaz Ahmad. A CUDA-based parallel implementation of a test vectors encoding algorithm in compression-based scan designs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):280– 293, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Adamatzky:2013:MRS

[AAAS^{+}13] Andrew Adamatzky, Selim Akl, Ramon Alonso-Sanz, Wesley van Dessel, Zuwairie Ibrahim, Andrew Ilachinski, Jeff Jones, Anne V. D. M. Kayem, Genaro J. Martínez, Pedro de Oliveira, Mikhail Prokopenko, Theresa Schubert, Peter Sloot, Emanuele Strano, and Xin-She Yang. Are motorways rational from slime mould’s point of view? International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(3):230– 248, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Alsalih:2008:CAH

[AAH08] Waleed Alsalih, Selim Akl, and Hossam Hassanein. Cooperative ad hoc computing: towards enabling cooperative processing in wireless environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(1):59–79, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a789667438.

Alsarhan:2016:RTC

[AAK16] Ayoub Alsarhan and Ahmad Al-Khasawneh. Resource trading in cloud environments for utility maximisation using game theoretic modelling approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(4):319– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Abawajy:2015:CAB

[Aba15] Jemal Abawajy. Comprehensive analysis of big data variety landscape. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):5–14, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ababneh:2019:TLA

[Aba19] Ahmad A. Ababneh. Target localization accuracy improvement via sensor mobility. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):594– 614, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Alamri:2015:TAM

[ABF15] Abdullah Alamri, Peter Bertok, and Adil Fahad. Towards an architecture for managing semantic knowledge in semantic repositories. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5):411– 425, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ababneh:2012:ECP

[ABMM^{+}12] Ismail Ababneh, Saad Bani-Mohammad, Wail Mardini, Hilal Alawneh, and Mohammad Hamed. The effect of communication on the performance of allocation request shape changes in 2D mesh-connected multicomputers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(5):409– 433, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ababneh:2010:ASC

[ABMOK10] Ismail Ababneh, Saad Bani-Mohammad, and Mohamed Ould-Khaoua. All shapes contiguous submesh allocation for 2D mesh multicomputers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(5):411– 421, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Agu:2006:MAM

[ABN^{+}06] Emmanuel Agu, Kutty Banerjee, Shirish Nilekar, Oleg Rekutin, and Diane Kramer. A middleware architecture for mobile 3D graphics. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3):183– 197, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a743847333.

Anceaume:2021:BTU

[ABS21] Emmanuelle Anceaume, Yann Busnel, and Bruno Sericola. Byzantinetolerant uniform node sampling service in large-scale networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):412– 439, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Angjeli:2013:LMF

[ACL13] Ariana Angjeli, Eddie Cheng, and László Lipták. Linearly many faults in augmented cubes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(5): 475–483, 2013.

Altomare:2019:DAE

[ACV19] Albino Altomare, Eugenio Cesario, and Andrea Vinci. Data analytics for energy-efficient clouds: design, implementation and evaluation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(6):690– 705, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akl:2005:AEP

[ACY05] Selim G. Akl, Brendan J. Cordy, and Weiguang Yao. An analysis of the effect of parallelism in the control of dynamical systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(2):147–168, ???? 2005. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a713729132.

Abell:2016:GLB

[ADL16a] Stephen Abell, Nhan Do, and John Jaehwan Lee. GPU-LMDDA: a bit-vector GPU-based deadlock detection algorithm for multiunit resource systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6):562–590, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Abell:2016:GOB

[ADL16b] Stephen Abell, Nhan Do, and John Jaehwan Lee. GPU-OSDDA: a bit-vector GPU-based deadlock detection algorithm for singleunit resource systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):450–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Al-Dubai:2010:ESI

[ADOKM10] Ahmed Y. Al-Dubai, Mohamed Ould-Khaoua, and Geyong Min. Editorial: Special issue: Performance evaluation of ubiquitous computing and networked systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):237– 239, 2010. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Agarwal:2011:ECT

[Aga11] Sachin Agarwal. Experimental comparison of a tree-based and a meshbased peer-to-peer video streaming system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(2):107–120, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akkaya:2010:AAP

[AGB10] Kemal Akkaya, Ismail Guneydas, and Ali Bicak. Autonomous actor positioning in wireless sensor and actor networks using stable-matching. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6): 439–464, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Andriani:2019:ASC

[AGK^{+}19] Gil Andriani, Eduardo Godoy, Guilherme Koslovski, Rafael Obelheiro, and Mauricio Pillon. An architecture for synchronising cloud file storage and organisation repositories. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):538–555, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ames:2013:QFS

[AGM13] Sasha Ames, Maya Gokhale, and Carlos Maltzahn. QMDS: a file system metadata management service supporting a graph data model-based query language. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(2):159– 183, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Asai:2005:QOC

[AIHA05] Tetsuya Asai, Masayuki Ikebe, Tetsuya Hirose, and Yoshihito Amemiya. A quadrilateral-object composer for binary images with reaction-diffusion cellular automata. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1): 57–67, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009267.

Ago:2013:FIN

[AIN13] Yuki Ago, Yasuaki Ito, and Koji Nakano. An FPGA implementation for neural networks with the FDFM processor core approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(4):308– 320, 2013.

Athanasiou:2018:URC

[AK18] Vasileios Athanasiou and Zoran Konkoli. On using reservoir computing for sensing applications: exploring environmentsensitive memristor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):367– 386, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Arshad:2020:EFU

[AKA^{+}20] Hafsa Arshad, Hasan Ali Khattak, Zoobia Ameer, Assad Abbas, and Samee U. Khan. Estimation of fog utility pricing: a bioinspired optimisation techniques’ perspective. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):309–322, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

AlAghbari:2013:EED

[AKE13] Zaher Al Aghbari, Ibrahim Kamel, and Walid Elbaroni. Energy-efficient distributed wireless sensor network scheme for cluster detection. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):1–28, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akl:2014:WC

[Akl14] Selim G. Akl. What is computation? International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(4): 337–345, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akl:2016:NE

[Akl16] Selim G. Akl. Nonuniversality explained. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):201–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akl:2020:CJT

[Akl20a] Selim G. Akl. A computational journey in the True North. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2):132– 142, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Akl:2020:HEG

[Akl20b] Selim G. Akl. How to encrypt a graph. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6): 668–681, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ando:2021:CPM

[AKT21] Ruo Ando, Youki Kadobayashi, and Hiroki Takakura. Choice of parallelism: multi-GPU driven pipeline for huge academic backbone network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):609– 622, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Asif:2011:PFM

[AM11] Muhammad Asif and Shikharesh Majumdar. Partitioning frameworks for mobile web services provisioning. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(6):519–544, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Al-Mouhamed:2015:EAO

[AMuHK15] Mayez Al-Mouhamed and Ayaz ul Hassan Khan. Exploration of automatic optimisation for CUDA programming. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(4):309–324, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.953158.

Adda:2007:RRM

[AMV07] Mehdi Adda, Rokia Missaoui, and Petko Valtchev. Relation rule mining. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6):439–449, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a784716115.

Anonymous:2005:E

[Ano05] Anonymous. Editorial. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1):1–3, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009269.

Anonymous:2006:SIW

[Ano06] Anonymous. Special issue:

Wireless networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(2):147, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a727649554.

Anonymous:2009:E

[Ano09] Anonymous. Editorial. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(2):105–106, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2013:EB

[Ano13]  Anonymous.  Editorial

board. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6):ebi, 2013.

Anonymous:2014:EB

[Ano14]  Anonymous.  Editorial

Board. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(6):ebi, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2015:A

[Ano15a] Anonymous. Announcement. International Journal of Parallel, Emergent and Distributed Systems:

IJPEDS, 30(1):79, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2015:EB

[Ano15b]  Anonymous.  Editorial

Board. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):ebi, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2016.1123449.

Anonymous:2016:EB

[Ano16]  Anonymous.  Editorial

Board. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6):ebi, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2017:EC

[Ano17] Anonymous. Eddie Cheng.

International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):307–308, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2018:Ca

[Ano18a]  Anonymous.  Corrigen-

dum. International Journal of Parallel, Emergent and Distributed Systems:

IJPEDS, 33(5):545, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2018:Cb

[Ano18b]  Anonymous.  Corrigen-

dum. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):546, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2018:Cc

[Ano18c]  Anonymous.  Corrigen-

dum. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):547, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2018:EN

[Ano18d]  Anonymous.  Editorial

notes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):346, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Anonymous:2018:ME

[Ano18e] Anonymous. Meet editor. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):524, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Alzeidi:2008:NMA

[AOKK08] N. Alzeidi, M. Ould-Khaoua, and A. Khonsari. A new modelling approach of wormholeswitched networks with finite buffers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(1):45– 57, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a789668424.

Adamatzky:2012:SME

[AP12] Andrew Adamatzky and Mikhail Prokopenko. Slime mould evaluation of Australian motorways. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(4):275–295, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ashraf:2018:MOD

[AP18] Adnan Ashraf and Ivan Porres. Multi-objective dynamic virtual machine consolidation in the cloud using ant colony system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):103–120, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Aljafar:2020:MIV

[APA20] Muayad J. Aljafar, Marek A. Perkowski, and John M. Acken. Multi-input volistor logic XNOR gates. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):423–432, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ali:2014:PBR

[AR14] Md. Mohsin Ali and Mst. Shakila Khan Rumi. Parallel binary reflected Gray code sequence generation on multicore architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(5):513– 520, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Alonso-Sanz:2016:QPD

[AS16] Ramón Alonso-Sanz. A quantum prisoner’s dilemma cellular automaton with probabilistic updating. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(4):305–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Alonso-Sanz:2018:BM

[AS18] Ramón Alonso-Sanz. Biomorphs with memory. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):1–11, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Adamatzky:2007:E

[ATA07] Andrew Adamatzky, Christof Teuscher, and Tetsuya Asai. Editorial. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(2): 77–78, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768584107.

Amrouche:2009:ISR

[ATARY09] A. Amrouche, A. Taleb-Ahmed, J. M. Rouvaen, and M. C. E. Yagoub. Improvement of the speech recognition in noisy environments using a nonparametric regression. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):49–67, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Abusubaih:2008:NAP

[AWGW08] Murad Abusubaih, Sven Wiethoelter, James Gross, and Adam Wolisz. A new access point selection policy for multi-rate IEEE 802.11 WLANs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(4): 291–307, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a792998284.

Asabere:2015:SPR

[AXM^{+}15] Nana Yaw Asabere, Feng Xia, Qinxue Meng, Fengqi Li, and Haifeng Liu. Scholarly paper recommendation based on social awareness and folksonomy. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(3):211–232, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.904859.

Arugula:2017:ANB

[AZS17] Mary Arugula, Yuanyuan Zhang, and Alex Simonian. Architectures of nano-biointerfaces: relevance to future biosensing, environment and energy applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):3–16, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bag:2009:BCL

[BB09a] Anirban Bag and Mostafa A.

Bassiouni. Biocomm — a cross-layer medium access control (MAC) and routing protocol co-design for biomedical sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):85– 103, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Blatt:2009:CCD

[BB09b] Markus Blatt and Peter Bastian. C++ components describing parallel domain decomposition and communication. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6): 467–477, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bernard:2010:MRW

[BBFN10] Thibault Bernard, Alain Bui, Olivier Flauzac, and Florent Nolot. A multiple random walks based self-stabilizing k-exclusion algorithm in ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2):135– 152, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Baker:2006:EPT

[BCJ06] A. H. Baker, S. Crivelli, and E. R. Jessup. An efficient parallel termination detection algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4):293–301, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a747880321.

Buatois:2009:CNC

[BCL09] Luc Buatois, Guillaume Caumon, and Bruno Lévy. Concurrent number cruncher: a GPU implementation of a general sparse linear solver. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(3):205– 223, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Badia:2014:SQQ

[BD14] Antonio Badia and Michael Dobbs. Supporting quantified queries in distributed databases. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(5):421– 459, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bertini:2007:MEO

[BDT07] Marco Bertini, Alberto Del Bimbo, and Carlo Torniai. Multimedia enriched ontologies for video digital libraries. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6):407– 416, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a784716892.

Barisik:2021:TCE

[BE21] Seçkin Barisik and Ilke Ercan. Thermodynamic cost of edge detection in artificial neural network (ANN)- based processors. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3): 262–274, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bentley:2007:SCM

[Ben07] Peter J. Bentley. Systemic computation: a model of interacting systems with natural characteristics. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(2):103–121, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768590342.

Brayner:2006:MTP

[BF06] Angelo Brayner and Jos De Aguiar Moraes Filho. On mobile transaction processing in dynamically configurable mobile database communities. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3):199– 213, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a743847336.

Becerra:2021:SBA

[BGLB21] A. Gargantilla Becerra, M. Gutiérrez, and R. Lahoz-Beltra. A synthetic biology approach for the design of genetic algorithms with bacterial agents. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3):275–292, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Brocco:2011:SPN

[BH11] Amos Brocco and Béat Hirsbrunner. Service provisioning for a nextgeneration adaptive grid. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(2):85–106, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Brunelle:2017:BAF

[BH17] Erica Brunelle and Jan Halámek. Biocomputing approach in forensic analysis. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):17– 29, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bistouni:2019:DRI

[BJ19] Fathollah Bistouni and Mohsen Jahanshahi. Determining the reliability importance of switching elements in the shuffleexchange networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4):448–476, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Barakat:2021:MID

[BKWA21] Basel Barakat, Simeon Keates, Ian J. Wassell, and Kamran Arshad. Modelling IoT devices communication employing representative operation modes to reveal traffic generation characteristics. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2): 117–129, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Blazy:2007:PFE

[BM07] Stephan Blazy and Oliver Marquardt. Parallel finite element computations of three-dimensional flow problems using padfem 2. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):257– 274, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a779508700.

Bocharov:2019:MDV

[BMB^{+}19] G. Bocharov, A. Meyerhans, N. Bessonov, S. Trofimchuk, and V. Volpert. Modelling the dynamics of virus infection and immune response in space and time. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4):341– 355, 2019. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Biswas:2012:IBA

[BMM12] Subir Biswas, Jelena Misić, and Vojislav Misić. An identity-based authentication scheme for safety messages in WAVE-enabled VANETs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(6):541– 562, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Buono:2014:PAS

[BMPV14] Daniele Buono, Gabriele Mencagli, Alessio Pascucci, and Marco Vanneschi. Performance analysis and structured parallelisation of the space-time adaptive processing computational kernel on multicore architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(5):460–498, 2014. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Belcastro:2019:PMS

[BMT19] Loris Belcastro, Fabrizio Marozzo, and Domenico Talia. Programming models and systems for Big Data analysis. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(6): 632–652, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Barsi:2009:ECP

[BP09] Ferruccio Barsi and Maria Cristina Pinotti. Error control by product codes in arithmetic units. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5):407– 419, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bakwad:2011:FME

[BPS^{+}11] K. M. Bakwad, S. S. Pattnaik, B. S. Sohi, S. Devi, Sastry V. R. S. Gollapudi, Ch. Vidya Sagar, and P. K. Patra. Fast motion estimation using small population-based modified parallel particle swarm optimisation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(6):457– 476, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bianchi:2018:ABC

[BPV18] Alessandro Bianchi, Sebastiano Pizzutilo, and Gennaro Vessio. An ASMbased characterisation of starvation-free systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):35–51, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Burnap:2015:CTI

[BRW^{+}15] Peter Burnap, Omer Rana, Matthew Williams, William Housley, Adam Edwards, Jeffrey Morgan, Luke Sloan, and Javier Conejero. COSMOS: Towards an integrated and scalable service for analysing social media on demand. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(2):80–100, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.902057.

Bontzorlos:2017:BAA

[BS17] Tilemachos Bontzorlos and Georgios Ch. Sirakoulis. Bioinspired algorithm for area surveillance using autonomous robots. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):368–385, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bauer:2016:PEM

[BSG^{+}16] Martin Bauer, Florian Schornbaum, Christian Godenschwager, Matthias Markl, Daniela Anderl, Harald Köstler, and Ulrich Rüde. A Python extension for the massively parallel multiphysics simulation framework waLBerla. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6):529–542, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bendechache:2019:PDC

[BTK19] Malika Bendechache, A-Kamel Tari, and M-Tahar Kechadi. Parallel and distributed clustering framework for big spatial data mining. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(6):671– 689, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bull:2009:DGP

[Bul09] Larry Bull. On dynamical genetic programming: simple Boolean networks in learning classifier systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5):421– 442, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bull:2016:EBN

[Bul16] Larry Bull. On the evolution of Boolean networks for computation: a guide RNA mechanism. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(2):101–113, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bull:2020:EDC

[Bul20] Larry Bull. Exploring distributed control with the NK model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):413– 422, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Burgin:2021:MDC

[Bur21] Mark Burgin. Modelling distributive computation by selective machines. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):395–411, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Bergen:2007:HHG

[BWHR07] B. Bergen, G. Wellein, F. Hülsemann, and U. Rijde. Hierarchical hybrid grids: achieving Teraflop performance on large scale finite element simulations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):311– 329, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a779508658.

Caravelli:2018:MSM

[Car18] Francesco Caravelli. The mise en scéne of memristive networks: effective memory, dynamics and learning. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):350– 366, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chu:2010:EPP

[CB10] Chunbo Chu and Monica Brockmeyer. Eventually perfect predicate detection in crash-affected finite average response time systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):37– 50, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cao:2008:MUB

[CCLT08] Lei Cao, Jian Cao, Minglu Li, and Feilong Tang.

Method of using BDI agents to implement serviceoriented workflow mapping in AGWMS. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):171– 195, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a791169318.

Carvalho:2019:ICA

[CCO19] Tiago I. Carvalho, Murillo G.

Carneiro, and Gina M. B. Oliveira. Improving cellular automata scheduling through dynamics control. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):115– 141, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

s:2007:CCW

[CDF^{+}07] P. Csomós, I. Dimov, I. Faragó, Á. Havasi, and T. Z. Ostromsky. Computational complexity of weighted splitting schemes on parallel computers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3):137–147, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

                                        smpp/content~content=
                                        a776022539.

Carvalho:2016:ESE

[CdO16] Leopoldo B. Carvalho and Pedro P. B. de Oliveira. Extending standard evolutionary programming with self-adaptive stable distributions. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(2):114– 142, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Conte:2022:ECI

[CEF^{+}22] Mattia Conte, Andrea Esposito, Luca Fiorillo, Raffaele Campanile, Carlo Annunziatella, Alfonso Corrado, Maria Gabriella Chiariello, Simona Bianco, and Andrea M. Chiariello. Efficient computational implementation of polymer physics models to explore chromatin structure. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):91–102, 2022. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cao:2010:PSE

[CEW^{+}10] Zhenwei Cao, David R. Easterling, Layne T. Watson, Dong Li, Kirk W. Cameron, and Wu-Chun Feng. Power saving experiments for large-scale global optimisation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(5):381–400, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chowdhury:2017:PAS

[CFK17] Farida Chowdhury, Jamie Furness, and Mario Kolberg. Performance analysis of structured peerto-peer overlays for mobile networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5): 522–548, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Casteigts:2012:TVG

[CFQS12] Arnaud Casteigts, Paola Flocchini, Walter Quattrociocchi, and Nicola Santoro. Time-varying graphs and dynamic networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(5):387–408, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2010:IFA

[CFZ^{+}10] Lanxiang Chen, Dan Feng, Yu Zhang, Lingfang Zeng, and Zhongying Niu. Integrating FPGA/ASIC into cryptographic storage systems to avoid re-encryption. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2):105–122, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Che:2012:WAM

[CG12] Xiangjiu Che and Zhanheng Gao. Watermarking algorithm for 3D mesh based on multi-scale radial basis functions. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(2): 133–141, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2014:SWS

[CH14] Pingsheng Chen and Weidong Hu. Sleep-wake up scheduling with probabilistic coverage model in sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(1):1–16, 2014.

Chelius:2006:LUB

[Che06] Guillaume Chelius. Lower and upper bounds for minimum energy broadcast and sensing problems in sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(6):405– 422, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768573272.

Chen:2007:DRA

[Che07] Xiao Chen. Data replication approaches for ad hoc wireless networks satisfying time constraints. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3): 149–161, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a776022348.

Chi:2009:PQN

[Chi09]  Hongmei Chi.  Paral-

lel quasirandom number generations for heterogeneous computing environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):21– 29, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cheng:2014:MPC

[CHJ^{+}14] Eddie Cheng, Philip Hu, Roger Jia, László Lipták, Brian Scholten, and James Voss. Matching preclusion and conditional matching preclusion for pancake and burnt pancake graphs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(5):499–512, 2014. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chou:2008:SIA

[Cho08] Chun Tung Chou. Special issue on advances in wireless networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(4):289, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a793005333.

Castillo:2009:CSL

[CIM^{+}09] Maribel Castillo, Francisco D. Igual, Mercedes Marqués, Rafael Mayo, Enrique S. Quintana-Ortí, Gregorio Quintana-Ortí, Rafael Rubio, and Robert van de Geijn. Out-of-core solution of linear systems on graphics processors. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6):521–538, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cheng:2016:SMP

[CKOX16] Eddie Cheng, Justin T.

Kelm, Roi Orzach, and Brian Xu. Strong matching preclusion of burnt pancake graphs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3): 220–??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cao:2006:PHB

[CL06] Chun Cao and Jian Lu.

Path-history-based access control for mobile agents. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3):215–225, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a743847335.

Christophe:2017:BWS

[CLM^{+}17] François Christophe, Teemu Laukkarinen, Tommi Mikkonen, Jonathan Massera, and Vafa Andalibi. Building wireless sensor networks with biological cultures: components and integration challenges. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):56–73, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cao:2010:SMW

[CML10] Jiuxin Cao, Bo Mao, and Junzhou Luo. A segmentation method for Web page analysis using shrinking and dividing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2):93–104, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cervera:2017:NAI

[CMM17] J. Cervera, J. A. Manzanares, and S. Mafe. Noise assisted image processing by ensembles of R-SETs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):295– 305, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Corrente:2021:RPC

[Cor21] Giuseppe Corrente. Reflections on probabilistic compared to quantum computational devices. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3): 251–261, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2008:DWM

[COZ08] Xiaowu Chen, Haifeng Ou, and Qinping Zhao. A distributed workflow management model for Grid middleware. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):107– 120, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a791169761.

Cheng:2020:DIN

[CQS20a] Eddie Cheng, Ke Qiu, and Zhizhang Shen. Diagnosability of interconnection networks: past, present and future. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):2–8, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cheng:2020:P

[CQS20b] Eddie Cheng, Ke Qiu, and Zhizhang Shen. Preface. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):1, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2013:EEC

[CR13] Xiao Chen and Neil C.

Rowe. Energy-efficient communication in wireless cable sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(4):289–307, 2013.

Chatziagorakis:2016:CAS

[CS16] Prodromos Chatziagorakis and Georgios Ch. Sirakoulis. Cellular automata simulation of saltwater intrusion in coastal aquifer. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6):517– 528, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2019:STG

[CS19] Yong Chen and Weijia Shang. Supernode transformation on GPGPUs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(2):181–202, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Choudhury:2014:CAB

[CSA14] Salimur Choudhury, Kai Salomaa, and Selim G. Akl. Cellular automatonbased algorithms for the dispersion of mobile wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(2):147–177, 2014.

Cuadra-Sanchez:2015:PNI

[CSAdS15] Antonio Cuadra-Sanchez, Javier Aracil, and Javier Ramos de Santiago. Proposal of a new information theorybased technique based on traffic anomaly detection analysis. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6): 464–477, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1044002.

Chatterjee:2011:WAA

[CSNB11] Monish Chatterjee, Swagato Sanyal, Mita Nasipuri, and Uma Bhattacharya. A wavelength assignment algorithm for de Bruijn WDM networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(6):477–491, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chen:2011:EDF

[CSW11] Xiao Chen, Jian Shen, and Jie Wu. Enhanced delegation forwarding in delay tolerant networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):331–345, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chaudhuri:2008:SSA

[CT08] Pranay Chaudhuri and Hussein Thompson. A selfstabilizing algorithm for the st-order problem. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3):219–234, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a790726562.

Chen:2011:FAR

[CTS^{+}11] Chi-Tung Chen, Sirin Tekinay, Cem Saraydar, Hsing-Chung Chen, Ming-Yuan Hsieh, and Jyu-Wei Wang. Flexible architecture of relay-based wireless network for network lifetime extension with hop-count constraint. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(2):121– 148, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chliamovitch:2019:ACC

[CVFC19] Gregor Chliamovitch, Lino Velasquez, Jean-Luc Falcone, and Bastien Chopard.

Assessing complexity in cellular automata using information theory. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1): 142–160, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chang:2012:WDC

[CWT12] Wei Chang, Jie Wu, and Chiu C. Tan. Wormhole defense for cooperative trajectory mapping. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(5): 459–480, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chang:2014:ICT

[CWT14] Wei Chang, Jie Wu, and Chiu C. Tan. Improving cooperative trajectory mapping applications with encounter-based error correction. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(1):68– 89, 2014.

Chen:2022:RMC

[CYG22] Yang Chen, Shuping Yuan, and Fenfei Gu. Research on the multi-channel operation of MAC protocol in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):187– 197, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Cao:2006:CCC

[CZXC06] J. Cao, Y. Zhang, L. Xie, and G. Cao. Consistency of cooperative caching in mobile peer-to-peer systems over MANET. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3): 151–168, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a743847337.

Dang:2009:RAA

[DA09] Minh Quan Dang and Jörn Altmann. Resource allocation algorithm for light communication grid-based workflows within an SLA context. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):31– 48, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dhanalakshmi:2016:EEC

[DA16] N. Dhanalakshmi and P. Alli. Efficient energy conservation in MANET using Energy Conserving Advanced Optimised Link State Routing model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):469–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

dosAnjos:2019:FSA

[dAGG^{+}19] Julio C. S. dos Anjos, Tatiana Galibus, Cláudio F. R. Geyer, Gilles Fedak, João Paulo C. L. Costa, Rubem Pereira, and Edison Pignaton de Freitas. Fast-Sec: an approach to secure Big Data processing in the cloud. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(3):272–287, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Derhab:2008:DME

[DB08] Abdelouahid Derhab and Nadjib Badache. A distributed mutual exclusion algorithm over multirouting protocol for mobile ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3): 197–218, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

                                        smpp/content~content=
                                        a790632848.

Djemame:2021:E

[DB21] Karim Djemame and Mohamed Bettaz. Editorial. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):51– 52, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dey:2019:OSR

[DBBC19] Vishal Dey, Abhishek Bandyopadhyay, Uma Bhattacharya, and Monish Chatterjee. An offline scheme for reducing cost of protection in all-optical WDM mesh networks with fast recovery. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):572– 593, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

De:2011:EGL

[DBD11] Sourav De, Siddhartha Bhattacharyya, and Paramartha Dutta. Efficient grey-level image segmentation using an optimised MUSIG (OptiMUSIG) activation function. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(1): 1–39, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

De:2017:CMP

[DCL17] Abhishek De, V. Srinivasa Chakravarthy, and Michael Levin. A computational model of planarian regeneration. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):331–347, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dolan:2011:CBL

[DD11] Ryanne Dolan and Guilherme DeSouza. CNNbased language and interpreter for image processing on GPUs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(3): 207–222, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dridi:2020:GTA

[DEBF20] S. Dridi, S. El Yacoubi, F. Bagnoli, and A. Fontaine. A graph theory approach for regional controllability of Boolean cellular automata. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5):499–513, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Devillard:2020:MSF

[DH20] F. Devillard and B. Heit. Multi-scale filters implemented by cellular automaton for retinal layers modelling. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6):617– 640, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dehne:2010:PCM

[DHRCB10] Frank Dehne, Glenn Hickey, Andrew Rau-Chaplin, and Mark Byrne. Parallel catastrophe modelling on a Cell/BE. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(5):401– 410, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Distefano:2009:HCD

[Dis09] Salvatore Distefano. How to capture dynamic behaviours of dependable systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(2):127– 150, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Djeraba:2007:DMM

[Dje07] Chabane Djeraba. Data mining from multimedia. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6):405– 406, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a784716816.

Du:2010:GRP

[DKN10] Jingzhe Du, Evangelos Kranakis, and Amiya Nayak. A geometric routing protocol in disruption tolerant network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6):489–508, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

deMello:2009:EPC

[dMDBS09] Rodrigo Fernandes de Mello, Evgueni Dodonov, Ricardo Bertagna, and Luciano José Senger. Extracting and predicting the communication behaviour of parallel applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(3):225–242, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Das:2021:SBP

[DMN21] Deepanwita Das, Srabani Mukhopadhyaya, and Debashis Nandi. Swarm-based painting of an area cluttered with obstacles. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(4):359–379, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Diakite:2012:ANA

[DNPT12] Sékou Diakité, Jean-Marc Nicod, Laurent Philippe, and Lamiel Toch. Assessing new approaches to schedule a batch of identical intreeshaped workflows on a heterogeneous platform. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(1):79–107, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dhawan:2009:DAF

[DP09] Akshaye Dhawan and Sushil K. Prasad. A distributed algorithmic framework for coverage problems in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):331– 348, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Davis:2009:PHP

[DS09] Kei Davis and Jöerg Striegnitz. Parallel/High Performance Object-Oriented Scientific Computing 2008. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6):463–465, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Duy:2011:IAH

[DSI11] Truong Vinh Truong Duy, Yukinori Sato, and Yasushi Inoguchi. Improving accuracy of host load predictions on computational grids by artificial neural networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(4): 275–290, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dilixiati:2021:SSC

[DSM21] Subinur Dilixiati, Eminjan Sabir, and Jixiang Meng. Star structure connectivities of pancake graphs and burnt pancake graphs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):440–448, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

daSilva:2017:ARA

[dSOK17] Marcelo Pereira da Silva, Rafael Rodrigues Obelheiro, and Guilherme Piegas Koslovski. Adaptive Remus: adaptive checkpointing for Xen-based virtual machine replication. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):348–367, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dehshibi:2020:RIB

[DSP20] Mohammad Mahdi Dehshibi, Jamshid Shanbehzadeh, and Mir Mohsen Pedram. A robust image-based cryptology scheme based on cellular nonlinear network and local image descriptors. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5):514–534, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Doroshenko:2021:EPD

[DTY21] Anatoliy Doroshenko, Eugene Tulika, and Olena Yatsenko. Enhancing parallelism of distributed algorithms with the actor model and a smart data movement technique. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):565–578, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Djukic:2008:GMW

[DV08] Petar Djukic and Shahrokh Valaee. Getting the most of WiFi mesh networks with 802.16 mesh emulation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(6):461–475, 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dai:2005:HRM

[DW05] Fei Dai and Jie Wu. A highly reliable multi-path routing scheme for ad hoc wireless networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4): 205–219, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648752.

Duan:2020:GAC

[DWZ20] Yubin Duan, Jie Wu, and Huanyang Zheng. A greedy approach for carpool scheduling optimisation in smart cities. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5):535–549, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dong:2009:FHF

[DYZ09] Qiang Dong, Xiaofan Yang, and Juan Zhao. Fault Hamiltonicity and fault Hamiltonian-connectivity of generalised matching networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5):455– 461, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Egwutuoha:2014:COP

[ECL^{+}14] Ifeanyi P. Egwutuoha, Shiping Chen, David Levy, Bran Selic, and Rafael Calvo. Cost-oriented proactive fault tolerance approach to high performance computing (HPC) in the cloud. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(4):363– 378, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Eberle:2016:AEP

[EdM16] André M. Eberle and Rodrigo F. de Mello. Application execution path analysis for the automatic parallelisation of binary codes in the Intel x86 platform. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):405–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Elhadef:2005:PGA

[EDN05] Mourad Elhadef, Shantanu Das, and Amiya Nayak. A parallel genetic algorithm for identifying faults in large diagnosable systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(2):113– 125, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a713729130.

El-Fakih:2018:PAR

[EFBAY18] Khaled El-Fakih, Gerassimos Barlas, Mustafa Ali, and Nina Yevtushenko. Parallel algorithms for reducing derivation time of distinguishing experiments for nondeterministic finite state machines. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2): 197–210, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Esquivel-Flores:2012:BCB

[EFBPMMOA12] O. Esquivel-Flores, H. Benítez-Pérez, P. Méndez-Monroy, and J. Ortega-Arjona. Bounded communication between nodes of a networked control system as a strategy of scheduling. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(6):481–502, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

ElYacoubi:2019:MPM

[EFF^{+}19] Samira El Yacoubi, Mireille Fargette, Abdoulaye Faye, Waldir de Carvalho Junior, Thérèse Libourel, and Maud Loireau. A multilayer perceptron model for the correlation between satellite data and soil vulnerability in the Ferlo, Senegal. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):3–12, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Elhadef:2018:PSB

[EG18] Mourad Elhadef and Sofiane Grira. Partial syndromebased system-level fault diagnosis using game theory. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):69– 86, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Essaid:2019:GPS

[EILB19] Mokhtar Essaid, Lhassane Idoumghar, Julien Lepagnot, and Mathieu Brévilliers. GPU parallelization strategies for metaheuristics: a survey. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):497– 522, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

ElYacoubi:2019:P

[El 19] Samira El Yacoubi. Preface. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):1–2, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

El-Moursy:2015:FMA

[EMERF15] Ali A. El-Moursy, Walid El-Reedy, and Hossam A. H. Fahmy. Fair memory access scheduling algorithms for multicore processors. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(4):286–308, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.922560.

El-Moursy:2012:HBP

[EMSY12] Ali El-Moursy, Sheif Saif, and Akmal Younis. Hierarchicalbased parallel technique for HMM 3D MRI brain segmentation algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(4):297–316, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Elhadef:2016:USV

[EN16] Mourad Elhadef and Amiya Nayak. Using support vector machines to solve the comparison-based systemlevel fault diagnosis problem. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(2):176– 200, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Essig:2006:NNG

[EPR06] Kai Essig, Marc Pomplun, and Helge Ritter. A neural network for 3D gaze recording with binocular eye trackers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(2):79– 95, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

smpp/content~content=
a727649555.

Elhadef:2014:FDU

[ER14] Mourad Elhadef and Lotfi Ben Romdhane. Fault diagnosis using partial syndromes: a modified Hopfield neural network approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(2): 119–146, 2014.

Elhadef:2007:PAE

[ERA07] M. Elhadef, L. B. Romdhane, and B. Ayeb. Performance analysis of an evolutionary algorithm for fault detection in t-diagnosable multi-processor systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(5):387–404, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a782139046.

El-Shafei:2018:HAS

[ESAA18] Mohammed El-Shafei, Imtiaz Ahmad, and Mohammad Gh. Alfailakawi. Hardware accelerator for solving 0–1 knapsack problems using binary harmony search. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):87– 102, 2018. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Elbidweihy:2009:MSD

[ET09] Mostafa Elbidweihy and Jerry L. Trahan. Maximal strips data structure to represent free space on partially reconfigurable FP-GAs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):349– 366, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fraser:2008:AMR

[FA08] Robert Fraser and Selim G. Akl. Accelerating machines: a review. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(1): 81–104, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a789666600.

Fatoohi:2010:ANA

[Fat10] Rod Fatoohi. Assessment of NSF application benchmarks on SGI Altix machines and Cray Opteron cluster. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):315– 329, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fernandez:2015:ROC

[FBC^{+}15] Carolina Fernandez, Carlos Bermudo, Gino Carrozzo, Roberto Monno, Bartosz Belter, Kostas Pentikousis, Umar Toseef, Tomohiro Kudoh, Atsuko Takefusa, Jason Haga, Bart Puype, and Jin Tanaka. A recursive orchestration and control framework for largescale, federated SDN experiments: the FELIX architecture and use cases. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):428–446, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1044003.

Feng:2011:CLS

[FCL^{+}11] Yuhong Feng, Jiannong Cao, Ivan Chuen Ho Lau, Zhong Ming, and Joseph Kee-Yin Ng. A componentlevel self-configuring personal agent platform for pervasive computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(3):223–238, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ferner:2006:RCC

[Fer06]  Clayton Ferner.  Re-

visiting communication code generation algorithms for message-passing systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(5):323– 344, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747881464.

Fazlali:2019:ADM

[FFHK19] Mahmood Fazlali, Mohammad K. Fallah, Naemeh Hosseinpour, and Ali Katanforoush. Accelerating datapath merging by task parallelisation on multicore systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5): 615–628, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fraser:2021:SIA

[FHC21] Bradley Fraser, Robert Hunjet, and Andrew Coyle. A swarm intelligent approach to data ferrying in sparse disconnected networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):630– 651, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fratto:2017:UFI

[FK17] Brian E. Fratto and Evgeny Katz. Utilization of a fluidic infrastructure for the realization of enzymebased Boolean logic operations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):139– 156, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fan:2022:ESD

[FMD^{+}22] Naijie Fan, Gang Mei, Zengyu Ding, Salvatore Cuomo, and Nengxiong Xu. Effects of spatial decomposition on the efficiency of kNN search in spatial interpolations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):103– 121, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fujiwara:2011:SIA

[FNB11] Akihiro Fujiwara, Koji Nakano, and Jacir L. Bordim. Special issue: Advances in parallel and distributed computational models. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(4):249, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Femminella:2015:NIR

[FNRV15] Mauro Femminella, Emilia Nunzi, Gianluca Reali, and Dario Valocchi. Networking issues related to delivering and processing genomic big data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):46–64, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Freitas:2016:TDP

[Fre16] Miguel Freitas. Twister: the development of a peerto-peer microblogging platform. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1): 20–33, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1053808.

Fang:2005:PBB

[FVCD05] Qiu Fang, Susan V. Vrbsky, Hui-Chuan Chen, and Yu Dang. Pull-based broadcasting with timing constraints. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4): 235–252, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648754.

Fan:2020:DAC

[FY20] Liqiang Fan and Jun Yuan.

The diagnosability of kary n-cubes with missing edges. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):57– 68, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Gabriel:2016:MHB

[GACdM16] Paulo H. R. Gabriel, Marcelo K. Albertini, Antonio Castelo, and Rodrigo F. de Mello. Minheap-based scheduling algorithm: an approximation algorithm for homogeneous and heterogeneous distributed systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1):64–84, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1009067.

Gibbins:2008:GIE

[GB08] Hussein Gibbins and Rajkumar Buyya. Gridscape II: An extensible Grid monitoring portal architecture and its integration with Google Maps. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2): 153–170, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a791169878.

Grimes:2019:DCB

[GBMB19] Sean Grimes, Linge Bai, Andrew W. E. McDonald, and David E. Breen. Directing chemotaxis-based spatial self-organisation via biased, random initial conditions. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4):380– 399, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Garnier:2011:CIB

[GG11] Ilias Garnier and Frédéric Gava. CPS implementation of a BSP composition primitive with application to the implementation of algorithmic skeletons. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(4):251–273, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Gouvy:2012:MEC

[GHMSR12] Nicolas Gouvy, Essia Hamouda, Nathalie Mitton, and David Simplot-Ryl. Minimising energy consumption through mobility with connectivity preservation in sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(6):521–540, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Grimaud:2007:UMS

[GHSR07] Gilles Grimaud, Yann Hodique, and Isabelle Simplot-Ryl. On the use of metatypes for safe embedded operating system extension. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(1):1– 13, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a773248289.

Ghafarian:2015:DWS

[GJB15] Toktam Ghafarian, Bahman Javadi, and Rajkumar Buyya. Decentralised workflow scheduling in volunteer computing systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5):343–365, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Garg:2011:LPD

[GKB11] Saurabh Kumar Garg, Pramod Konugurthi, and Rajkumar Buyya. A linear programming-driven genetic algorithm for metascheduling on utility grids. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(6):493–517, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ghadge:2010:SCP

[GKST10] Shilpa Ghadge, Timothy Killingback, Bala Sundaram, and Duc A. Tran. A statistical construction of power-law networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(3):223–235, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Gordon:2006:PAI

[Gor06]  Dan Gordon.  Parallel

ART for image reconstruction in CT using processor arrays. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(5):365– 380, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747881463.

Guarracino:2008:SNE

[GPZ08] Mario R. Guarracino, Francesca Perla, and Paolo Zanetti. A sparse nonsymmetric eigensolver for distributed memory architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3):259– 270, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a790636744.

Ghomari:2007:TPN

[GR07] A. Ghomari and M. K.

Rahmouni. A time Petri net-based approach for synchronization, analysis and management of multimedia scenarios. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6):451– 466, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a784715901.

Garofalakis:2011:AMP

[GS11] John Garofalakis and Eleftherios Stergiou. Analytical model for performanceevaluation of multistage interconnection networks supporting multi-class priority service. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):399– 427, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Goel:2021:CPD

[GSKD21] Lipika Goel, Mayank Sharma, Sunil Kumar Khatri, and D. Damodaran. Crossproject defect prediction using data sampling for class imbalance learning: an empirical study. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):130–143, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Giridhar:2017:AMO

[GSSR17] Munigoti S. Giridhar, S. Sivanagaraju, Chintalapudi V. Suresh, and P. Umapathi Reddy. Analyzing the multi objective analytical aspects of distribution systems with multiple multi-type compensators using modified cuckoo search algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):549– 571, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

ddeke:2007:PAH

[GST07] Dominik Göddeke, Robert Strzodka, and Stefan Turek. Performance and accuracy of hardware-oriented native-, emulated- and mixed-precision solvers in FEM simulations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4): 221–256, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a779508744.

Guo:2006:CVT

[Guo06] Jinhua Guo. Consistency verification in transactionbased optimistic steering of distributed computations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(6):385– 404, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768573890.

Guo:2014:SIC

[Guo14] Song Guo. Special issue on cyber-physical systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3):213–214, 2014.

Gottschling:2009:GSA

[GWJ09] Peter Gottschling, David S. Wise, and Adwait Joshi. Generic support of algorithmic and structural recursion for scientific computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6):479– 503, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hsaini:2021:TBA

[HAC21] Sara Hsaini, Salma Azzouzi, and My El Hassan Charaf. A temporal based approach for MapReduce distributed testing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(4):293–311, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

He:2014:NHT

[HB14] Qian He and Rick S. Blum. New hypothesis testingbased rapid change detection for power grid system monitoring. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3):239– 263, 2014.

Huang:2017:SMH

[HBB^{+}17] Jie Huang, Michel Barbeau, Stéphane Blouin, Craig Hamm, and Martin Taillefer. Simulation and modeling of hydro acoustic communication channels with wide band attenuation and ambient noise. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5):466–485, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Huang:2008:ITU

[HBS08] Yangcheng Huang, Saleem Bhatti, and Søren-Aksel Sørensen. The impact of topology update strategies on the performance of a proactive MANET routing protocol. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(6): 447–460, 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hwang:2005:EEO

[HDK05] Joengmin Hwang, David H. C. Du, and Ewa Kusmierek. Energy efficient organization of mobile sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4):

221–233, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648753.

Hermosillo-Gomez:2019:RET

[HGBP19] J. Angel Hermosillo-Gomez and Hector Benitez-Perez. RUN enhancement through Bayesian networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):523–537, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hong:2013:EEC

[HH13] Won-Sin Hong and Sun-Yuan Hsieh. Extra edge connectivity of hypercubelike networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(2): 123–133, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hamdan:2021:DSA

[HHA21] Salam Hamdan, Amjad Hudaib, and Arafat Awajan. Detecting Sybil attacks in vehicular ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):69– 79, 2021. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Hong:2013:MCB

[HHY13] Yang Hong, Changcheng Huang, and James Yan. Modelling chaotic behaviour of SIP retransmission mechanism. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(2):95–122, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

He:2016:FDA

[HL16] Li He and Furong Li.

A fast diagnosis algorithm for interconnection network in high performance computers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1):34–46, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.958823.

Haase:2007:HOM

[HLP07] Gundolf Haase, Manfred Liebmann, and Gernot Plank. A Hilbert-order multiplication scheme for unstructured sparse matrices. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):213– 220, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a779509037.

Hu:2011:HAI

[HLY11] Xiaomei Hu, Lilan Liu, and Tao Yu. A hierarchical architecture for improving scalability and consistency in CVE systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(3):179–205, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hu:2012:OEP

[HLZ^{+}12] Liang Hu, Lin Lin, Jia Zhao, Xilong Che, and Xiaohui Wei. Optimisation to the execution performance of grid job based on distributed file system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(2):109–121, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hu:2020:TNS

[HMW^{+}20] Wei Hu, Tianao Ma, Yonghao Wang, Fangfang Xu, and Joshua Reiss. TDCS: a new scheduling framework for real-time multimedia OS. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):396– 411, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Holtgen:2020:FYH

[Höl20] Stefan Höltgen. Fifty years in home computing, the digital computer and its private use(er)s. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2): 170–184, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hoadjli:2021:SMC

[HR21] Abir Hoadjli and Khaled Rezeg. A scalable mobile context-aware recommender system for a smart city administration. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):97–116, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hernandez:2010:ECO

[HRF10] Saul Eduardo Pomares Hernandez, Luis Alberto Morales Rosales, and Jean Fanchon. An efficient causal ordering algorithm for multicast communication channels. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):17– 36, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hailat:2014:PMC

[HRR^{+}14] Eyad Hailat, Vincent Russo, Kamel Rushaidat, Jason Mick, Loren Schwiebert, and Jeffrey Potoff. Parallel Monte Carlo simulation in the canonical ensemble on the graphics processing unit. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(4):379– 400, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hoefler:2010:AMO

[HSL10] Torsten Hoefler, Timo Schneider, and Andrew Lumsdaine. Accurately measuring overhead, communication time and progression of blocking and nonblocking collective operations at massive scale. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):241–258, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Heinzl:2009:GPL

[HSSS09] René Heinzl, Philipp Schwaha, Franz Stimpfl, and Siegfried Selberherr. GUIDE: Parallel library-centric application design by a generic scientific simulation environment. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6):505– 520, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Hu:2018:PSD

[Hu18] Fangxia Hu. A probabilistic solution discovery algorithm for solving 0–1 knapsack problem. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6): 618–626, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ho:2008:SLD

[HWS08] Kevin Ho, Jie Wu, and John Sum. On the session lifetime distribution of Gnutella. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(1):1– 15, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a789667014.

Hong:2016:QAM

[HYJ16] Jungha Hong, Taewan You, and Heeyoung Jung. Qualitative analysis of mapping systems in the Internet architectures based on identifier/locator separation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1):85–99, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2015.1044001.

Islam:2011:LAC

[IA11] Kamrul Islam and Selim G. Akl. A local algorithm to compute multiple connected dominating sets in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):369– 380, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Immich:2015:QDV

[IBCC15] Roger Immich, Pedro Borges, Eduardo Cerqueira, and Marilia Curado. QoEdriven video delivery improvement using packet loss prediction. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):478–493, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1044004.

Ihaddadene:2008:EBP

[Iha08] Nacim Ihaddadene. Extraction of business process models from workflow events logs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3):247– 258, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a790794803.

Immanuel:2021:PCA

[IJT^{+}21] S. James Immanuel, S. Jayasankar, D. Gnanaraj Thomas, Meenakshi Paramasivan, Robinson Thamburaj, and Atulya K. Nagar. Parallel contextual array insertion deletion P systems and Siromoney matrix grammars. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(4): 335–358, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ilachinski:2012:MIT

[Ila12] Andrew Ilachinski. Modelling insurgent and terrorist networks as selforganised complex adaptive systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(1):45– 77, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ibarra:2008:CCM

[IP08] Oscar H. Ibarra and Andrei Păun. Computing with cells: membrane systems — some complexity issues. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(5):347–365, 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Indirani:2014:SBE

[IS14] G. Indirani and K. Selvakumar. A swarm-based efficient distributed intrusion detection system for mobile ad hoc networks (MANET). International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(1):90– 103, 2014.

Jiang:2021:PSO

[JCL21] Jehn-Ruey Jiang, Yen-Chung Chen, and Ting-Yu Lin. Particle swarm optimization for charger deployment in wireless rechargeable sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):652– 667, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jendrsczok:2009:SCA

[JEH09] J. Jendrsczok, P. Ediger, and R. Hoffmann. A scalable configurable architecture for the massively parallel GCA model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):275–291, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jedda:2012:TBU

[JJZ12] Ahmed Jedda, Guy-Vincent Jourdan, and Nejib Zaguia. Towards better understanding of the behaviour of Bluetooth networks distributed algorithms. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(6):563–586, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jones:2017:RSM

[JMA17] Jeff Jones, Richard Mayne, and Andrew Adamatzky. Representation of shape mediated by environmental stimuli in Physarum polycephalum and a multiagent model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):166– 184, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jones:2016:AMA

[Jon16] Jeff Jones. Applications of multi-agent slime mould computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):420– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jazayeri:2020:PTD

[JS20] Ali Jazayeri and Hiroki Sayama. A polynomialtime deterministic approach to the travelling salesperson problem. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):454–460, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jia:2005:MST

[JTZX05] Weijia Jia, Wanqing Tu, Wei Zhao, and Gaochao Xu. Multi-shared-trees based multicast routing control protocol using anycast selection. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1):

69–84, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009270.

Jiang:2006:LGI

[JW06] Zhen Jiang and Jie Wu. A limited-global information model for fault-tolerant routing in dual-cube. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(1):61–77, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727649168.

Jiang:2022:RRG

[JW22] Suhan Jiang and Jie Wu.

A reward response game in the blockchain-powered federated learning system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):68– 90, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jiang:2016:BTF

[JWWZ16] Wenjun Jiang, Jie Wu, Guojun Wang, and Huanyang Zheng. Blood typing for families: a novel hybrid human-computer application. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(4):354– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jin:2013:DOC

[JWZL13] Ruofan Jin, Bing Wang, Peng Zhang, and Peter B. Luh. Decentralised online charging scheduling for large populations of electric vehicles: a cyber-physical system approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):29–45, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Jaffar:2011:DDI

[JZZ11] Humzah Jaffar, Xiaobo Zhou, and Liqiang Zhang. DIBS: a dual interval bandwidth scheduling algorithm for short-term differentiation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(3):165– 178, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Krontiris:2012:MIP

[KA12] Ioannis Krontiris and Andreas Albers. Monetary incentives in participatory sensing using multiattributive auctions. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(4):317–336, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Karakostas:2012:IRW

[Kar12] Bill Karakostas. Implementing robust workflow management in Erlang. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(6):503–520, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kassogue:2019:CAA

[KBAO19] Hamidou Kassogué, Abdes Samed Bernoussi, Mina Amharref, and Mustapha Ouardouz. Cellular automata approach for modelling climate change impact on water resources. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):21–36, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Komann:2007:REI

[KF07] Marcus Komann and Dietmar Fey. Realising emergent image preprocessing tasks in cellularautomaton-alike massively parallel hardware. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(2): 79–89, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a768584729.

Khadem:2020:IAP

[KG20] Behrooz Khadem and Reza Ghasemi. Improved algorithms in parallel evaluation of large cryptographic S-boxes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):461–472, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Khedr:2016:DAC

[Khe16] Ahmed M. Khedr. Decomposable algorithm for computing k-nearest neighbours across partitioned data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(4):334– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kuznetsov:2020:BCC

[KHUC20] A. V. Kuznetsov, A. V. Halaimova, M. A. Ufimtseva, and E. S. Chelebieva. Blocking a chemical communication between Trichoplax organisms leads to their disorderly movement. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):473– 482, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kim:2016:RTA

[Kim16] Dohan Kim. Representations of task assignments in distributed systems using Young tableaux and symmetric groups. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(2): 152–175, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kuroda:2017:INC

[KK17] Ryota Kuroda and Seiya Kasai. Implementation of a noise-coexistence threshold logic architecture on a GaAs-based nanowire FET network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):287– 294, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kale:2018:CIA

[KK18] Ishaan R. Kale and Anand J.

Kulkarni. Cohort intelligence algorithm for discrete and mixed variable engineering problems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):627–662, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kiseleva:2018:HSE

[KKK^{+}18] M. A. Kiseleva, E. V. Kudryashova, N. V. Kuznetsov, O. A. Kuznetsova, G. A. Leonov, M. V. Yuldashev, and R. V. Yuldashev. Hidden and self-excited attractors in Chua circuit: synchronization and SPICE simulation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):513– 523, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kulkarni:2021:SOM

[KKK21] Umesh M. Kulkarni, Harish H. Kenchannavar, and Umakant P. Kulkarni. Selforganising map-based dynamic decision-making algorithm for heterogeneous wireless sensor network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(4):312–334, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kulkarni:2018:CCI

[KKKK18] Omkar Kulkarni, Ninad Kulkarni, Anand J. Kulkarni, and Ganesh Kakandikar. Constrained cohort intelligence using static and dynamic penalty function approach for mechanical components design. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6): 570–588, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kanjani:2010:SBF

[KLMW10] Khushboo Kanjani, Hyunyoung Lee, Whitney L. Maguffee, and Jennifer L. Welch. A simple Byzantinefault-tolerant algorithm for a multi-writer regular register. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(5):423– 435, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Khara:2010:GGB

[KMS10] Sibaram Khara, Iti Saha Mishra, and Debashis Saha. Global gatewaybased UMTS/WLAN integration for improved delay performance. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2):153–170, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Koutsogiannis:2018:OFU

[KMT18] Efthymios Koutsogiannis, Lefteris Mamatas, and Vassilis Tsaoussidis. Opportunistic forwarding for user-provided networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):717–741, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Khasymski:2015:SBB

[KN15] Aleksandr Khasymski and Dimitrios S. Nikolopoulos. Scalable black-box prediction models for multidimensional adaptation on NUMA multi-cores. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(3):193–210, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2014.895346.

Konkoli:2018:DTR

[Kon18] Zoran Konkoli. On developing theory of reservoir computing for sensing applications: the state weaving environment echo tracker (SWEET) algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2):121–143, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kuckuk:2013:IPD

[KPK13] Sebastian Kuckuk, Tobias Preclik, and Harald Köstler. Interactive particle dynamics using OpenCL and Kinect. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6):519– 536, 2013.

Kuschel:2017:HPS

[KR17] Christian Kuschel and Ulrich Rüde. Highperformance simulation of nation-sized smart grids. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):647–668, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kohno:2017:CSS

[KT17] Yu Kohno and Tatsuji Takahashi. A cognitive satisficing strategy for bandit problems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):232– 242, 2017. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Kohl:2019:HFE

[KTD^{+}19] Nils Kohl, Dominik Thönnes, Daniel Drzisga, Dominik Bartuschat, and Ulrich Rüde. The HyTeG finiteelement software framework for scalable multigrid solvers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):477– 496, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kuznetsov:2018:MMA

[Kuz18] Alexander V. Kuznetsov. Model of the motion of agents with memory based on the cellular automaton. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):290– 306, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kona:2005:IMA

[KX05] Manoj Kumar Kona and Cheng-Zhong Xu. An integrated mobile agent framework for distributed network management. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1):39–55, ???? 2005. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009271.

Kojecky:2017:ESA

[KZS17] Lumír Kojecký, Ivan Zelinka, and Petr Saloun. Evolutionary synthesis of automatic classification on astroinformatic big data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5):429–447, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kostikov:2021:RDS

[KZS21] Alexander A. Kostikov, Nikolay D. Zaitsev, and Oleg V. Subotin. Realisation of the double sweep method by using a Sleptsov net. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):516– 534, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Kengyel:2019:TSL

[KZWS19] Daniela Kengyel, Payam Zahadat, Franz Wotawa, and Thomas Schmickl. Towards swarm level optimisation: the role of different movement patterns in swarm systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(3): 241–259, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Larrea:2013:IWF

[LAA13] Mikel Larrea, Antonio Fernández Anta, and Sergio Arévalo. Implementing the weakest failure detector for solving the consensus problem. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6):537–555, 2013.

Li:2014:NPP

[LAWS14] Depeng Li, Zeyar Aung, John R. Williams, and Abel Sanchez. No peeking: privacy-preserving demand response system in smart grids. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3):290– 315, 2014.

Lin:2015:DAN

[LC15] Wei-Tee Lin and Chih-Ping Chu. Determining the appropriate number of nodes for fast mining of frequent patterns in distributed computing environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5): 380–392, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Liu:2019:MBF

[LC19] Yuan Liu and Junjie Chen.

A K-means based firefly algorithm for localization in sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4): 364–379, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lin:2018:AAP

[LCG18] Jian (Denny) Lin, Albert M. K. Cheng, and Gokhan Gercek. Approximation algorithms in partitioning real-time tasks with replications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2):211– 232, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lopes:2014:OMI

[LDB^{+}14] Frederico Lopes, Flavia C.

Delicato, Thais Batista, Everton Cavalcante, Thiago Pereira, Paulo F. Pires, Paulo Ferreira, and Reginaldo Mendes. OpenCOPI: middleware integration for ubiquitous computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(2):178–212, 2014.

Levene:2019:HDL

[LFL19] Mark Levene, Trevor Fenner, and George Loizou. Human dynamics with limited complexity. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4): 356–363, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2020:OMA

[LG20] Xiuhong Li and Altenbek Gulila. Optimised memory allocation for less false abortion and better performance in hardware transactional memory. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4): 483–491, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Le:2021:SCG

[LG21] Tuan Le and Mario Gerla. Social contact graph-based group communication schemes for delay tolerant networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(4):380– 394, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2013:DLB

[LGW13] Fan Li, Jinnan Gao, and Yu Wang. Distributed load balancing mechanism for detouring schemes of geographic routing in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(2):184–197, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Liu:2010:HHP

[LH10] Baolin Liu and Bo Hu.

HPRD: a high performance RDF database. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2): 123–133, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lukas:2009:DWM

[LHIN09] Georg Lukas, André Herms, Svilen Ivanov, and Edgar Nett. Dependable wireless mesh networks: an integrated approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(2):151–169, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2020:ECE

[LHL^{+}20] Jin’e Li, Yanze Huang, Limei Lin, Hui Yu, and Riqing Chen. The extra connectivity of enhanced hypercubes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):91– 102, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2005:SDR

[Li05]  Keqin Li.  Scheduling

DAGs with random parallel tasks on binarily partitionable systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(2): 85–97, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a713729128.

Li:2010:DPE

[Li10a] Keqin Li. Design and performance evaluation of communication algorithms in multihop wireless networks with multiple channels. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6):465– 488, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2010:PEH

[Li10b] Keqin Li. Performance evaluation of heuristic algorithms for routing and wavelength assignment in WDM optical networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):273–292, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2010:GEN

[Li10c] Keqiu Li. Guest editorial: Network and parallel computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(2):91–92, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2008:GAS

[LK08] Ming Li and David Kotz.

Group-aware stream filtering for bandwidth-efficient data dissemination. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(6):429–446, 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2008:ASM

[LL08] Yawei Li and Zhiling Lan. Analytical study of migration-enhanced fault tolerance for long-running applications in IFR systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(5):409– 426, 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2018:NAS

[LL18] Xu Li and Jia-Bao Liu. A novel approach to speed up ant colony algorithm via applying vertex coloring. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):608– 617, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Luo:2016:PDM

[LLCZ16] Tao Luo, Yin Liao, Guoliang Chen, and Yunquan Zhang. P-DOT: a model of computation for big data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):233– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Luo:2006:NUT

[LLW06] Junzhou Luo, Yong Lee, and Jun Wu. A nonuniform traffic oriented scheduling algorithm in combined input-crosspoint-queued (CICQ) switches. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4):279–292, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747880317.

Li:2020:SPS

[LLWZ20] Jun Li, Yangkun Luo, Chong Wang, and Zhigao Zeng. Simplified particle swarm algorithm based on nonlinear decrease extreme disturbance and Cauchy mutation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):236– 245, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2020:CBD

[LLZX20] Peng Li, Baozhou Luo, Wenjun Zhu, and He Xu. Cluster-based distributed dynamic cuckoo filter system for Redis. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3): 340–353, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lim:2016:SST

[LMS16] Norman Lim, Shikharesh Majumdar, and Vineet Srivastava. Security sieve: a technique for enhancing the performance of secure sockets layer-based distributed systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):481–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lof:2006:AOC

[LR06] Henrik Löf and Jarmo Rantakokko. Algorithmic optimizations of a conjugate gradient solver on shared memory architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(5):345– 363, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a747881462.

Li:2013:DBG

[LS13] Ze Li and Haiying Shen. A direction-based geographic routing scheme for intermittently connected mobile networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(5):449– 474, 2013.

Lan:2014:ILS

[LS14] Kun-Chan Lan and Wen-Yuah Shih. An indoor location-tracking system for smart parking. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3): 215–238, 2014.

Lewis:2019:FRA

[LS19] Timothy B. Lewis and Quentin F. Stout. A framework for recursive algorithms in low-energy broadcast networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(3): 321–340, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Limmer:2014:SNS

[LSB^{+}14] Steffen Limmer, André Schneider, Christian Boehme, Dietmar Fey, Simon Schmitz, Achim Graupner, and Martin Sülzle. Services for numerical simulations and optimisations in grids. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(6):521–543, 2014. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lipskoch:2015:RDC

[LT15] Kinga Lipskoch and Oliver Theel. Relaxing data consistency along different dimensions for increasing operation availabilities. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(3):233–261, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2014.906599.

Lukasik:2018:QMC

[Luk18] Andrzej Lukasik. Quantum models of cognition and decision. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3): 336–345, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Lampart:2018:DPP

[LVZM18] Marek Lampart, Tomás Vantuch, Ivan Zelinka, and Stanislav Misák. Dynamical properties of partialdischarge patterns. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):474–489, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2007:HSS

[LW07] Xiuqi Li and Jie Wu. A hybrid searching scheme in unstructured P2P networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(1):15– 38, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a773248244.

Li:2021:RSD

[LW21] Yanli Li and Xiaonan Wang. Road-safety data delivery in IP-based vehicular sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3):199–220, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2008:IIS

[LWZ08] Xiuqi Li, Jie Wu, and Shi Zhong. ISRL: intelligent search by reinforcement learning in unstructured peer-to-peer networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(1):17–44, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a789662112.

Liu:2008:ASP

[LX08] X. Liu and Z. Xu. Approaching simple and powerful service-computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):121–135, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a791169488.

Liu:2020:DMD

[LYH^{+}20] Xiaoming Liu, Zhou Yang, Wei Hu, Jun Liu, and Kai Zhang. Detection of macular diseases in optical coherence tomography image. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):260– 272, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Liu:2005:PAG

[LZ05] Huazhou Liu and Qing-An Zeng. Performance analysis of a guard channel reservation handoff scheme for wireless and mobile networks using smart antennas. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4):169–184, ???? 2005. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648750.

Lu:2007:FRA

[LZ07] Enyue Lu and S. Q. Zheng.

Fast reconfiguration algorithms for time, space, and wavelength dilated optical Benes networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(1): 39–58, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a773248163.

Liu:2020:IEE

[LZG^{+}20] Weifeng Liu, Jie Zhou, Bin Gong, Hongjun Dai, and Meng Guo. Improving the energy efficiency of dataintensive applications running on clusters. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3): 246–259, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Liu:2020:BAC

[LZL20] Junnan Liu, Xing Zhai, and Xianfu Liao. Bibliometric analysis on cardiovascular disease treated by traditional Chinese medicines based on big data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):323–339, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Liu:2021:IHP

[LZQL21] Shuo Liu, Kang Zhou, Huaqing Qi, and Jiangrong Liu. Improved hybrid particle swarm optimisation for image segmentation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(1):44–50, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2010:PMT

[LZS^{+}10] Ruixuan Li, Cuihua Zuo, Haiying Shen, Kunmei Wen, and Xiwu Gu. PercolationNET: a multitree P2P overlay network supporting high coverage search. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):73– 89, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2015:VMP

[LZWD15] Kangkang Li, Huanyang Zheng, Jie Wu, and Xiaojiang Du. Virtual machine placement in cloud systems through migration process. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5):393– 410, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mangeas:2019:SEI

[MAG^{+}19] M. Mangeas, J. André, C. Gomez, M. Despinoy, G. Wattelez, and T. Touraivane. A spatially explicit integrative model for estimating the risk of wildfire impacts in New Caledonia. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):37–52, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mahafzah:2011:PMI

[Mah11] Basel A. Mahafzah. Parallel multithreaded IDA* heuristic search: algorithm design and performance evaluation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(1):61– 82, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Merabet:2019:DPT

[MAO19] A. Merabet, A. Ayadi, and A. Omrane. Detection of pollution terms in nonlinear second order wave systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1): 13–20, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Marowka:2009:BCT

[Mar09] Ami Marowka. BSP2OMP: a compiler for translating BSP programs to OpenMP. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):293– 310, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Margenstern:2017:WUC

[Mar17] Maurice Margenstern. A weakly universal cellular automaton on the pentagrid with two states. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):309–330, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Margenstern:2018:NTC

[Mar18] Maurice Margenstern. Navigation tools for cellular automata in two families of tilings of the hyperbolic plane. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):12– 34, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Margenstern:2022:WCW

[Mar22] Maurice Margenstern. What can we learn from universal Turing machines? International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):123–151, 2022. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Migas:2010:AHR

[MB10] N. Migas and W. J.

Buchanan. Ad hoc routing metrics and applied weighting for QoS support. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):293–314, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mian:2010:REE

[MBB10] Adnan Noor Mian, Roberto Baldoni, and Roberto Beraldi. A robust and energy efficient protocol for random walk in ad hoc networks with IEEE 802.11. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(5):363– 379, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Maamri:2018:PAM

[MBCB18] F. Maamri, S. Bououden, M. Chadli, and I. Boulkaibet. The Pachycondyla Apicalis metaheuristic algorithm for parameters identification of chaotic electrical system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):490–502, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mayer:2012:RPS

[MBCK12] Tobias R. Mayer, Lionel Brunie, David Coquil, and Harald Kosch. On reliability in publish/subscribe systems: a survey. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(5):369–386, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mal:2013:EVS

[MCYG13] Siddhartha Mal, Arunabh Chattopadhyay, Albert Yang, and Rajit Gadh. Electric vehicle smart charging and vehicleto-grid operation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(3):249–265, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Margariti:2013:SRF

[MD13] Spiridoula V. Margariti and Vassilios V. Dimakopoulos. A study on the redundancy of flooding in unstructured p2p networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(3):214– 229, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mohamed:2010:EEO

[MDJ10] M. A. Maluk Mohamed, V. R. Devanathan, and D. Janakiram. EOMP: an exactly once multicast protocol for distributed mobile systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(3): 183–207, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Michailidis:2019:EMC

[Mic19] Panagiotis D. Michailidis.

An efficient multi-core implementation of the Jaya optimisation algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(3):288–320, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Misic:2007:DPI

[Miš07] Jelena Mišić. Design and performance issues of slaveslave cluster interconnection in 802.15.4 beacon enabled networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(1): 59–75, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a773248224.

Mohammadian:2021:IET

[MJ21] Masoud Mohammadian and Zafer Javed. Intelligent evaluation of test suites for developing efficient and reliable software. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2): 159–188, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mafarja:2021:WOA

[MJAT21] Majdi Mafarja, Iyad Jaber, Sobhi Ahmed, and Thaer Thaher. Whale Optimisation Algorithm for highdimensional small-instance feature selection. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):80–96, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mayne:2017:COD

[MJGA17] Richard Mayne, Jeff Jones, Ella Gale, and Andrew Adamatzky. On coupled oscillator dynamics and incident behaviour patterns in slime mould Physarum polycephalum: emergence of wave packets, global streaming clock frequencies and anticipation of periodic stimuli. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1): 95–118, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mandal:2015:NPB

[MM15] Monalisa Mandal and Anirban Mukhopadhyay. A novel PSO-based graphtheoretic approach for identifying most relevant and non-redundant gene markers from gene expression data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(3): 175–192, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.889138.

Maiti:2021:CQO

[MMK21] Ananda Maiti, Andrew Maxwell, and Alexander Kist. Comparison of QoS optimisation techniques in adaptive smart device peer-to-peer overlay networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):668– 686, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Matsumoto:2017:NET

[MMS^{+}17] Takuya Matsumoto, Haruka Matsuo, Saki Sumida, Yoshiaki Hirano, Dock-Chil Che, and Hiroshi Ohyama. Nonlinear electric transport in macromolecular system for stochastic computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):252– 258, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Masip:2015:SCN

[MMT^{+}15] Xavi Masip, Abdelhamid Mellouk, Stefano Testa, Eva Marín, Guang-Jie Ren, and Said Hoceini. Smart Communications in Network Technologies. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):426–427, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1076999.

Mongy:2007:SVV

[Mon07] Sylvain Mongy. A study on video viewing behavior: application to movie trailer miner. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3):163– 172, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a776022475.

Motoike:2006:BR

[Mot06] Ikuko Motoike. Book review. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(5):381– 383, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747881461.

Mezhoud:2020:HKM

[MOZH20] Naima Mezhoud, Mourad Oussalah, Abdelouahab Zaatri, and Zoheir Hammoudi. Hybrid kriging and multilayer perceptron neural network technique for coverage prediction in cellular networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6): 682–706, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Marin-Perianu:2008:CBS

[MPSHH08] R. S. Marin-Perianu, J. Scholten, P. J. M. Havinga, and P. H. Hartel. Cluster-based service discovery for heterogeneous wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(4):325– 346, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a793001220.

Mocito:2006:CAC

[MRA^{+}06] Jos Mocito, Liliana Rosa, Nuno Almeida, Hugo Miranda, Lu s Rodrigues, and Ant nia Lopes. Context adaptation of the communication stack. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3): 169–181, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a743847334.

Ma:2016:HHP

[MS16] Xinpei Ma and Hiroki Sayama. Hierarchical heterogeneous particle swarm optimization: algorithms and evaluations. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(5):504–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mackey:2012:UUS

[MSM^{+}12] Grant Mackey, Saba Sehrish, Christopher Mitchell, John Bent, and Jun Wang. USFD: a unified storage framework for SOAR HPC scientific workflows. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(4):347–367, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Martinez:2020:BNH

[MSTMCVL20] Genaro J. Martínez, Juan C.

Seck-Tuoh-Mora, Sergio V. Chapa-Vergara, and Christian Lemaitre. Brief notes and history of computing in Mexico during 50 years. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2):185–192, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mamatas:2008:EES

[MT08] Lefteris Mamatas and Vassilis Tsaoussidis. Exploiting energy-saving potential in heterogeneous networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(4):309– 324, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a793001410.

Munakata:2010:CCU

[MTSA10] Toshinori Munakata, Jun Takahashi, Munehisa Sekikawa, and Kazuyuki Aihara. Chaos computing: a unified view. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):3–16, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ma:2012:MOH

[MTW12] Yonggang Ma, Guozhen Tan, and Wei Wang. A multi-objective hypergraph partitioning model for parallel computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(4): 337–346, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Man:2013:ACE

[MUIN13] Duhu Man, Kenji Uda, Yasuaki Ito, and Koji Nakano. Accelerating computation of Euclidean distance map using the GPU with efficient memory access. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(5):383–406, 2013.

Misic:2008:BPW

[MW08] Vojislav Mišić and Yu Wang.

Best papers from the WWASN2007 workshop. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(6):427–428, 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ma:2020:EVP

[MWHK20] Rongxuan Ma, Yonghao Wang, Wei Hu, and Mahir Payyanil Karalakath. Evaluation of video payload over low latency networks: Flexilink. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):273–287, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nagy:2006:QCQ

[NA06] Marius Nagy and Selim G.

Akl. Quantum computation and quantum information. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(1):1– 59, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727649169.

Nagy:2007:QCB

[NA07] Marius Nagy and Selim G. Akl. Quantum computing: beyond the limits of conventional computation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(2):123–135, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768584067.

Nagy:2012:CUI

[NA12] Naya Nagy and Selim G. Akl. Computing with uncertainty and its implications to universality. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(2):169–192, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nakano:2014:OIA

[Nak14a] Koji Nakano. Optimal implementations of the approximate string matching and the approximate discrete signal matching on the memory machine models. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(2):104– 118, 2014.

Nakano:2014:SMM

[Nak14b]  Koji Nakano.  Simple

memory machine models for GPUs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(1):17– 37, 2014.

Nance:2020:PRL

[Nan20] Richard E. Nance. Personal reflections … on over 50 years in computer simulation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2): 118–131, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nasri:2010:MCP

[Nas10]  Wahid Nasri.  Mod-

elling communication performances in heterogeneous and hierarchical environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):259– 272, 2010. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Nene:2012:UAM

[NDP12] Manisha J. Nene, Rajendra S. Deodhar, and Lalit M. Patnaik. UREA: an algorithm for maximisation of coverage in stochastic deployment of wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(3):249–274, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nishiyama:2013:CBC

[NGA13] Yuta Nishiyama, Yukio-Pegio Gunji, and Andrew Adamatzky. Collisionbased computing implemented by soldier crab swarms. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):67– 74, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Newton:2019:MMP

[NJ19] Todd A. Newton and Eugene B. John. A metric for measuring power efficiency and data throughput in mobile ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(5):556–571, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Novikov:2018:RMM

[NKC18] Dmitry Novikov, Vsevolod Korepanov, and Alexander Chkhartishvili. Reflexion in mathematical models of decision-making. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3): 319–335, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nugent:2018:TRT

[NM18] M. Alexander Nugent and Timothy W. Molter. Thermodynamic-RAM technology stack. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):430– 444, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nandi:2015:HTD

[NNRS15] Mrinal Nandi, Amiya Nayak, Bimal Roy, and Santanu Sarkar. Hypothesis testing and decision theoretic approach for fault detection in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(4):262–285, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2014.910514.

Netto:2020:IRC

[NOdORA20] Hylson Netto, Caio Pereira Oliveira, Luciana de Oliveira Rech, and Eduardo Alchieri. Incorporating the Raft consensus protocol in containers managed by Kubernetes: an evaluation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(4):433–453, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Nosonovsky:2018:LIA

[Nos18] Michael Nosonovsky. Logical and information aspects in surface science: friction, capillarity, and superhydrophobicity. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3): 307–318, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Niu:2021:SPB

[NZCX21] Yunyun Niu, Jieqiong Zhang, Yulin Chen, and Jianhua Xiao. Simulation of pedestrian behaviours in high-rise buildings based on Intelligence Decision P System. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(1):28– 43, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Oya:2017:SIE

[OA17] Takahide Oya and Tetsuya Asai. Special issue of emerging stochastic computing and hardware. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):243, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Oulmahdi:2015:TPL

[OCV15] Mohamed Oulmahdi, Christophe Chassot, and Nicolas Van Wambeke. Transport protocols: limitations, evolution obstacles and solutions for an actual deployment in the Internet. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):515–535, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1053807.

Ogawa:2017:NSS

[Oga17]  Takuji Ogawa.  Non-

symmetric single-molecule electric properties towards stochastic molecular computation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):271– 277, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Othman:2011:ASG

[OGD11] A. Othman, I. Gourlay, and K. Djemame. An adaptive service for grid applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(1):41– 60, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ogston:2010:PPA

[OJ10] Elth Ogston and Stephen A. Jarvis. Peer-to-peer aggregation techniques dissected. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):51– 71, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

ONeill:2017:MAM

[OMKN17] Eoghan O’Neill, John Mc-Glone, Peter Kilpatrick, and Dimitrios Nikolopoulos. Managed acceleration for in-memory database analytic workloads. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4): 406–427, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Oyo:2017:ECV

[OT17] Kuratomo Oyo and Tatsuji Takahashi. Efficacy of a causal value function in game tree search. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):206–217, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Pancerz:2018:PNM

[Pan18] Krzysztof Pancerz. Petri net models for Physarum machines built to realise Boolean functions. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):276–289, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Pokahr:2013:ACA

[PB13] Alexander Pokahr and Lars Braubach. The active components approach for distributed systems development. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(4):321– 369, 2013.

Pereira:2018:NRT

[PdP18] Diogo Augusto Pereira, Wagner Ourique de Morais, and Edison Pignaton de Freitas. NoSQL realtime database performance comparison. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2):144– 156, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Prati:2017:NAT

[PGFA17] Enrico Prati, Ernesto Giussani, Giorgio Ferrari, and Tetsuya Asai. Noiseassisted transmission of spikes in Maeda-Makino artificial neuron arrays. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):278–286, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Privman:2016:LPA

[PGL16] Vladimir Privman, Vyacheslav Gorshkov, and Sergiy Libert. Lattice percolation approach to numerical modelling of tissue aging. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1):1–19, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/

doi/abs/10.1080/17445760.
2015.1057588.

Praxmarer:2007:PCG

[PK07] Peter Praxmarer and Dieter Kranzlmüller. Parallel computing and the Grid—experiences and applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):293– 309, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a779508858.

Parkhomenko:2019:SVC

[PL19] S. S. Parkhomenko and T. M. Ledeneva. Scheduling in volunteer computing networks, based on neural network prediction of the job execution time. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4):430–447, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Privman:2017:P

[Pri17a] Vladimir Privman. Preface. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):1–2, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Privman:2017:TME

[Pri17b] Vladimir Privman. Theoretical modeling expressions for networked enzymatic signal processing steps as logic gates optimized by filtering. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):30–43, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Privman:2017:VP

[Pri17c] Vladimir Privman. Vladimir Privman. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):157– 158, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Puglisi:2017:WUT

[PRMF17] Silvia Puglisi, David Rebollo-Monedero, and Jordi Forné. On web user tracking of browsing patterns for personalised advertising. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5):502–521, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Pluhacek:2020:CEM

[PSVK20] Michal Pluhacek, Roman Senkerik, Adam Viktorin, and Tomas Kadavy. Chaosenhanced multiple-choice strategy for particle swarm optimisation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6):603– 616, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ponce:2009:IPC

[PWL09] Victor Ponce, Jie Wu, and Xiuqi Li. Improve peer cooperation using social networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(3):189– 204, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Qiu:2007:FMS

[QA07] K. Qiu and S. G. Akl. Finding the maximum subsequence sum on interconnection networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(5): 371–385, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a782125751.

Qiu:2020:SIS

[QK20] Meikang Qiu and Sun-Yuan Kung. Special issue on ‘smart computing and communication’ in International Journal of Parallel, Emergent and Distributed Systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):217–218, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Qin:2020:CVC

[QTCM20] Dejin Qin, Yingzhi Tian, Laihuan Chen, and Jixiang Meng. Cyclic vertexconnectivity of Cartesian product graphs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):81–90, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Rawal:2020:PRE

[Raw20] Bharat S. Rawal. Proxy re-encryption architect for storing and sharing of cloud contents. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):219– 235, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Rago:2013:UES

[RBU13] Stephen Rago, Aniruddha Bohra, and Cristian Ungureanu. Using eager strategies to improve NFS I/O performance. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(2):134–158, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Rashid:2019:LAR

[RCS19] Nafiul Rashid, Salimur Choudhury, and Kai Salomaa. Localized algorithms for redundant readers elimination in RFID networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(3):260– 271, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Romano:2009:AVE

[RCSQ09] Paolo Romano, Bruno Ciciani, Andrea Santoro, and Francesco Quaglia. Accuracy versus efficiency of hyper-exponential approximations of the response time distribution of MMPP/M/1 queues. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(2):107–125, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Raman:2010:PTA

[RG10] Vijay Raman and Indranil Gupta. Performance tradeoffs among percolationbased broadcast protocols in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6):509– 530, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ruiz:2006:MAH

[RGS06] Pedro M. Ruiz and Antonio F. Gomez-Skarmeta. Multicast ad hoc routing through mobility-aware Steiner tree meshes with consistency across different mobility models. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4): 257–277, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747880322.

Rijde:2007:SIA

[RH07] Ulrich Rijde and Frank Hülsemann. Special issue: Applied parallel computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):211, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

smpp/content~content=
a779509030.

Richter:2018:SIR

[Ric18] Hendrik Richter. Scaleinvariance of ruggedness measures in fractal fitness landscapes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):460– 473, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ristov:2019:NMC

[RMK^{+}19] Sasko Ristov, Roland Mathá, Dragi Kimovski, Radu Prodan, and Marjan Gusev. A new model for cloud elastic services efficiency. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(6):653– 670, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Rashid:2012:EPS

[RNMQ12] Hammad Rashid, Clara Novoa, Mark McKenney, and Apan Qasem. Efficient parallel solutions to the integral knapsack problem on current chipmultiprocessor systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(1):19–44, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Roqueiro:2007:CPU

[RP07] Damian Roqueiro and Valery A. Petrushin. Counting people using video cameras. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3):193– 209, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a776022382.

Rupanetti:2021:TEA

[RS21] Dulana Rupanetti and Hassan Salamy. Thermal and energy-aware utilisation management on MP-SoC architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):449–469, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Rayati:2015:OOC

[RSR15] Mohammad Rayati, Aras Sheikhi, and Ali Mohammad Ranjbar. Optimising operational cost of a smart energy hub, the reinforcement learning approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(4):325–341, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.974600.

Ren:2020:SNS

[RW20] Yunxia Ren and Shiying Wang. A short note on strong local diagnosability property of exchanged hypercubes under the comparison model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):9–15, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shubbar:2019:ENM

[SA19] Roaa Shubbar and Mahmood Ahmadi. Efficient name matching based on a fast two-dimensional filter in named data networking. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(2):203– 221, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Safonov:2018:CNE

[Saf18] Alexander A. Safonov.

Computing via natural erosion of sandstone. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):742–751, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sanghera:2017:IBD

[SAN^{+}17] Narinder Sanghera, Alexander Anderson, Nick Nuar, Can Xie, Daniel Mitchell, and Judith Klein-Seetharaman. Insulin biosensor development: a case study. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):119–138, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sabaghian:2022:FTS

[SB22] Alireza Sabaghian and Saeed Balochian. Fault tolerant synchronisation of integer and fractional order 6D hyper-chaotic systems via two control signals. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):152–166, 2022. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stepney:2005:JNC

[SBC^{+}05] Susan Stepney, Samuel L. Braunstein, John A. Clark, Andy Tyrrell, Andrew Adamatzky, Robert E. Smith, Tom Addis, Colin Johnson, Jonathan Timmis, Peter Welch, Robin Milner, and Derek Partridge. Journeys in nonclassical computation I: a grand challenge for computing research. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1): 5–19, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009268.

Stepney:2006:JNC

[SBC^{+}06] Susan Stepney, Samuel L.

Braunstein, John A. Clark, Andy Tyrrell, Andrew Adamatzky, Robert E. Smith, Tom Addis, Colin Johnson, Jonathan Timmis, Peter Welch, Robin Milner, and Derek Partridge. Journeys in nonclassical computation II: initial journeys and waypoints. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(2):97– 125, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727649556.

Sarr:2006:NBA

[SCCrL06] Cheikh Sarr, Claude Chaudet, Guillaume Chelius, and Isabelle Gu rin Lassous. A node-based available bandwidth evaluation in IEEE 802.11 ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(6):423–440, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a768574030.

Schumann:2017:CUR

[Sch17] Andrew Schumann. Conventional and unconventional reversible logic gates on Physarum polycephalum. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):218–231, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Schumann:2018:LUC

[Sch18] Andrew Schumann. Logics for unconventional computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):233– 236, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Samanthula:2014:STB

[SCJM14] Bharath K. Samanthula, Hu Chun, Wei Jiang, and Bruce M. McMillin. Secure and threshold-based power usage control in smart grid environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3): 264–289, 2014.

Singh:2015:CDS

[SCK15] Ramesh Singh, Sandip Chakraborty, and Sushanta Karmakar. Concurrent deterministic 1–2 skip list in distributed message passing systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(2):135–174, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2013.876637.

Sulistio:2009:GES

[SCPB09] Anthony Sulistio, Uros Cibej, Sushil K. Prasad, and Rajkumar Buyya. GarQ: an efficient scheduling data structure for advance reservations of grid resources. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):1–19, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Subramani:2007:EAA

[SD07] K. Subramani and D. Desovski. An empirical analysis of algorithms for partially clairvoyant scheduling. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(5):331– 353, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a779055773.

Sibai:2014:PEC

[SEM14] Fadi N. Sibai and Ali El-Moursy. Performance evaluation and comparison of parallel conjugate gradient on modern multi-core accelerator and massively parallel systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(1): 38–67, 2014.

Sakiyama:2017:MEG

[SG17] Tomoko Sakiyama and Yukio-Pegio Gunji. Modulation effect with global ambiguity in 2-dimensional random walk. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):159– 165, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shalinie:2006:MPF

[Sha06] S. Mercy Shalinie. Modeling parallel feed-forward based compression network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4):227– 237, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747880320.

Shapiro:2020:CVL

[Sha20] Linda G. Shapiro. Computer vision: the last 50 years. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2):112– 117, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sahli:2017:BMS

[SHB17] Hamza Sahli, Nabil Hameurlain, and Faiza Belala. A bigraphical model for specifying cloud-based elastic systems and their behaviour. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):593– 616, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Schwaha:2009:SSC

[SHSS09] Philipp Schwaha, René Heinzl, Franz Stimpfl, and Siegfried Selberherr. Synergies in scientific computing by combining multiparadigmatic languages for high-performance applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(6):539– 549, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Siebert:2020:YMG

[Sie20] Rainer Siebert. 50 years (and more) of German computer history. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2): 193–208, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Singh:2017:GVN

[Sin17] Rachhpal Singh. Geneticvariable neighborhood search with thread replication for mobile cloud computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5):486– 501, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Simovici:2006:NMS

[SJ06] Dan A. Simovici and Szymon Jaroszewicz. A new metric splitting criterion for decision trees. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4): 239–256, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747880318.

Safonov:2017:PCT

[SJ17] Alexander Safonov and Jeff Jones. Physarum computing and topology optimisation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(5):448– 465, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sarr:2014:AEE

[SK14] Cheikh Sarr and Sofiane Khalfallah. Average endto-end delay computation in IEEE 802.11 ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(6):544– 561, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sadikin:2015:ITS

[SK15] Mohammad Fal Sadikin and Marcel Kyas. IMAKA-Tate: secure and efficient privacy preserving for indoor positioning applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):447–463, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2015.1058939.

Schumann:2018:FMU

[SK18a] Andrew Schumann and Alexander V. Kuznetsov. Foundations of mathematics under neuroscience conditions of lateral inhibition and lateral activation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):237–256, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shastri:2018:MCI

[SK18b] Apoorva S. Shastri and Anand J. Kulkarni. Multicohort intelligence algorithm: an intra- and intergroup learning behaviour based socio-inspired optimisation methodology. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):675–715, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Srinivasan:2022:SBD

[Sk22] Sathyanarayanan Srinivasan and Ramesh kandukoori. A synod based deterministic and indulgent leader election protocol for asynchronous large groups. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):220– 247, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shivaprakasha:2013:PAE

[SKP13] K. S. Shivaprakasha, Muralidhar Kulkarni, and Raghavendra Patkar. Performance analysis of energy efficient modulation and coding schemes for wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6): 576–589, 2013.

Stojmenovic:2012:BST

[SKZ12] Ivan Stojmenovic, Adnan Afsar Khan, and Nejib Zaguia. Broadcasting with seamless transition from static to highly mobile wireless ad hoc, sensor and vehicular networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(3):225–234, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Suresh:2006:DLS

[SMOK06] S. Suresh, V. Mani, S. N.

Omkar, and H. J. Kim. Divisible load scheduling in distributed system with buffer constraints: genetic algorithm and linear programming approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(5):303–321, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747881465.

Sridevi:2015:OTM

[SMS15] M. Sridevi, C. Mala, and Siddhant Sanyam. Optimised template matching techniques in parallel environment with exhaustive search and swarm intelligence. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(2): 118–134, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/17445760.
2013.844235.

Sharma:2014:END

[SN14] Rashmi Sharma and Nitin. Entropy, a new dynamics governing parameter in real time distributed system: a simulation study. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(6): 562–586, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sabbaghi-Nadooshan:2010:SNH

[SNMSA10] Reza Sabbaghi-Nadooshan, Mehdi Modarressi, and Hamid Sarbazi-Azad. The 2D SEM: a novel highperformance and low-power mesh-based topology for networks-on-chip. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):331–344, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stovold:2017:RDC

[SO17] James Stovold and Simon O’Keefe. Reaction– diffusion chemistry implementation of associative memory neural network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):74– 94, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stepic:2018:LFA

[SO18] Angelina Ilić Stepić and Zoran Ognjanović. Logics to formalise p-adic valued probability and their applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(3):257– 275, 2018. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Satomi:2019:DSM

[SO19] Kota Satomi and Takahide Oya. Design of slime-mold-inspired multi-layered singleelectron circuit. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4): 400–411, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shahrabi:2005:PMB

[SOKM05] A. Shahrabi, M. Ould-Khaoua, and L. M. Mackenzie. Performance modelling of broadcast communication in multicomputer networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(1): 21–37, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a714009266.

Shmeleva:2021:DLS

[SOL21] Tatiana R. Shmeleva, Jan W. Owsiński, and Abdulmalik Ahmad Lawan. Deep learning on Sleptsov nets. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):535–548, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Scarpa:2009:DHQ

[SP09] M. Scarpa and A. Puliafito. Developing high quality software. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(2): 171–187, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Saxena:2005:DHS

[SPBD05] Navrati Saxena, Cristina M.

Pinotti, Kalyan Basu, and Sajal K. Das. A dynamic hybrid scheduling algorithm for heterogeneous asymmetric environments. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4): 185–204, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648751.

Stetsenko:2021:PAD

[SPD21] Inna V. Stetsenko, Alexander A. Pavlov, and Oleksandra Dyfuchyna. Parallel algorithm development and testing using Petriobject simulation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):549–564, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Srinivasan:2012:ICA

[SR12] S. Srinivasan and R. Rajaram. An improved, centralised algorithm for detection and resolution of distributed deadlock in the generalised model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(3):205–224, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stella:2013:ODW

[SR13] I. Johnsi Stella and S. Radha. Optimal design of wireless sensor network for ionising radiation detection using Neyman–Pearson criteria. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):75– 94, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Somasundaram:2011:VTV

[SRG^{+}11] Thamarai Selvi Somasundaram, Kumar Rangasamy, Kannan Govindarajan, Balakrishnan Ponnuraman, Balachandar R. Amarnath, Rajendar Kandan, Rajiv Rajaian, Mahendran Ellappan, Madhusudhanan Bairappan, and Rajesh Britto Gnanaprakasam. VCDE: a toolkit for virtual cluster creation for grid environment. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(6):443–455, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Simplot-Ryl:2009:DAE

[SRTE09] Isabelle Simplot-Ryl, Issa Traoré, and Patricia Everaere. Distributed architectures for electronic cash schemes: a survey. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(3):243–271, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sumathi:2012:QAR

[SS12] R. Sumathi and R. Srinivasan. QoS aware routing protocol to improve reliability for prioritised heterogeneous traffic in wireless sensor network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(2):143–168, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sarma:2013:SAP

[SS13] Mousmita Sarma and Kandarpa Kumar Sarma. Segmentation of Assamese phonemes using hybrid soft-computational approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(4):370–382, 2013.

Singh:2017:FFD

[SS17] Kulwardhan Singh and T. P. Sharma. FDR: fault detection and recovery scheme for wireless sensor networks using virtual grid. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):617– 631, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sharma:2018:EAE

[SS18] Shantanu Sharma and Awadhesh Kumar Singh. An election algorithm to ensure the high availability of leader in large mobile ad hoc networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2):172–196, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Souravlas:2019:TDR

[SS19] Stavros Souravlas and Angelo Sifaleras. Trends in data replication strategies: a survey. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(2): 222–239, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sato:2017:APS

[SSHM17] Hiroshi Sato, Tomohiro Shirakawa, Akitoshi Hagihara, and Kento Maeda. An analysis of play style of advanced mahjong players toward the implementation of strong AI player. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):195–205, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Shirakawa:2017:GAM

[SSS^{+}17] Tomohiro Shirakawa, Naruhisa Sugiyama, Hiroshi Sato, Kazuki Sakurai, and Eri Sato. Gait analysis and machine learning classification on healthy subjects in normal walking. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(2):185–194, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Syed:2018:STS

[SSS18] Mahaboob Shareef Syed, C. V. Suresh, and S. Sivanagaraju. Short term solar insolation prediction: P-ELM approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6): 663–674, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Smaglichenko:2018:SAC

[SSZC18] Tatyana A. Smaglichenko, Alexander V. Smaglichenko, Ivan Zelinka, and Boris Chigarev. Seismic attractor can assist in finding of geothermal area? International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):503–512, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sasaki:2011:MCO

[ST11] Shigero Sasaki and Atsuhiro Tanaka. Modelling coherence overhead of multi-versioned caches for random accesses. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(4):291–311, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stergiou:2021:SMI

[Ste21] Eleftherios Stergiou. A study of multistage interconnection networks operating with wormhole routing and equipped with multi-lane storage. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3):221–239, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stojmenovic:2006:LCO

[Sto06] Ivan Stojmenovic. Listing combinatorial objects in parallel. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(2):127– 146, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727649557.

Stojmenovic:2010:ECV

[Sto10] Ivan Stojmenovic. Editorial: Celebration of the 25th volume. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(1):1–2, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Stojmenovic:2015:E

[Sto15] Ivan Stojmenovic. Editorial. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):1–2, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Suzuki:2005:BIA

[Suz05] Junichi Suzuki. Biologicallyinspired adaptation of autonomic network applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(2):127– 146, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a713729131.

Suzuki:2007:IBM

[Suz07] Yasuhiro Suzuki. An investigation of the Brusselator on the mesoscopic scale. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(2):91–102, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a768585540.

Sun:2021:FLD

[SWW21] Zhang Sun, Qing Wang, and Zhongjun Wei. Fault location of distribution network with distributed generations using electrical synaptic transmissionbased spiking neural P systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(1):11– 27, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Sem:2007:RDL

[SY07] David Sem and Sidney Youlou. Repetitions detection on a linear array with reconfigurable pipelined bus system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3):173–183, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a776022432.

Song:2007:TOP

[SYLS07] Wen-Zhan Song, Fenghua Yuan, Richard LaHusen, and Behrooz Shirazi. Timeoptimum packet scheduling for many-to-one routing in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(5):355– 370, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

                                        smpp/content~content=
                                        a782122888.

Sun:2020:REC

[SZG^{+}20] Xueli Sun, Shuming Zhou, Zhendong Gu, Yihong Wang, and Min Li. Reliability evaluation of complete cubic networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):42–56, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tofigh:2015:NAB

[TAMM15] Tom Tofigh, Sasan Adibi, Amin Mobasher, and Masood Mortazavi. Novel approach to big data collaboration with network operators network function virtualisation (NFV). International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):65–78, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Takano:2017:DES

[TAO17] Makoto Takano, Tetsuya Asai, and Takahide Oya. Design and evaluation of single-electron associative memory circuit. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3): 259–270, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tarakanov:2016:SST

[TB16] Alexander O. Tarakanov and Alla V. Borisova. Sea surface temperature simulation and forecast. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(2):143–151, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tiab:2018:NQA

[TBMB18] Amal Tiab, Louiza Bouallouche-Medjkoune, and Samra Boulfekhar. A new QoS aware and energy efficient opportunistic routing protocol for wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):52– 68, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Talebi:2010:PFF

[TJKY10] Mohammad S. Talebi, Fahimeh Jafari, Ahmad Khonsari, and Mohammad Hossien Yaghmaee. Proportionally fair flow control mechanism for best effort traffic in network-on-chip architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4):345–362, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tyutyunnik:2021:AIC

[TLL21]  Alexander  Tyutyunnik,

Ekaterina Lobaneva, and Alexey Lazarev. Algorithm for identifying clients based on dynamic MAC addresses in narrowly targeted secure networks using deep learning neural networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):470–481, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tinoco:2019:IMS

[TLO19] Claudiney R. Tinoco, Danielli A. Lima, and Gina M. B. Oliveira. An improved model for swarm robotics in surveillance based on cellular automata and repulsive pheromone with discrete diffusion. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1): 53–77, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Takeuchi:2016:CAG

[TNTI16] Yuji Takeuchi, Koji Nakano, Daisuke Takafuji, and Yasuaki Ito. A character art generator using the local exhaustive search, with GPU acceleration. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(1):47–63, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.962026.

Traskas:2011:MAS

[TP11] Dimitrios Traskas and Julian Padget. A multi-agent systems approach to callcentre management. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):347–367, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tran:2011:SIG

[Tra11] Duc A. Tran. Special issue on Grid and P2P Systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(2):83–??, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tyagi:2021:RBD

[TSK21] Shobha Tyagi, Subhranil Som, and S. K. Khatri. Reliability-based dynamic multicast group formation provisioning local adjustment ensuring quality of service globally in MANETs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(2):144– 158, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tantalaki:2020:RBD

[TSR20] Nicoleta Tantalaki, Stavros Souravlas, and Manos Roumeliotis. A review on big data real-time stream processing and its scheduling techniques. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5): 571–601, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Theu:2021:IMR

[TTS^{+}21] Luong Thi Theu, Quang-Huy Tran, Vijender Kumar Solanki, Tatiana R. Shemeleva, and Duc-Tan Tran. Influence of the multi-resolution technique on tomographic reconstruction in ultrasound tomography. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):579– 593, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Trunfio:2019:CSB

[TV19] Paolo Trunfio and Vladimir Vlassov. Clouds for scalable Big Data processing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(6):629– 631, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tran:2020:GAS

[TV20] Duc A. Tran and Quynh Vo. A geo-aware server assignment problem for mobile edge computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6):652–667, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Tran:2016:RFF

[TZG16] Duc A. Tran, Ting Zhang, and Siyuan Gong. A regularization framework for fingerprint-based reconstruction of mobile trajectories. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):268– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Riaz-ud-Din:2016:QVS

[uDZD16] Faizal Riaz ud Din, Wanlei Zhou, and Robin Doss.

Query verification schemes for cloud-hosted databases: a brief survey. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6): 543–561, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Uchida:2014:AAC

[UIN14] Akihiro Uchida, Yasuaki Ito, and Koji Nakano. Accelerating ant colony optimisation for the travelling salesman problem on the GPU. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(4):401– 420, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Urruty:2008:KRM

[Urr08] Thierry Urruty. KpyrRec: a recursive multidimensional indexing structure. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3):235–245, ???? 2008. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a790637363.

Umrao:2016:FTR

[US16] Lokendra Singh Umrao and Ravi Shankar Singh. Faulttolerant routing over shortest node-disjoint paths in hypercubes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):294– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vijayakumar:2012:EAB

[VBA12] V. Vijayakumar, R. S.

D. Wahida Banu, and J. H. Abawajy. An efficient approach based on trust and reputation for secured selection of grid resources. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(1):1–17, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Velykodniy:2021:MBC

[VBZV21] Stanislav Velykodniy, Zhanna Burlachenko, and Svitlana Zaitseva-Velykodna. Modelling the behavioural component of the emergent parallel processes of working with graph databases using Petri net-tools. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):498–515, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Volkov:2017:BMA

[Vol17] Alexander G. Volkov. Biosensors, memristors and actuators in electrical networks of plants. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(1):44–55, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Viswanathan:2014:OTC

[VPS14] N. Viswanathan, K. Paramasivam, and K. Somasundaram. An optimised 3D topology for on-chip communications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(4): 346–362, 2014. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vreeswijk:2017:EDM

[Vre17] Gerard Vreeswijk. The effect of different motion types in simple discrete particle systems with quantitative stigmergy. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):386–405, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vasiliadis:2012:CBW

[VRV12] D. C. Vasiliadis, G. E.

Rizos, and C. Vassilakis. Class-based weighted fair queuing scheduling on quad-priority Delta Networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(5):435– 457, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vasiliadis:2011:MPE

[VRVG11] D. C. Vasiliadis, G. E.

Rizos, C. Vassilakis, and E. Glavas. Modelling and performance evaluation of a novel internal-priority routing scheme for finitebuffered multistage interconnection networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):381–397, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vourkas:2018:ME

[VS18a] Ioannis Vourkas and Georgios Ch. Sirakoulis. Meet the editors. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):445– 447, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vourkas:2018:SIA

[VS18b] Ioannis Vourkas and Georgios Ch. Sirakoulis. Special issue on ‘Advances in Memristive Networks’. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):347–349, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Vatsavai:2007:ESS

[VSB07] Ranga Raju Vatsavai, Shashi Shekhar, and Thomas E. Burk. An efficient spatial semi-supervised learning algorithm. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6): 427–437, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a784715888.

Viktorin:2018:MPR

[VSPK18] Adam Viktorin, Roman Senkerik, Michal Pluhacek, and Tomas Kadavy. Modified progressive random walk with chaotic PRNG. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):450–459, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Watson:2020:HCC

[WA20] Layne T. Watson and Andrew Adamatzky. A half century of computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2):111, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wacker:2019:GET

[Wac19] Simon Wacker. The Garden of Eden theorem for cellular automata on group sets. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(1):78–114, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wallace:2008:TFM

[Wal08] Rodrick Wallace. Toward formal models of biologically inspired, highly parallel machine cognition. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(5):367–408, 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wallace:2009:PCM

[Wal09] Rodrick Wallace. Programming coevolutionary machines: the emerging conundrum. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5):443– 453, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wallace:2010:TEC

[Wal10] Rodrick Wallace. Tunable epigenetic catalysis: programming real-time cognitive machines. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(3): 209–222, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Watson:2020:PRY

[Wat20] Layne T. Watson. Personal reflections on 50 years of scientific computing: 1967–2017. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2):209–215, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2015:RPC

[WB15] Jing Wang and Miodrag Bolic. Reducing the phase cancellation effect in augmented RFID system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(6):494–514, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2015.1035718.

Wang:2016:DAP

[WBP16] MingXue Wang, Kosala Yapa Bandara, and Claus Pahl. A distributed architecture for policycustomisable multi-tenant processes-as-a-service. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(4):377–??, 2016. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2005:NGC

[WCC05] Guojun Wang, Jiannong Cao, and Keith C. C. Chan. A novel group communication protocol using the RingNet hierarchy in mobile Internet. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4): 253–280, ???? 2005. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a727648755.

Wang:2014:NHI

[WCL14] Xiaofei Wang, Ying Cai, and Zhuo Li. A novel hybrid incentive mechanism for node cooperation in mobile cyber-physical systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 29(3):316– 336, 2014.

Wu:2011:DSA

[WCW11] Dan Wu, Chi Hong Cheong, and Man Hon Wong. Distributed snapshots for ad hoc network systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(2):149–164, 2011. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2013:CLN

[WCW13] Jun Wang, Lu Cheng, and Lizhe Wang. Concentric layout, a new scientific data layout for matrix data-set in Hadoop file system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(5):407–433, 2013.

Wang:2015:MSI

[WCY15] Xiaonan Wang, Hongbin Cheng, and Yufeng Yao. Mobility support for IPv6based VANET. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5): 366–379, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wust:2018:PTC

[WFK18] Daniel Wust, Dietmar Fey, and Johannes Knödtel. A programmable ternary CPU using hybrid CMOS/ memristor circuits. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):387–407, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wakrime:2020:MBA

[WJH20] Abderrahim Ait Wakrime, Said Jabbour, and Nabil Hameurlain. A MaxSAT based approach for QoS cloud services. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(6): 641–651, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2020:IED

[WJZK20] Ying Wang, Qiuping Jiang, Qian Zhou, and Yunfeng Kong. Implementation and evaluation distributed mixed pixels analysis algorithm for hyperspectral image based on constraint non-negative matrix factorization. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):365– 375, 2020. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2006:USC

[WL06] Ju Wang and Jonathan Liu. Uplink scheduling in clustered 3G network with mobile relaying. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(6):441– 463, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a768573074.

Wang:2020:IAP

[WLL20] Xiaowei Wang, Kezhi Lv, and Bo Li. IPART: an automatic protocol reverse engineering tool based on global voting expert for industrial protocols. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):376–395, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2012:TRE

[WLWL12] Xiaoyan Wang, Jie Li, Kui Wu, and Huaibei Liu. Transmission rate enhancement via adaptive relaying in wireless networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(3):235–247, 2012. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2021:IIM

[WLX21] Shaolin Wang, Xiyu Liu, and Laisheng Xiang. An improved initialisation method for K-means algorithm optimised by tissue-like P system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(1):3–10, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wu:2009:EMB

[WM09] Jiang Wu and D. Manivannan. An enhanced modelbased checkpointing protocol for preventing useless checkpoints. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5): 383–406, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wu:2013:FIM

[WM13] Jiang Wu and D. Manivannan. A fully informed model-based checkpointing protocol for preventing useless checkpoints. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6): 485–518, 2013.

Wang:2020:DAG

[WM20] Shiying Wang and Xiaolei Ma. Diagnosability of arrangement graphs with missing edges under the MM* model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1): 69–80, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wei:2020:AEW

[WQZZ20] Mengfan Wei, Baojun Qiao, Jianhui Zhao, and Xianyu Zuo. The area extraction of winter wheat in mixed planting area based on Sentinel-2 a remote sensing satellite images. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):297–308, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Won:2013:DBA

[WS13] Myounggyu Won and Radu Stoleru. A destinationbased approach for cut detection in wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(3): 266–288, 2013. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wenisch:2007:CSD

[WvTB^{+}07] Petra Wenisch, Christoph van Treeck, André Borrmann, Ernst Rank, and Oliver Wenisch. Computational steering on distributed systems: Indoor comfort simulations as a case study of interactive CFD on supercomputers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):275– 291, ???? 2007. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a779508970.

Wu:2005:FTD

[WW05] Jie Wu and Dajin Wang.

Fault-tolerant and deadlockfree routing in 2-D meshes using rectilinear-monotone polygonal fault blocks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(2):99–111, ???? 2005. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a713729129.

Wang:2017:RSM

[WW17] Jun Wang and Qilin Wu.

Relay selection for maximizing the number of successive transmission in cooperative networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):632–646, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Walkowiak:2015:ONC

[WWKK15] Krzysztof Walkowiak, Michal Woźniak, Miroslaw Klinkowski, and Wojciech Kmiecik. Optical networks for costefficient and scalable provisioning of big data traffic. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):15– 28, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Wang:2018:ORS

[WWLW18] Ningkui Wang, Zhixuan Wang, Xianming Liu, and Daijun Wei. Optimal route selection based on Monte Carlo method and adaptive amoeba algorithm under uncertain environment. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(2):157–171, 2018. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xue:2012:FNS

[XC12] Weilian Xue and Zhongxian Chi. A flexible node scheduling scheme of minimum delay and energy efficient for wireless sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(2): 123–131, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xie:2013:DLB

[XCW13] Kun Xie, Jiannong Cao, and Jigang Wen. Distributed load-balancing algorithm for fast tag reading. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(5):434– 448, 2013.

Xiao:2008:SOM

[XFCH08] Nong Xiao, Wei Fu, Tao Chen, and Qian Huang. A service-oriented monitoring system with a forecasting algorithm of a time sequence-based hybrid model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):137– 151, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a791170086.

Xiao:2008:GEI

[Xia08] Nong Xiao. Guest Editor’s introduction: Best papers from the GCC 2006 Conference. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(2):105– 106, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

                                        /www.informaworld.com/
                                        smpp/content~content=
                                        a791170372.

Xing:2010:WAH

[XM10] Guoliang Xing and Vojislav B. Misić. Wireless Ad Hoc and Sensor Networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6): 437–438, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xiang:2018:CLB

[XM18] Yande Xiang and Jianyi Meng. A Cross-layer based mapping for spiking neural network onto network on chip. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):526– 544, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xu:2006:SIM

[Xu06] Cheng-Zhong Xu. Special issue on mobile distributed computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(3):149– 150, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a743847338.

Xu:2015:CCB

[XYC^{+}15] Guobin Xu, Wei Yu, Zhijiang Chen, Hanlin Zhang, Paul Moulema, Xinwen Fu, and Chao Lu. A cloud computing based system for cyber security management. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):29– 45, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xu:2020:SPH

[XZL^{+}20] Lizhong Xu, Jia Zhao, Chenming Li, Changli Li, Xin Wang, and Zhifeng Xie. Simulation and prediction of hydrological processes based on firefly algorithm with deep learning and support vector for regression. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):

288–296, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Xu:2020:CCC

[XZY20] Liqiong Xu, Shuming Zhou, and Weihua Yang. Component connectivity of Cayley graphs generated by transposition trees. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1): 103–110, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamashita:2020:HPC

[Yam20] Ryozo Yamashita. History of personal computers in Japan. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(2): 143–169, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamato:2021:AOM

[Yam21a] Yoji Yamato. Automatic offloading method of loop statements of software to FPGA. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(5):482– 494, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamato:2021:SEI

[Yam21b] Yoji Yamato. Study and evaluation of improved automatic GPU offloading method. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6):594– 608, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamato:2022:SEA

[Yam22] Yoji Yamato. Study and evaluation of automatic GPU offloading method from various language applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):22– 39, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yu:2020:GNC

[YHL^{+}20] Hui Yu, Yanze Huang, Limei Lin, Jin’e Li, and Riqing Chen. The g-good-neighbour conditional diagnosability of enhanced hypercube under PMC model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):29– 41, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yakopcic:2018:FMB

[YHT18] Chris Yakopcic, Raqibul Hasan, and Tarek M.

Taha. Flexible memristor based neuromorphic system for implementing multi-layer neural network algorithms. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(4):408– 429, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yang:2012:REO

[YHWZ12] WenZhong Yang, ChuanHe Huang, Bo Wang, and ZhenYu Zhang. Reliable and efficient opportunistic multicast routing for MANETs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 27(3):193– 203, 2012. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

You:2005:PCQ

[YHY05] Tiantong You, Hossam Hassanein, and Chi-Hsiang Yeh. A paradigm for controllable QoS-based medium access control in local and ad hoc wireless networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 20(3–4):281–302, ???? 2005. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/

smpp/content~content=
a727648756.

Ying:2015:PLP

[YM15] Bidi Ying and Dimitrios Makrakis. Protecting location privacy in vehicular networks against locationbased attacks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(2):101–117, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/17445760.
2014.910665.

Yang:2007:GMN

[YMLC07] Xiaofan Yang, Graham M.

Megson, Xiaofeng Liao, and Jianqiu Cao. Generalized matching networks and their properties. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(3):185–192, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a776022501.

Yengisetty:2011:AVC

[YR11] Subba Rao V. Yengisetty and Bimal K. Roy. Applications of visual cryptography. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(5):429– 442, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamagiwa:2009:MPS

[YS09] Shinichi Yamagiwa and Leonel Sousa. Modelling and programming streambased distributed computing based on the metapipeline approach. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):311–330, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yu:2015:SIN

[Yu15] Shui Yu. Special issue on networking aspects in Big Data. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(1):3–4, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yamagiwa:2011:PIR

[YW11] Shinichi Yamagiwa and Koichi Wada. Performance impact on resource sharing among multiple CPU- and GPU-based applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(4):313– 329, 2011. CODEN ????

ISSN 1744-5760 (print), 1744-5779 (electronic).

Yuan:2009:SSM

[YWC09] Quan Yuan, Jie Wu, and Ionut Cardei. SMRS: a scalable multi-path routing scheme. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(1):69– 84, 2009. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yu:2013:OSI

[YWGH13] Yulong Yu, Yuxin Wang, He Guo, and Xubin He. Optimisation schemes to improve hybrid coscheduling for concurrent virtual machines. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):46–66, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yang:2011:CDD

[YYD11] Erjie Yang, Xiaofan Yang, and Qiang Dong. Conditional diagnosability of DC-CLC graphs under the comparison model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(3): 239–248, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yang:2010:DAG

[YYN10] Hui Yang, Xiaofan Yang, and Amiya Nayak. A (4n − 9)/3 diagnosis algorithm for generalised cube networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(3):171– 182, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Yang:2013:CFT

[YZ13] Xiaoxue Yang and Shuming Zhou. On conditional fault tolerant of dual-cubes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(3):199–213, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zenil:2019:CTL

[ZBHOHQ19] Hector Zenil, Liliana Badillo, Santiago Hernández-Orozco, and Francisco Hernández-Quiroz. Coding-theorem like behaviour and emergence of the universal distribution from resourcebounded algorithmic probability. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(2):161– 180, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zheng:2007:LML

[ZCT07] Huicheng Zheng, Pádraig Cunningham, and Alexey Tsymbal. Learning multiple linear manifolds with self-organizing networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(6):417–426, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a786941094.

Zareei:2019:EHM

[ZD19] Sophie Zareei and Jeremiah D.

Deng. Energy harvesting modelling for self-powered fitness gadgets: a feasibility study. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(4):412– 429, 2019. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zelinka:2018:P

[Zel18] Ivan Zelinka. Preface. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(5):449, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2016:OBF

[Zha16]  Wenbing Zhao.  Opti-

mistic Byzantine fault tolerance. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):254– ??, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhang:2018:ECA

[Zha18]  Xiaoge Zhang.  Emer-

gent computing and its applications. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):548– 549, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2020:CED

[Zha20] Liang Zhao. Communicationefficient decentralised algorithms for seismic tomography with sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(5):550– 570, 2020. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhang:2021:MC

[Zha21] Gexiang Zhang. Membrane computing. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(1):1–2, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhou:2006:ISA

[Zho06] Xiaobo Zhou. Internet services and architectures: a performance perspective. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4):303– 304, ???? 2006. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a747880319.

Zhang:2017:DHN

[ZKG17] Bruce Zhang, Laszlo Bela Kish, and Claes-Göran Granqvist. Drawing from hats by noise-based logic. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(3):244–251, 2017. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhang:2021:FDE

[ZM21a] Hong Zhang and Jixiang Meng. Faulty diagnosability and g-extra connectivity of DQcube. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3): 189–198, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2021:SEE

[ZM21b] Shuang Zhao and Jixiang Meng. Super extra edgeconnectivity in regular networks with edge faults. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(3):240–250, 2021. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhu:2021:CCD

[ZM21c] Hongzhou Zhu and Jixiang Meng. Cyclic connectivity of the data center network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6): 623–629, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhou:2013:HRM

[ZMX^{+}13] Jiang Zhou, Can Ma, Jin Xiong, Weiping Wang, and Dan Meng. Highly reliable message-passing mechanism for cluster file system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6):556– 575, 2013.

Zaitsev:2021:PSI

[ZP21] Dmitry A. Zaitsev and David E. Probert. Preface for special issue Petri/ Sleptsov net based technology of programming for parallel, emergent and distributed systems. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 36(6): 495–497, 2021. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhuang:2010:EOG

[ZPW10] Yanyan Zhuang, Jianping Pan, and Guoxing Wu. Energy-optimal grid-based clustering in wireless microsensor networks with data aggregation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(6):531–550, 2010. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2018:DCD

[ZS18] Liang Zhao and WenZhan Song. Decentralized consensus in distributed networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):550– 569, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zaitsev:2022:ACS

[ZSL22] Dmitry A. Zaitsev, Tatiana R. Shmeleva, and Piotr Luszczek. Aggregation of clans to speed-up solving linear systems on parallel architectures. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):198–219, 2022. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zaitsev:2022:SSH

[ZSP22] Dmitry A. Zaitsev, Tatiana R. Shmeleva, and Birgit Pröll. Spatial specification of hypertorus interconnect by infinite and reenterable coloured Petri nets. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(1):1–21, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2018:ABB

[ZSYG18] Liang Zhao, Wen-Zhan Song, Xiaojing Ye, and Yujie Gu. Asynchronous broadcast-based decentralized learning in sensor networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(6):589– 607, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zurawski:2008:FTS

[ZW08] Jason Zurawski and Dajin Wang. Fault-tolerance schemes for clusterheads in clustered mesh networks. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 23(3):271– 287, ???? 2008. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:/

/www.informaworld.com/
smpp/content~content=
a790636068.

Zhou:2020:DEA

[ZWMR20] Zhipeng Zhou, Shiying Wang, Xiaolei Ma, and Yunxia Ren. Diagnosability of expanded k-ary n-cubes with missing edges under the comparison model. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(1):16–28, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhang:2009:RAH

[ZYLH09] Weiwen Zhang, Xiaofan Yang, Yan Liang, and Zhi Huang. Routing algorithms in honeycomb meshes. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(5):367–382, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zou:2017:FFM

[ZYM17] Yuanping Zou, Bidi Ying, and Dimitrios Makrakis. FMSLPP: fake-message based sink location privacy preservation for WSNs against global eavesdroppers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(6):572– 592, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zhao:2020:SER

[ZYW20] Huijuan Zhao, Ning Ye, and Ruchuan Wang. Speech emotion recognition based on hierarchical attributes using feature nets. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 35(3):354–364, 2020. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Zaitsev:2022:SEP

[ZZ22] Dmitry A. Zaitsev and MengChu Zhou. From strong to exact Petri net computers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 37(2):167– 186, 2022. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).